### v1

# Sample Compression for Continual Learning

Abstract
Continual learning algorithms aim to learn from a sequence of tasks, making the training distribution non-stationary. The majority of existing continual learning approaches in the literature rely on heuristics and do not provide learning guarantees for the continual learning setup. In this paper, we present a new method called ‘Continual Pick-to-Learn’ (CoP2L), which is able to retain the most representative samples for each task in an efficient way. The algorithm is adapted from the Pick-to-Learn algorithm, rooted in the sample compression theory. This allows us to provide high-confidence upper bounds on the generalization loss of the learned predictors, numerically computable after every update of the learned model. We also empirically show on several standard continual learning benchmarks that our algorithm is able to outperform standard experience replay, significantly mitigating catastrophic forgetting.
Machine Learning, ICML
1 Introduction
A common assumption in traditional machine learning is that the underlying data distribution does not evolve with time. In Continual Learning (De Lange et al., 2021; Wang et al., 2024), the goal is to develop machine learning algorithms that are able to learn under settings where this assumption is replaced by a set-up where the model is trained on an evolving training data distribution, in such a way that samples are revealed one task at a time.
However, when neural networks are trained on evolving data distributions, networks tend to forget the training from earlier tasks. This fact is known as the catastrophic forgetting (McCloskey & Cohen, 1989; French, 1999; Goodfellow et al., 2014). In order to cope with forgetting, various methodologies have been developed, such as the regularization based approaches (e.g., Kirkpatrick et al., 2017), architectural approaches (e.g., Rusu et al., 2016), or rehearsal based approaches (e.g., Rolnick et al., 2019).
It has been shown on various applications that rehearsal-based approaches typically perform very competitively compared to the other methods in terms of performance and forgetting (Libera et al., 2023; Mai et al., 2020; Chaudhry et al., 2019; Aljundi et al., 2019). However, the selection of the replay buffer typically remains a heuristic, and precise upper bounds for the generalization error are not available.
Our proposed continual learning scheme builds upon the Pick-to-Learn algorithm (P2L) of Paccagnan et al. (2024). P2L is a meta algorithm that selects a compression set, which amounts to express the dataset in terms of ‘core’ examples. This meta-algorithm was developed specifically with sample compression theory (Floyd & Warmuth, 1995; Campi & Garatti, 2023; Laviolette et al., 2005) in mind. By finding a small subset of the data such that a predictor learned on this data achieves low error on the whole training set, Pick-to-Learn enables us to compute tight upper bounds for the generalization error of the learned predictor.
In this paper, we propose ‘Continual Pick-to-Learn‘ (CoP2L), an algorithm that leverages the sample compression theory to intelligently select the training data from earlier tasks to mitigate forgetting. We derive high-confidence upper bounds on the generalization error for each task simultaneously, estimated directly from the training set. We also empirically show that CoP2L mitigates the forgetting significantly, and outperforms the standard replay baseline significantly in the small replay buffer size settings. Our contribution overall can be summarized as follows :
  • We propose the algorithm Continual Pick-to-Learn (CoP2L), that integrates the sample compression theory within the continual learning setup. To the best of our knowledge, we are the first to integrate the theoretical results from sample compression to continual learning.
  • CoP2L is able to provide high confidence non-trivial upper bounds for the generalization error by means of the sample compression theory. We experimentally showcase that the bounds are numerically computable, have non-trivial values and follow the general error trends observed on the test set. Therefore, they can actually be used as risk certificates on the model behavior on the learned tasks.
  • We experimentally show that CoP2L significantly mitigates forgetting, while also being able to obtain better performance compared to vanilla replay, in the settings where the replay buffer is small.
1.1 Related Work
Practical Continual learning
In the continual learning literature, various practical approaches have been developed to mitigate forgetting. Broadly, the approaches could be divided into three categories which include regularization approaches (Kirkpatrick et al., 2017; Zenke et al., 2017; Li & Hoiem, 2018; Aljundi et al., 2018), architecture-based approaches (Rusu et al., 2016; Mallya et al., 2018; Aljundi et al., 2017; Mallya & Lazebnik, 2018; Wang et al., 2022) and rehearsal based approaches (Rolnick et al., 2019; Chaudhry et al., 2019; Shin et al., 2017; Rebuffi et al., 2017; Buzzega et al., 2020). In this paper, we use experience replay as a baseline as it is shown in the literature to be a very strong continual learning baseline to mitigate forgetting (Libera et al., 2023).
Theoretical Continual Learning
There have been several works in theoretical continual learning in the recent years. For example, Bennani & Sugiyama (2020); Doan et al. (2021) study the generalization error obtained with Neural Tangent Kernel (NTK) models when the model is trained with Orthogonal Gradient Descent (Farajtabar et al., 2020).
Yin et al. (2020) provides generalization bounds for regularization-based continual learning methods. Asanuma et al. (2021) studies the impact of task similarity in the student-teacher setup. Li et al. (2022) provides a sample complexity bound in the context of continual representation learning for an adapted PackNet (Mallya & Lazebnik, 2018). In Chen et al. (2022), the authors provide a theoretical bound on memory using PAC learning theory. These works however do not provide bounds that are computable in practice.
A line of work also investigates the generalization of continual learning models when linear models are employed. In Goldfarb & Hand (2023), the authors investigate the effect of overparametrized linear models. In Evron et al. (2022); Lin et al. (2023), the authors analyze the generalization behavior of linear regression models, even though Lin et al. (2023) conjectures that their results could be extended to neural networks as well.
Different from all the listed works above, in this paper, we provide a practical non-trivial generalization bound that is computed on the training set, and predicts the test error for each task separately. Furthermore, also unlike the approaches above, our bound is applicable to any neural network architecture and does not have limitations on the continual learning setup that is employed.
Sample Compression
The sample compress theory has been shown effective for providing tight generalization bounds (Laviolette et al., 2005; Marchand et al., 2003; Marchand & Shawe-Taylor, 2002; Marchand & Sokolova, 2005; Shah, 2007). However, most sample compress approaches are limited to low-complexity models. A notable exception is the method Pick-to-Learn (P2L), which successfully provides guarantees for deep neural networks (Campi & Garatti, 2023; Paccagnan et al., 2024). In this paper, we adapt the key ideas introduced in the sample compression theory to the continual learning case, which enables us to provide an upper bound on the test error from training samples.
2 Background
2.1 Continual Learning
Notation.
Let us consider a series of task distributions ${\mathcal{D}}_{1},{\mathcal{D}}_{2},\ldots,{\mathcal{D}}_{T}$ on an input-output space ${\mathcal{X}}\times{\mathcal{Y}}$, where the amount of tasks $T\in{\mathbb{N}}^{*}$ can be (countably) infinite. We are given a predictive model $f_{\theta}:{\mathcal{X}}\to{\mathcal{Y}}$ (e.g., a neural network architecture), with learnable parameters $\theta\in\Theta$. From randomly initialized parameters $\theta_{0}$, the aim of the continual learning process at step $t\in\{1,\ldots,T\}$ is to update the parameters $\theta_{t-1}$ into $\theta_{t}$ by learning from a new task sample $S_{t}\sim{\mathcal{D}}_{t}$.
We denote the training set at task $t$ as $S_{t}=\{({\mathbf{x}}_{t,i},y_{t,i})\}_{i=1}^{n_{t}}$, with ${\mathbf{x}}_{t,i}\in{\mathcal{X}}$ and $y_{t,i}\in{\mathcal{Y}}_{t}\subseteq{\mathcal{Y}}$; that is, the same input space ${\mathcal{X}}$ is shared among all tasks, and the output space ${\mathcal{Y}}$ may be the union of several task specific output spaces ${\mathcal{Y}}_{t}$.
Given a loss function $\ell:\Theta\times{\mathcal{X}}\times{\mathcal{Y}}\to{\mathbb{R}}^{+}$, we want the last updated predictor $f_{\theta_{T}}$ to maintain a low generalization loss on all observed tasks $t\in\{1,\ldots,T\}$:
${\mathcal{L}}_{\operatorname{\mathcal{D}}_{t}}(\theta_{T})\ \coloneqq\operatorname*{\mathbf{E}}_{({\mathbf{x}},y)\sim{\mathcal{D}}_{t}}\ell(\theta_{T},{\mathbf{x}},y)\,.$ (1)
The empirical loss counterparts on observed training samples are given by
^$\widehat{\mathcal{L}}_{S_{t}}(\theta_{T})\ \coloneqq\ \frac{1}{n_{t}}\sum_{i=1}^{n_{t}}\ell(\theta_{T},{\mathbf{x}}_{t,i},y_{t,i})\,.$ (2)
The challenge of continual learning lies in the fact that the learner observes the task datasets $S_{1},S_{2},\ldots,S_{t},\ldots,S_{T}$ sequentially, and we assume that the system cannot keep all observed data in memory. Nevertheless, at task $t$, we would like the system to perform well on all tasks, including the previous ones $11,\dots,t-$ 1. However, simply updating the parameters $\theta_{t-1}$ learned from a previous tasks to optimize the loss ^$\widehat{\mathcal{L}}_{S_{t}}(\theta_{t})$ on a current task $t$ would lead to catastrophic forgetting (French, 1999). Therefore, as indicated in Section 1.1, numerous empirical methods have been developed to mitigate forgetting. A simple yet very effective strategy is to simply store a small subset of data from the earlier tasks, coined as the replay buffer, and to include them in the objective function of the experience replay strategy for learning task $t$:
${\mathcal{F}}_{t}^{\ \text{replay}}(\theta_{t})\ \coloneqq\ \widehat{\mathcal{L}}_{S_{t}}(\theta_{t})+\frac{1}{|\mathcal{B}|}\sum_{j\in\mathcal{B}}\ell(\theta_{T},{\mathbf{x}}_{j},y_{j})\,,$ (3)
where the second term is the loss on the replay buffer ${\mathcal{B}}$. In our experiments, we use this method as the baseline as it’s shown to be a simple and very effective continual learning strategy (Libera et al., 2023; Mai et al., 2020).
2.2 Sample Compression Theory
Given a dataset $S=\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{n}$, a family of learnable parameters $\Theta$ and a learning algorithm $A$ such that $A(S)\in\Theta$, sample compression theory provides generalization bounds for any predictor $A(S)$ on the condition that the predictor can be provably represented as a function of a small subset of the dataset, called the compression set, and an additional source of information, called the message. If this is the case, we call $A(S)$ a sample-compressed predictor.
The compression set is denoted $S^{\operatorname{\mathbf{i}}}$ and is parameterized by a vector of indices $\operatorname{\mathbf{i}}\subset\{1,\ldots,n\}$. Each vector $\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)$, with $\operatorname{\mathscr{P}}(n)$ the set of all the possible subsets of $\{1,\ldots,n\}$, is ordered such that we have
$\operatorname{\mathbf{i}}=(i_{1},i_{2},\ldots,i_{\operatorname{|\mathbf{i}|}})\,,\mbox{ with }1\leq i_{1}<\ldots<i_{\operatorname{|\mathbf{i}|}}\leq n\,.$ (4)
The compression set is defined such that
$S^{\operatorname{\mathbf{i}}}=S^{(i_{1},\ldots,i_{\operatorname{|\mathbf{i}|}})}\!=\{(\boldsymbol{x}_{i_{1}},y_{i_{1}}),\ldots,(\boldsymbol{x}_{i_{\operatorname{|\mathbf{i}|}}},y_{i_{\operatorname{|\mathbf{i}|}}})\}\!\subseteq S.$ (5)
We define the complement vector $\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}$ such that $\operatorname{\mathbf{i}}\cap\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}=\emptyset$ and $\operatorname{\mathbf{i}}\cup\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}=\{1,\ldots,n\}$. Thus, we have the complement set $S_{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}=S\setminus S_{\operatorname{\mathbf{i}}}$ and $|\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}|=n-\operatorname{|\mathbf{i}|}$.
In addition to the compression set, a message $\mu$ is sometimes needed to compress the predictor. Although generally defined as a binary sequence, the message set can be defined by any set of countable sequences of symbols. Let $\Sigma=\{\sigma_{1},\sigma_{2},\ldots,\sigma_{N}\}$ be the alphabet used to construct the messages and $\Sigma^{*}$ be the set of all possible sequences, of length 00 to $\infty$, constructed using the alphabet $\Sigma$. For all $\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)$, we choose $\mathscr{M}(\operatorname{\mathbf{i}})$ a countable subset of $\Sigma^{*}$, which represents all the possible messages that can be chosen for $\operatorname{\mathbf{i}}$.
Given a learning algorithm $A$ and a dataset $S$, to prove that the predictor $A(S)$ is a sample-compressed one, we need to define two functions: a compression function and a reconstruction function. The compression function $\mathscr{C}:\bigcup_{1\leq n\leq\infty}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{n}\to\bigcup_{m\leq n}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{m}\times\bigcup_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\mathscr{M}(\operatorname{\mathbf{i}})$ outputs a compression set and a message that are used to represent $A(S)$. Given a dataset $S$, we have $\mathscr{C}(S)=(S^{\operatorname{\mathbf{i}}},\mu)$. Then, the reconstruction function $\operatorname{\mathscr{R}}:\bigcup_{m\leq n}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{m}\times\bigcup_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\mathscr{M}(\operatorname{\mathbf{i}})\to\Theta$ is defined such that $A(S)=\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},\mu)$. Both the reconstruction function and the compression function must be deterministic and data-independent.
The forthcoming results rely on a probability distribution over the set of sample-compressed predictors $\overline{\Theta}\subseteq\Theta$. For any sample-compressed predictor $\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)$, this data-independent distribution is expressed as $P_{\overline{\Theta}}\quantity(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu))=P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{i}})}(\mu)$, with $P_{\operatorname{\mathscr{P}}(n)}$ a probability distribution over $\operatorname{\mathscr{P}}(n)$ and $P_{\mathscr{M}(\operatorname{\mathbf{i}})}$ a probability distribution over $\mathscr{M}(\operatorname{\mathbf{i}})$. Following the work of Marchand & Sokolova (2005), we consider $P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})=\smallmatrixquantity(n\\ \operatorname{|\mathbf{i}|})^{-1}\zeta(\operatorname{|\mathbf{i}|})$, with $\zeta(k){=}\tfrac{6}{\pi^{2}}(k+1)^{-2}$. Given a vector of indices $\operatorname{\mathbf{i}}$, we consider $P_{\mathscr{M}(\operatorname{\mathbf{i}})}$ to be a uniform distribution over the messages.
We now present the sample compression bound of Bazinet et al. (2025). We use the shorthand notation ^$\operatorname{\widehat{\mathcal{L}}}_{S}^{\,{}_{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}}(\cdot)$ to denote the empirical loss on the training sample that don’t belong to the compression set: $S^{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}=S\setminus S^{\mathbf{i}}$; this is mandatory to obtain an unbiased estimate of the loss, as the reconstructed predictor $\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)$ relies on $S^{\mathbf{i}}$.
Theorem 2.1.
For any distribution $\operatorname{\mathcal{D}}$ over $\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}$, for any family of set of messages $\{\mathscr{M}(\operatorname{\mathbf{i}})|\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)\}$, for any deterministic reconstruction function $\operatorname{\mathscr{R}}$, for any loss $\ell:\Theta\times\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}\to[0,1]$ and for any $\delta\in(0,1]$, with probability at least $1-\delta$ over the draw of $S\sim\operatorname{\mathcal{D}}^{n}$, we have
$\forall\mathbf{i}\in\operatorname{\mathscr{P}}(n),\,\mu\in\mathscr{M}(\operatorname{\mathbf{i}}):$
$\mathrm{kl}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,{}_{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}}\quantity(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)),\mathcal{L}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)))\leq\epsilon(n,\operatorname{\mathbf{i}},\mu,\delta)$
with $\mathrm{kl}(q,p)=q\ln\frac{q}{p}+(1-q)\ln\frac{1-q}{1-p}$ the binary Kullback-Leibler divergence and
$\epsilon(n,\operatorname{\mathbf{i}},\mu,\delta)=\frac{1}{n-\operatorname{|\mathbf{i}|}}\quantity[\log\!\matrixquantity(n\\ \operatorname{|\mathbf{i}|})\!\!+\!\log\!\quantity(\frac{2\sqrt{n-\operatorname{|\mathbf{i}|}}}{\zeta(\operatorname{|\mathbf{i}|})P_{\mathscr{M}(\operatorname{\mathbf{i}})}(\mu)\delta})].$
Given the parameters $\theta$ of a reconstructed predictor, i.e., $\theta{=}\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)$, Theorem 2.1 states that the $\mathrm{kl}$-discrepancy between the true risk ^$\operatorname{\widehat{\mathcal{L}}}_{S}^{\,{}_{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}}(\theta)$ and the empirical risk $\mathcal{L}_{\operatorname{\mathcal{D}}}(\theta)$ is upper bounded by the complexity term $\epsilon(n,\operatorname{\mathbf{i}},\mu,\delta)$. From this result, one can compute an upper bound on the true risk $\mathcal{L}_{\operatorname{\mathcal{D}}}(\theta)$ by inverting the Kullback-Leibler divergence. Indeed, Theorem 2.1 can be rewritten as
$\mathcal{L}_{\operatorname{\mathcal{D}}}\quantity(\theta)\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,{}_{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}}\quantity(\theta),\epsilon(n,\operatorname{\mathbf{i}},\mu,\delta))$ (6)
with
$\mathrm{kl}^{-1}(q,\varepsilon)\ =\ \operatorname*{arg\,sup}_{0\leq p\leq 1}\{\mathrm{kl}(q,p)\leq\varepsilon\}.$ (7)
On the one hand, given a fixed compression set size $\operatorname{|\mathbf{i}|}$, the bound decreases when the training set size $n$ increase. On the other hand, given a fixed $n$, the bound increases with $\operatorname{|\mathbf{i}|}$.
2.3 Pick-to-Learn
To obtain sample compression guarantees for a class of predictors, it is necessary to prove that the learned predictor only depends on a small subset of the data and (optionally) a message. Some predictors, such as the SVM (Boser et al., 1992) and the perceptron (Rosenblatt, 1958; Moran et al., 2020), have straightforward compression scheme, as they only depend on a subset of the data after training. Some algorithms, such as the set covering machine (SCM) (Marchand & Shawe-Taylor, 2002; Marchand et al., 2003; Marchand & Sokolova, 2005) and decision trees (Shah, 2007), necessitate hand-crafted compression schemes involving a message. Up until recently, there was no sample compression scheme for deep neural networks.
The meta-algorithm Pick-To-Learn (P2L) was proposed by Paccagnan et al. (2024) as a compression scheme for any class of predictors, with a specific focus to deep neural networks. The meta-algorithm modifies the training loop of the predictor, by iteratively choosing datapoints over which the model is updated. Starting with initial parameters $\theta_{\rm init}$, P2L evaluates the predictor on the whole training dataset, adds the datapoints with the highest losses to the compression set and then updates the predictor using the compression set. The meta-algorithm stops once the losses of training examples not in the compression set are lower than a given stopping criteria ($-\ln(0.5)$ for the cross-entropy loss, which is equivalent to achieving zero errors on the complement set $S^{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}$). We provide the pseudo-code of Pick-To-Learn in Algorithm 1.
Algorithm 1 Original Pick-To-Learn (P2L)
0:  $S=\{({\mathbf{x}}_{i},y_{i})\}_{i=1}^{n}$Training set}
0:  $\theta_{\rm init}$Initialization parameters of a given model $f_{\theta}$}
0:  $\gamma$Stopping criteria}
1:  $k\leftarrow 0$ ; $C_{0}\leftarrow\emptyset$ ; $\theta_{0}\leftarrow\theta_{\rm init}$
2:  $(\overline{{\mathbf{x}}},\overline{y})\leftarrow\operatorname*{\mathrm{argmax}}_{({\mathbf{x}},y)\in S}\ell(\theta_{0},{\mathbf{x}},y)$
3:  while $\ell(\theta_{k},\overline{{\mathbf{x}}},\overline{y})\geq\gamma$ do
4:     $k\leftarrow k+1$
5:     $C_{k}\leftarrow C_{k-1}\cup\{(\overline{{\mathbf{x}}},\overline{y})\}$
6:     $\theta_{k}\leftarrow\mbox{update}(\theta_{k-1},C_{k})$
7:     $(\overline{{\mathbf{x}}},\overline{y})\leftarrow\operatorname*{\mathrm{argmax}}_{({\mathbf{x}},y)\in S\setminus C_{k}}\ell(\theta_{k},{\mathbf{x}},y)$
8:  end while
9:  return $\theta_{k},C_{k}$.
10:  {Learned parameters and compression set}
3 Sample Compression for Continual Learning
3.1 A New Training Scheme
A striking realization coming from the P2L algorithm is that only a small fraction of the training set – the compression set – needs to be provided to the learner in order to achieve good generalization. The winning strategy is to select this compression set to ensure a low risk on the training samples not being retained in the compression set (i.e., the complement of the compression set). This observation motivates our strategy for managing the replay buffer of our new algorithm Continual Pick-To-Learn (CoP2L): we subsample the complement set to create the replay buffer, in order to maintain a low risk on these examples while learning subsequent tasks. When needed, CoP2L can choose to add a datapoint from the buffer, thus mitigating the forgetting by adding a well-chosen datapoint.
Algorithm 2 Modified Pick-To-Learn (mP2L)
0:  $\theta_{\rm init}$Initialization parameters of a given model $f_{\theta}$}
0:  $S=\{({\mathbf{x}}_{i},y_{i},w_{i})\}_{i=1}^{n}$Training set (with weights)}
0:  $B^{\star}=\{({\mathbf{x}}_{i},y_{i},w_{i})\}_{i=1}^{m}$Buffer set (with weights)}
0:  $\gamma$Stopping criteria}
0:  $K^{\star}$Maximum number of iterations}
1:  $k\leftarrow 0$ ; $C_{0}\leftarrow\emptyset$ ; $\theta_{0}\leftarrow\theta_{\rm init}$
2:  $S^{\star}\leftarrow S\cup B^{\star}$
3:  $(\overline{{\mathbf{x}}},\overline{y},\overline{w})\leftarrow\operatorname*{\mathrm{argmax}}_{({\mathbf{x}},y,w)\in S^{\star}}\ell(\theta_{0},{\mathbf{x}},y)\cdot w$
4:  while $\ell(\theta_{k},\overline{{\mathbf{x}}},\overline{y})\cdot\overline{w}\geq\gamma$ and $k\leq K^{\star}$ do
5:     $k\leftarrow k+1$
6:     $C_{k}\leftarrow C_{k-1}\cup\{(\overline{{\mathbf{x}}},\overline{y},\overline{w})\}$
7:     $\theta_{k}\leftarrow\mbox{update}(\theta_{k-1},C_{k})$
8:     $(\overline{{\mathbf{x}}},\overline{y},\overline{w})\leftarrow\operatorname*{\mathrm{argmax}}_{({\mathbf{x}},y,w)\in S^{\star}\setminus C_{k}}\ell(\theta_{k},{\mathbf{x}},y)\cdot w$
9:  end while
10:  if $k<K^{\star}$ then
11:     $k\leftarrow\operatorname*{\mathrm{argmin}}_{0\leq k^{\prime}\leq k}\Psi(S,\theta_{k^{\prime}},C_{k^{\prime}}\cap S)$
12:  end if
13:  return $\theta_{k},C_{k}$.
14:  {Learned parameters and compression set}
Algorithm 3 Continual Pick-To-Learn (CoP2L)
0:  $S_{1},S_{2},\ldots,S_{T}$Training sets}
0:  $\theta_{0}$Initialization parameters of a given model}
0:  $\gamma$P2L’s stopping criteria}
0:  $m$Buffer’s max sampling size}
0:  $\omega$Weight for buffer tasks}
1:  $B_{i}\leftarrow\emptyset\quad\forall i=1,\ldots,T$
2:  $B^{\star}\leftarrow\emptyset$
3:  for $t\in\{1,\ldots,T\}$ do
4:     $\hat{S_{t}}\leftarrow\{({\mathbf{x}},y,1)\}_{({\mathbf{x}},y)\in S_{t}}$
5:     $\theta_{t},C^{\star}\leftarrow\mbox{mP2L}(\theta_{t-1},\hat{S_{t}},B^{\star},\gamma,\infty)$
6:     $B_{i}\leftarrow\mbox{sample}(B_{i},\lfloor\frac{m}{t}\rfloor)\quad\forall i=1,\ldots,t-1$
7:      $B_{t}\leftarrow\mbox{sample}(\hat{S}_{t}\setminus C^{\star},\lfloor\frac{m}{t}\rfloor)$
8:     $B^{\star}\leftarrow\bigcup_{i=1}^{t}\{({\mathbf{x}},y,\omega)\}_{({\mathbf{x}},y,\cdot)\in B_{i}}$
9:  end for
10:  return $\theta_{T}$. {Learned parameters}
Each new task is learned by a modified version of the P2L meta-algorithm (Algorithm 2, entitled mP2L), as explained further down. That is, for each new task, the proposed continual learning algorithm CoP2L calls mP2L and updates the buffer using the datapoints that were not chosen in the compression set. Similarly to the replay buffer method, at the end of each task $t$, the buffer contains $\lfloor\tfrac{m}{t}\rfloor$ datapoints of each previous task, with $m$ the maximum size of the buffer. We present CoP2L in Algorithm 3.
We modified Pick-To-Learn in two significant ways. First, we introduced weights to the loss functions to tackle the imbalance problem between the current task and the previous tasks. When working with a replay buffer, the class imbalance problem is taken care of by training on a number of datapoints from previous tasks. As our model is only trained on a very limited subset of the data, we need to find another way to mitigate the effect of class imbalance. Before starting the training on a new task, we set the weight of the datapoints from the previous task to $\omega$ and the weight of the current task to $1$. For the stopping criteria of mP2L to be satisfied, the worst loss on the current task must be less than $\gamma$ and the worst loss on the previous task must be less than $\tfrac{\gamma}{\omega}$. This ensures that the model has achieved perfect accuracy on the current task and all the datapoints from the previous tasks that are found in the buffer.
The second modification leading to mP2L is the choice of the returned predictor. The original P2L algorithm trains the model until it first achieves zero errors on the complement set. In presence of a noisy dataset, this might lead to an overfitting behavior. Instead, to perform early stopping, mP2L relies on the trade-off encoded in Theorem 2.1 between the accuracy on the complement set and the complexity of the model. More precisely, it returns the model’s parameters $\theta$ that minimizes
$\Psi(S,\theta,C)$ $=\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S\setminus C}\quantity(\theta),\ \tfrac{1}{|S\setminus C|}\log\quantity(\tfrac{2\sqrt{|S\setminus C|}{|S|\choose|C|}}{\zeta(|C|)\delta})),$ (8)
with $\mathrm{kl}^{-1}$ defined by Equation 7. Thanks to Theorem 2.1, we have that $\operatorname{\mathcal{L}}_{\mathcal{D}}(\theta)\leq\Psi(S,\theta,C)$ with probability at least $1-\delta$.
Note that for the first observed task, the mP2L procedure starts from randomly initialized parameters $\theta_{0}$ and is executed on the training sample $S_{1}$ (as it is done by the original P2L). Then, for every subsequent task $t\in\{2,\ldots,T\}$, the mP2L procedure is initialized to the previously learned parameters $\theta_{t-1}$. It then learns from the current task sample $S_{t}$ and a subset of randomly selected instances from previous tasks. The latter is obtained from the procedure sample$(S,m)$ (see Algorithm 3), which represents the random sampling of $m$ instances of $S$ without replacement.
From a theoretical standpoint, in contrast to the original P2L, both mP2L and CoP2L cannot be reconstructed straightforwardly without a message for two reasons: (1) the bound function $\Psi(S,\theta,C)$ cannot be computed on the whole dataset, when only the compression set is given as input; (2) the sampling procedure sample$(S,m)$ cannot be reproduced in the reconstruction step. In the following section, we define the compression and reconstruction functions of CoP2L, which will give rise to sample compression bounds as presented in Theorem 3.1.
3.2 Compression and Reconstruction Scheme
To obtain generalization bounds for CoP2L, we need to prove that it can be compressed. To do so, we provide the compression and reconstruction functions of CoP2L. The compression function is presented in Algorithm 4 (provided in supplementary material) and the reconstruction function is presented in Algorithm 5.
Compression function.
In line with the prevalent literature, we presented the sample compression framework (Section 2.2) based on a compression function providing a compression set and optionally a message. In the case of CoP2L, the reconstruction scheme relies on two compression sets $(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}})$ and a message pair $(\mu_{1},\mu_{2})$, which we explain later in this section. [^1] The idea of using multiple compression sets appeared in Marchand & Shawe-Taylor (2002) and Marchand et al. (2003), but in a setting without a message Two challenges prevent CoP2L from being used as its own compression and reconstruction function as the original P2L.
The first challenge comes from the sampling of the buffer ${\mathcal{B}}$. When training using CoP2L, datapoints from the buffer can be chosen to be part of the compression set. At the next task, these datapoints may still be available to train the model (if they are not excluded from the buffer by the sampling step). Therefore, the reconstruction function needs to know when to remove the datapoint. The message thus provide the task after which the datapoint was removed during CoP2L execution. As the reconstruction function only requires a message for datapoints who were removed from the buffer after being added to the compression set $S^{\operatorname{\mathbf{i}}}$, a second compression set $S^{\operatorname{\mathbf{j}}}$ is dedicated to datapoints that necessitate a message. The message $\mu_{1}$ given is chosen among $\{2,\ldots,T\}$ for each datapoint belonging to $S^{\operatorname{\mathbf{j}}}$.
The second one is the use of the bound as stopping criterion in mP2L. Recall that the reconstruction function does not have access to the whole dataset. Thus, it cannot recover the bound value computed during the initial training phase (see Line 11 of Algorithm 2) and use it as a stopping criterion. To address this inconvenience, the number of iterations to perform is provided to mP2L as a message $\mu_{2}=(\mu_{2}^{1},\ldots,\mu_{2}^{T})$. That is, for each task $t$, with a buffer $\mathcal{B}$, the message component $\mu_{2}^{t}$ is chosen in $\{1,\ldots,n_{t}+|\mathcal{B}|\}$, giving the number of mP2L’s iterations to perform.
After learning on $T$ tasks, the compression function provides the sample compression set $S^{\operatorname{\mathbf{i}}}$ and  $S^{\operatorname{\mathbf{j}}}$, along with a message pair $(\mu_{1},\mu_{2})$ chosen among the set of all possible messages, denoted as
$\mathscr{M}(\operatorname{\mathbf{j}},T)\,=\,\{2,\ldots,T\}^{|\operatorname{\mathbf{j}}|}\times\left[\bigtimes\limits_{t=1}^{T}\{1,\ldots,n_{t}+|{\mathcal{B}}|\}\right].$ (9)
Reconstruction function.
Under the assumption that the input parameters $\theta_{0},\gamma,m,\omega$ are the same for the compression function and for CoP2L, the reconstruction function must output the exact same predictor as CoP2L. However, it only has access to the compression sets and the messages. The first challenge of the compression function is addressed using $\mu_{1}$ from line 6 to 11 of Algorithm 5. For each datasets, we verify if a datapoint was previously excluded from the buffer by the sampling function. If so, we remove it from the buffer. The second challenge is adressed in line 4 of Algorithm 5 by giving the message $\mu_{2}$ to mP2L as stopping criterion.
Algorithm 4 Compression function of CoP2L
0:  $\theta_{0}$Initialization parameters of a given model}
0:  $\gamma$P2L’s stopping criteria}
0:  $m$Buffer’s max sampling size}
0:  $\omega$Weight for buffer tasks}
0:  $S_{1},S_{2},\ldots,S_{T}$Training sets}
1:  $B_{i}\leftarrow\emptyset,C_{i}\leftarrow\emptyset,C^{\mu}_{i}\leftarrow\emptyset,\mu_{1,i}\leftarrow\emptyset\quad\forall i=1,\ldots,T$
2:  $B^{\star}\leftarrow\emptyset,\mu_{2}\leftarrow\emptyset$
3:  for $t\in\{1,\ldots,T\}$ do
4:     $\hat{S_{t}}\leftarrow\{({\mathbf{x}},y,1)\}_{({\mathbf{x}},y)\in S_{t}}$
5:     $\theta_{t},C^{\star}\leftarrow\mbox{mP2L}(\theta_{t-1},\hat{S_{t}},B^{\star},\gamma,\infty)$
6:     $\mu_{2}^{t}\leftarrow|C^{\star}|$
7:     $C_{i}\leftarrow C_{i}\cup(C^{\star}\cap S_{i})\quad\forall i=1,\ldots,t$
8:     $B_{i}\leftarrow\mbox{sample}(B_{i},\lfloor\frac{m}{t}\rfloor)\quad\forall i=1,\ldots,t-1$
9:     $B_{t}\leftarrow\mbox{sample}(S_{t}\setminus C^{\star},\lfloor\frac{m}{t}\rfloor)$
10:     for $i\in\{1,\ldots,t-1\}$ do
11:        for $({\mathbf{x}}_{i,j},y_{i,j},\cdot)\in B^{\star}$ and $({\mathbf{x}}_{i,j},y_{i,j},\cdot)\notin B_{i}$ do
12:           $C^{\mu}_{i}\leftarrow C^{\mu}_{i}\cup\{({\mathbf{x}}_{i,j},y_{i,j})\}$
13:           $C_{i}\leftarrow C_{i}\setminus\{({\mathbf{x}}_{i,j},y_{i,j})\}$
14:           $\mu_{1,i}^{j}\leftarrow\{t\}$
15:        end for
16:     end for
17:     $B^{\star}\leftarrow\bigcup_{i=1}^{t}\{({\mathbf{x}},y,\omega)\}_{({\mathbf{x}},y,\cdot)\in B_{i}}$
18:  end for
19:  return $\theta_{T},\{C_{t}\}_{t=1}^{T},\{C^{\mu}_{t}\}_{t=1}^{T},\{M_{t}\}_{t=1}^{T},N$. {Learned parameters, compression sets and messages sets}
Algorithm 5 Reconstruction function of CoP2L
0:  $\theta_{0}$Initialization parameters of a given model}
0:  $\gamma$P2L’s stopping criteria}
0:  $m$Buffer’s max sampling size}
0:  $\omega$Weight for buffer tasks}
0:  $C_{1},C_{2},\ldots,C_{T},\ C^{\mu}_{1},C^{\mu}_{2},\ldots,C^{\mu}_{T}$Compression sets}
0:  $\mu_{1,1},\mu_{1,2},\ldots,\mu_{1,T},\,\mu_{2}${Message sets}
1:  $B^{\star}\leftarrow\emptyset$
2:  for $t\in\{1,\ldots,T\}$ do
3:     $\hat{S_{t}}\leftarrow\{({\mathbf{x}},y,1)\}_{({\mathbf{x}},y)\in C_{t}}$
4:     $\theta_{t},C^{\star}\leftarrow\mbox{mP2L}(\theta_{t-1},\hat{S_{t}},B^{\star},\gamma,\mu_{2}^{t})$
5:     $B_{t}\leftarrow(C_{t}\setminus C^{\star})\cup C^{\mu}_{t}$
6:     for $i\in\{1,\ldots,t-1\}$ do
7:        for $(\boldsymbol{x}_{i,j},y_{i,j},\omega_{i})\in C^{\mu}_{i}$ do
8:           if $\mu_{1,i}^{j}=t$ then
9:              $B_{i}\leftarrow B_{i}\setminus\{(\boldsymbol{x}_{i,j},y_{i,j},\omega_{i})\}$
10:           end if
11:        end for
12:        $B^{\star}\leftarrow\bigcup_{i=1}^{t}\{({\mathbf{x}},y,\omega)\}_{({\mathbf{x}},y,\cdot)\in B_{i}}$
13:     end for
14:  end for
15:  return $\theta_{T}$. {Learned parameters}
3.3 Generalization Bound
The last subsection provided the compression and reconstruction functions of CoP2L. Thus, we know CoP2L provides a sample-compressed predictor. Theorem 2.1 can be used to obtain generalization bounds on the last task, which we do in mP2L when we compute the bound $\Psi$, but it holds for only one distribution of data. Therefore, Theorem 2.1 cannot be applied to bound all tasks learned by our continual learning scheme.
Le us now consider a reconstruction function that takes as input two compression sets $\operatorname{\mathbf{i}},\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)$ and a message $\mu=(\mu_{1},\mu_{2})\in\mathscr{M}(\operatorname{\mathbf{j}},T)$. For any distribution $\operatorname{\mathcal{D}}_{t}$ and any dataset $S_{t}\sim\operatorname{\mathcal{D}}_{t}$, we denote the loss on both compression sets as ^$\widehat{{\mathcal{L}}}_{S_{t}}^{\ \operatorname{\mathbf{i}}}(\theta)$ and ^$\widehat{{\mathcal{L}}}_{S_{t}}^{\ \operatorname{\mathbf{j}}}(\theta)$. The loss on the joint compression sets is denoted ^$\widehat{{\mathcal{L}}}_{S_{t}}^{\ \operatorname{\mathbf{i}}\cup\operatorname{\mathbf{j}}}(\theta)$ and the loss on the complement set of the joint set $\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}$ is denoted ^$\widehat{{\mathcal{L}}}_{S_{t}}^{\ \operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\!(\theta)$. Moreover, we denote the reconstruction function of CoP2L as $\operatorname{\mathscr{R}}_{1:T}(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)=\operatorname{\mathscr{R}}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu;S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}).$ This formulation is important for the following theorem, as the reconstruction function should be fixed for all datasets $S_{1}$ to $S_{T}$, with the exception of $S_{t}$.
Theorem 3.1.
For any set of independent distributions $\left\{{\mathcal{D}}_{t}\right\}_{t=1}^{T}$ over ${\mathcal{X}}\times{\mathcal{Y}}$, with $\operatorname{\mathscr{R}}_{1:T}$ the reconstruction function of CoP2L (Algorithm 5), with the family of set of messages $\{\mathscr{M}(\operatorname{\mathbf{j}},T)|\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)\}$ defined for CoP2L, for any loss $\ell:\Theta\times{\mathcal{X}}\times{\mathcal{Y}}\to[0,1]$, and for any $\delta\in(0,1]$, with probability at least $1-\delta$ over the draw of $S_{t}\sim{\mathcal{D}}_{t},t=1,\ldots,T$, we have
$\forall t\in\{1,\ldots,T\},{\mathbf{i}},\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n_{t}),\mu\in\mathscr{M}(\operatorname{\mathbf{j}},T):$
$\mathrm{kl}\quantity(\widehat{{\mathcal{L}}}_{S_{t}}^{\ \operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\!\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)),\ {\mathcal{L}}_{\operatorname{\mathcal{D}}_{t}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)))$
$\leq\frac{1}{n_{t}-|{\mathbf{i}}|-|\mathbf{j}|}\bigg{[}\log\matrixquantity(n_{t}\\ |{\mathbf{i}}|)+\log\matrixquantity(n_{t}-\operatorname{|\mathbf{i}|}\\ |\mathbf{j}|)$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}+\sum_{i=1}^{T}\log\frac{1}{\zeta(\mu_{2}^{i})}+|\mathbf{j}|\log(T-1)+\log\frac{T}{\zeta(|{\mathbf{i}}|)\zeta(|\operatorname{\mathbf{j}}|)\delta}\bigg{]}$
with $\mu=(\mu_{1},\mu_{2})$.
The proof is given in Appendix B. It first relies on a tighter version of Theorem 2.1, where the $2\sqrt{n-\operatorname{|\mathbf{i}|}}$ term is removed and a second compression set is added. Then, this new result is adapted to the continual learning setting.
For all tasks $1$ through $T$, this new theorem encodes a tradeoff similar to Theorem 2.1. A more accurate predictor will achieve a tighter bound. However, a simpler predictor will achieve a tighter bound.
In the sample compression theory, we define the complexity of the model using the size of the compression sets $\operatorname{\mathbf{i}}$ and $\operatorname{\mathbf{j}}$ and the probability of choosing a message. The probability of the messages $\mu_{1}$ and $\mu_{2}$ are functions of the size of $\operatorname{\mathbf{j}}$ and the number of tasks. Indeed, the probability of $\mu_{1}$ is independent of its content but decays when $|\operatorname{\mathbf{j}}|$ and $T$ grow larger. Moreover, the probability of $\mu_{2}$ depends on the size of the compression set outputted by mP2L at each task $T$. Thus, the larger the compression sets are, the smaller the probability of both messages will be and the complexity of the model grows when the compression sets grow.
In conclusion, the model that will achieve the best possible bound is a model that minimizes the loss on the dataset whilst still keeping $\operatorname{\mathbf{i}}$ and $\operatorname{\mathbf{j}}$ the smallest possible.
4 Experiments
Figure 1: Accuracy (above) and forgetting (below) over replay buffer size for (left column) MNIST dataset, (middle column) Fashion-MNIST dataset, (right column) EMNIST dataset. The orange triangles show the average performance for the corresponding box plot.
Figure 2: CoP2L bounds with respect to tasks on the MNIST and Fashion-MNIST datasets
Datasets We carry-out experiments on MNIST (LeCun et al., 2010), Fashion-MNIST (Xiao et al., 2017), and EMNIST (Cohen et al., 2017) datasets on class-incremental learning. We divide the tasks into increments of 2 classes. This yields 5 tasks for MNIST and Fashion-MNIST datasets, and 13 tasks for the EMNIST dataset.
Implementation Details We have used the Avalanche continual learning framework for our experimental setup (Carta et al., 2023). Avalanche is a popular toolkit for continual learning. We have used their class-incremental MNIST and Fashion-MNIST dataset implementations, as well as the replay baseline in order to ensure compatibility with the results in the literature. We implement a block version of mP2L, which adds $k$ examples to the compression set at a time, as described in Algorithm 2 of Paccagnan et al. (2024). For the experiments, we use $k=8$. We provide the code for our experimental results in an anonymous repository [^2] https://anonymous.4open.science/r/sample_compression_continual_learning-00C7..
Model architecture and training details We use a simple multilayer perceptron (MLP) architecture for all experiments (SimpleMLP class from the Avalanche toolkit), and train using the SGD optimizer, with a learning rate of 0.001 for CoP2L and 0.01 for Replay. For all experiments with the Replay method, we train the model 20 epochs on each tasks. In contrast, CoP2L follows a different training strategy, continuing until it reaches zero errors on the complement set. We add a weight of $\omega=15$ in the cross-entropy loss when training on previous tasks.
Computation of the bound To compute the bound, we need to keep track of the size of the compression sets and of the messages. We thus implement the compression function as described in Algorithm 4. However, we need to adjust the code to account for mP2L adding $k$ datapoints to the compression set at each iterations. When computing the message $\mu_{2}$, at iteration $t$, $\mu_{2}^{t}$ is the number of iterations of the mP2L algorithm. Knowing that mP2L outputs a compression set $C^{\star}_{t}$, and adds $k$ datapoints to the compression set at each iteration, the number of iterations is $\mu_{2}^{t}=\frac{|C^{\star}_{t}|}{k}$. Finally, when computing the bound, we use the code provided by Viallard et al. (2021) to invert the $\mathrm{kl}$ divergence.
4.1 Metric Comparison over Buffer Size
In Figure 1, we compare the accuracy and forgetting obtained with different replay buffer sizes of (1000, 2000, 3000, 4000, 5000) samples for CoP2L and replay. This represents roughly 8% to 40% of the per-task dataset size, for MNIST and Fashion-MNIST, and 10% to 52% for EMNIST. For each replay buffer size, we report the variation of accuracy obtained at the end of training averaged over different seeds, and test accuracy for different tasks. Note that we denote the accuracy obtained on a task $t$, after training the model for $T$ tasks with $\text{Accuracy}_{T,t}$. The accuracy we report is defined as
$\text{Average Accuracy}_{T}=\frac{1}{s\cdot T}\sum_{i=1}^{s}\sum_{t=1}^{T}\text{Accuracy}_{T,t}^{i},$ (10)
where $i$ index is for different seeds, and $t$ is for the task index. In our experiments, we used $s=5$ seeds to measure the variability with respect to weight initialization. We also report the average forgetting at task $T$ which is defined as follows,
$\text{Average Forgetting}_{T}=$ (11)
$\frac{1}{(T-1)\cdot s}\sum_{t=1}^{T-1}\sum_{i=1}^{s}$ $\text{Accuracy}_{t,t}^{i}-\text{Accuracy}_{T,t}^{i},$
and can be calculated when $T>1$. Note that in Figure 1, we report the average accuracy obtained after the final task, such that $T=5$ for MNIST and Fashion-MNIST and $T=13$ for EMNIST.
We observe that for the MNIST dataset, on average CoP2L significantly outperforms replay (as evidenced by the variation over seeds). For the Fashion-MNIST and EMNIST datasets, we observe that for small buffer sizes, the average accuracy is obtained with our CoP2L is significantly better. For larger buffer sizes, in general CoP2L performs comparably or slightly better.
We would like to also note for all three datasets, CoP2L significantly outperforms replay in terms of forgetting. We attribute this due to the fact that CoP2L is able to intelligently select examples from the replay buffer. Namely, CoP2L is able to focus on examples that are most representative of the earlier tasks, which we observe to help significantly in mitigating the deterioration over earlier tasks as evidenced by the results presented in the right column in Figure 1.
4.2 Empirical Study of the CoP2L generalization bound
In Figure 2, we showcase the generalization bound of CoP2L. We observe that the bound values for the generalization error always remain larger than the actual error incurred after each task, and we observe that the reported bound values are non-trivial. We also observe that the bound values generally follow similar trends as the actual error we report on the test set, while the estimate for the bound is calculated exclusively on the training set. In Figure 2, we provide the bounds for the models trained with a buffersize of 1000, but we give the results for all buffersizes, and all three datasets in Appendix C.
Figure 3: Ablation study on MNIST and Fashion-MNIST on the additional weighting to ensure class balance we propose in Algorithm 2 on the CoP2L loss function. Accuracy (left) and forgetting (right) over replay buffer size for the MNIST dataset
Figure 4: Studying the variation of cumulative accuracy with respect to task ordering for CoP2L and Replay on MNIST dataset (left) and Fashion-MNIST dataset (right). We use 5 different task order, and show the variability with respect to each task.
4.3 Ablation on CoP2L loss terms
We have conducted an ablation study to understand the effect of additional weighting that counteracts class imbalance that arises in the continual learning setup when new tasks are introduced (Standard replay counteracts this by extending the batch size to have equal class representation). In the modified P2L algorithm (2), we incorporate an additional weighting term over classes to adapt P2L for class-balanced continual learning. In order to assess whether this idea is indeed helpful, we have run an experiment on the MNIST dataset with and without the additional weighting idea. In Figure 3, we show this experiment. We observe that for both accuracy (left panel) and forgetting (right panel), the inclusion of the additional loss weighting term significantly improves the results.
4.4 Sensitivity analysis with respect to the task ordering
In order to also assess the sensitivity of CoP2L with respect to ordering of the tasks, on MNIST and Fashion-MNIST experiments we have conducted an experiment where we randomly shuffle the task order. For both dataset, we have tested our algorithm with 5 different random task ordering.
For each task, we report the cumulative accuracy
[…]
where $\text{Accuracy}_{T,t}^{o}$ denotes the accuracy obtained at task $t$ after finishing training until task $T$, and the $o$ index denotes the task order. We have used $O=5$ different task orderings. We present the results of this experiment in Figure 4. We observe that on both MNIST and Fashion-MNIST datasets CoP2L is able outperform with more tasks. We also observe that the variability of the results decrase more significantly for CoP2L compared to replay.
5 Conclusions
In this paper, we have proposed CoP2L, an algorithm rooted in the sample compression theory. To the best of our knowledge, this is the first attempt in the literature to employ sample compression theory within the continual learning context. We provided sample compression bounds for CoP2L and verified empirically that they were non-vacuous. We furthermore showed that on three different datasets, our approach is able to outperform experience replay, which is known to be a strong continual learning baseline.
Acknowledgements
We also wish to thank Benjamin Leblanc for his help proof-reading the manuscript. Mathieu Bazinet is supported by a FRQNT B2X scholarship (343192). Pascal Germain is supported by the NSERC Discovery grant RGPIN-2020-07223. Cem Subakan is supported by NSERC Discovery grant RGPIN-2023-05759. This research was also enabled in part by support provided by the Digital Research Alliance of Canada.
Appendix A Tighter sample compression $\mathrm{kl}$ bound
In this section, we prove a tighter $\mathrm{kl}$ bound that is not derived from Theorem 2.1 but proved in a very similar way. To do so, we start by providing the Chernoff test-set bound (Langford, 2005) for losses in $[0,1]$ (Foong et al., 2022).
Theorem A.1 ((Foong et al., 2022)).
Let $X_{1},\ldots,X_{n}$ be i.i.d. random variables with $X_{i}\in[0,1]$ and $\operatorname*{\mathbb{E}}[X_{i}]=p$. Then, for any $\delta\in(0,1]$, with probability at least $1-\delta$
$p\leq\mathrm{kl}^{-1}\quantity(\frac{1}{n}\sum_{i=1}^{n}X_{i},\frac{1}{n}\log\frac{1}{\delta}).$
We now prove a tighter $\mathrm{kl}$ sample compression bound. As our algorithm CoP2L needs two compression sets, we consider a second compression set $\operatorname{\mathbf{j}}$. We define a conditional probability distribution $P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})$ that incorporates the knowledge that a vector $\operatorname{\mathbf{i}}$ was already drawn from $\operatorname{\mathscr{P}}(n)$ and that $\operatorname{\mathbf{i}}\cap\operatorname{\mathbf{j}}=\emptyset$. Thus, we have $\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})\leq 1.$ If the choice of $\operatorname{\mathbf{j}}$ isn’t conditional to the choice of $\operatorname{\mathbf{i}}$, we simply have $P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})=P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}})$.
Theorem A.2.
For any distribution $\operatorname{\mathcal{D}}$ over $\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}$, for any family of set of messages $\{\mathscr{M}(\operatorname{\mathbf{j}})\,|\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)\}$, for any deterministic reconstruction function $\operatorname{\mathscr{R}}$ , for any loss $\ell:\Theta\times\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}\to[0,1]$ and for any $\delta\in(0,1]$, with probability at least $1-\delta$ over the draw of $S\sim\operatorname{\mathcal{D}}^{n}$, we have
$\forall\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n),\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n),\mu\in\mathscr{M}(\operatorname{\mathbf{j}}):$
$\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta})$
Proof.
We prove the complement of the expression in Theorem A.2. We choose $\delta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}=P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta$.
$\operatorname*{\mathbb{P}}_{S\sim\operatorname{\mathcal{D}}^{n}}\quantity(\exists\operatorname{\mathbf{i}},\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n),\mu\in\mathscr{M}(\operatorname{\mathbf{j}}):\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))>\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{\delta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}}))$
$\leq\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\operatorname*{\mathbb{P}}_{S\sim\operatorname{\mathcal{D}}^{n}}\quantity(\exists\mu\in\mathscr{M}(\operatorname{\mathbf{j}}):\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))>\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{\delta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}}))$ (13)
$\leq\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\sum_{\mu\in\mathscr{M}(\operatorname{\mathbf{j}})}\operatorname*{\mathbb{P}}_{S\sim\operatorname{\mathcal{D}}^{n}}\quantity(\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))>\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{\delta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}}))$ (14)
$\leq\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\sum_{\mu\in\mathscr{M}(\operatorname{\mathbf{j}})}\operatorname*{\mathbb{E}}_{S^{\operatorname{\mathbf{i}}}\sim\operatorname{\mathcal{D}}^{\operatorname{|\mathbf{i}|}}}\operatorname*{\mathbb{E}}_{S^{\operatorname{\mathbf{j}}}\sim\operatorname{\mathcal{D}}^{|\operatorname{\mathbf{j}}|}}\operatorname*{\mathbb{P}}_{S^{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\sim\operatorname{\mathcal{D}}^{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}}\quantity(\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))>\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{\delta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}}))$ (15)
$\leq\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\sum_{\mu\in\mathscr{M}(\operatorname{\mathbf{j}})}\operatorname*{\mathbb{E}}_{S^{\operatorname{\mathbf{i}}}\sim\operatorname{\mathcal{D}}^{\operatorname{|\mathbf{i}|}}}\operatorname*{\mathbb{E}}_{S^{\operatorname{\mathbf{j}}}\sim\operatorname{\mathcal{D}}^{|\operatorname{\mathbf{j}}|}}\delta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}$ (16)
$\leq\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\sum_{\mu\in\mathscr{M}(\operatorname{\mathbf{j}})}\delta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}$
$\leq\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\sum_{\mu\in\mathscr{M}(\operatorname{\mathbf{j}})}P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta$
$\leq\delta.$
We use the union bound in Equation 13 and Equation 14. In Equation 15, we use the i.i.d. assumption. In Equation 16 we use the Chernoff Test-set bound of Theorem A.1 with $p=\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))=\operatorname*{\mathbb{E}}_{(\boldsymbol{x},y)\sim\operatorname{\mathcal{D}}}\ell(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu),\boldsymbol{x},y)$ and $X_{i}=\ell(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu),\boldsymbol{x}_{i},y_{i})\forall i\in\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}$. Finally, the last line is derived from the definition of $P_{\operatorname{\mathscr{P}}(n)}$ and $P_{\mathscr{M}(\operatorname{\mathbf{j}})}$. ∎
Appendix B Proof of Theorem 3.1
We now prove Theorem 3.1.
See 3.1
Proof.
Let us choose $t\in[1,T]$. Let us sample $S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}$. We denote the reconstruction function of CoP2L as $\operatorname{\mathscr{R}}_{1:T}(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)=\operatorname{\mathscr{R}}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu;S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}).$
Then, we have :
$\operatorname*{\mathbb{P}}_{S^{t}}\Bigg{(}\begin{split}&\forall\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n_{t}),\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n_{t}),\mu\in\mathscr{M}(\operatorname{\mathbf{j}}):\\ &\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n_{t}-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta})\end{split}\Biggr{)}$
As all the datasets (except $S^{t}$) are sampled beforehand, we can define $\operatorname{\mathscr{R}}_{1:T}(\cdot)=\operatorname{\mathscr{R}}\quantity(\cdot;S^{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T})$ before drawing $S^{t}$. Thus, the reconstruction function is only a function of the dataset $S^{t}$, which is the setting of the result of Theorem A.2. We can then lower bound this probability by $1-\delta$.
We finish the proof by applying the bound to all datasets.
With
$\epsilon(n_{k},\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu,\delta)=\frac{1}{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta},$
we have
$\operatorname*{\mathbb{P}}_{S_{1},\ldots,S_{T}}\Bigg{(}$ $\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)),\epsilon(n_{t},\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu,\delta))\Biggr{)}$
$=\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{T}}\operatorname{\mathbb{I}}\Bigg{(}$ $\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)),\epsilon(n_{t},\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu,\delta))\Biggr{)}$
$=\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}}\operatorname*{\mathbb{E}}_{S_{t}}\operatorname{\mathbb{I}}\Bigg{(}$ $\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)),\epsilon(n_{t},\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu,\delta))\Biggr{)}$
$=\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}}\operatorname*{\mathbb{P}}_{S_{t}}\Bigg{(}$ $\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)),\epsilon(n_{t},\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu,\delta))\Biggr{)}$
$\geq\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}}1-\delta$
$\geq 1-\delta$
We use a union bound argument to add the $\forall k\in[1,T]$, leading to the term $\frac{\delta}{T}$.
As this theorem holds specifically for CoP2L, we specify the distributions $P_{\operatorname{\mathscr{P}}(n_{t})}$ and $P_{\mathscr{M}(\operatorname{\mathbf{j}})}$. Following the work of Marchand et al. (2003), which used sample compression bounds with three compression sets, with $\zeta(k)=\frac{6}{\pi}(k+1)^{-1}$, we choose $P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{i}})=\matrixquantity(n_{t}\\ \operatorname{|\mathbf{i}|})^{-1}\zeta(\operatorname{|\mathbf{i}|})$ and $P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})=\matrixquantity(n_{t}-\operatorname{|\mathbf{i}|}\\ |\operatorname{\mathbf{j}}|)^{-1}\zeta(|\operatorname{\mathbf{j}}|)$.
Finally, we split the message $\mu$ into two messages $\mu_{1}$ and $\mu_{2}$. The first message is defined using the alphabet $\Sigma_{1}=\{2,\ldots,T\}$. This message indicates the task number where a datapoint is sampled out of the buffer. Thus, the size of $\Sigma_{1}$ is $T-1$ and the probability of a symbol is $\frac{1}{T-1}$. For any vector $\operatorname{\mathbf{j}}$, we have a sequence of length $|\operatorname{\mathbf{j}}|$, and thus the probability of choosing each sequence is $\quantity(\frac{1}{T-1})^{|\operatorname{\mathbf{j}}|}$.
The second message $\mu_{2}$ is defined using the alphabet $\Sigma_{2}=\{1,n_{t}\}$. We choose to use $\zeta(k)=\frac{6}{\pi^{2}}(k+1)^{-2}$ as the probability distribution over each character. We know that for any $N\geq 0$, $\sum_{k=1}^{N}\zeta(k)\leq 1$. Thus, we have the probability of a sequence $\mu_{2}=\mu_{2}^{1}\ldots\nu_{2}^{T}$ is $\prod_{i=1}^{T}\zeta(\mu_{2}^{i})$.
We define $\mathscr{M}(\operatorname{\mathbf{j}},T)$ the set of messages such that the sequence of symbols from $\Sigma_{1}$ is of length $|\operatorname{\mathbf{j}}|$ and the sequence of symbols from $\Sigma_{2}$ is of length $T$. ∎
Appendix C Bound plots with different buffer sizes
In this section, we introduce the empirical bound estimates over tasks on MNIST, Fashion-MNIST and EMNIST datasets. In figure 5, we present results for different replay buffer sizes (1000, 2000, 3000, 4000, 5000).
Figure 5: CoP2L Bounds with respect to tasks on the MNIST, Fashion-MNIST and EMNIST datasets

### v3

@@
-Sample Compression for Continual Learning
+Sample Compression for Self-Certified Continual Learning
 Abstract
 Continual learning algorithms aim to learn from a sequence of tasks, making the training distribution non-stationary.
-The majority of existing continual learning approaches in the literature rely on heuristics and do not provide learning guarantees for the continual learning setup.
-In this paper, we present a new method called ‘Continual Pick-to-Learn’ (CoP2L), which is able to retain the most representative samples for each task in an efficient way.
-The algorithm is adapted from the Pick-to-Learn algorithm, rooted in the sample compression theory.
-This allows us to provide high-confidence upper bounds on the generalization loss of the learned predictors, numerically computable after every update of the learned model.
-We also empirically show on several standard continual learning benchmarks that our algorithm is able to outperform standard experience replay, significantly mitigating catastrophic forgetting.
-Machine Learning, ICML
+The majority of existing continual learning approaches in the literature rely on heuristics and do not provide learning guarantees.
+In this paper, we present a new method called Continual Pick-to-Learn (CoP2L), which is able to retain the most representative samples for each task in an efficient way.
+CoP2L combines the Pick-to-Learn algorithm (rooted in the sample compression theory) and the experience replay continual learning scheme.
+This allows us to provide non-vacuous upper bounds on the generalization loss of the learned predictors, numerically computable after each task.
+We empirically evaluate our approach on several standard continual learning benchmarks across Class-Incremental, Task-Incremental, and Domain-Incremental settings.
+Our results show that CoP2L is highly competitive across all setups, often outperforming existing baselines, and significantly mitigating catastrophic forgetting compared to vanilla experience replay in the Class-Incremental setting.
+It is possible to leverage the bounds provided by CoP2L in practical scenarios to certify the predictor reliability on previously learned tasks, in order to improve the trustworthiness of the continual learning algorithm.
 1 Introduction
 A common assumption in traditional machine learning is that the underlying data distribution does not evolve with time.
-In Continual Learning (De Lange et al., 2021; Wang et al., 2024), the goal is to develop machine learning algorithms that are able to learn under settings where this assumption is replaced by a set-up where the model is trained on an evolving training data distribution, in such a way that samples are revealed one task at a time.
+In Continual Learning de2021continual ; wang2023comprehensive , the goal is to develop machine learning algorithms that are able to learn under settings where this assumption is replaced by a set-up where the model is trained on an evolving training data distribution, in such a way that samples are revealed one task at a time.
 However, when neural networks are trained on evolving data distributions, networks tend to forget the training from earlier tasks.
-This fact is known as the catastrophic forgetting (McCloskey & Cohen, 1989; French, 1999; Goodfellow et al., 2014).
-In order to cope with forgetting, various methodologies have been developed, such as the regularization based approaches (e.g., Kirkpatrick et al., 2017), architectural approaches (e.g., Rusu et al., 2016), or rehearsal based approaches (e.g., Rolnick et al., 2019).
-It has been shown on various applications that rehearsal-based approaches typically perform very competitively compared to the other methods in terms of performance and forgetting (Libera et al., 2023; Mai et al., 2020; Chaudhry et al., 2019; Aljundi et al., 2019).
+This fact is known as the catastrophic forgetting McCloskey1989CatastrophicII ; forgetting_french ; goodfellow2013empirical .
+In order to cope with forgetting, various methodologies have been developed, such as regularization based approaches ewc_kirkpatrick ; zenke2017continual ; cl_li ; mas , architectural approaches pnn_rusu ; mallya2018piggyback ; expert_aljundi ; packnet_mallya ; l2p_wang , or rehearsal based approaches rolnick2019experience ; chaudhry2019efficient ; replay_shin ; icarl_rebuffi ; der .
+It has been shown on various applications that rehearsal-based approaches typically perform very competitively compared to the other methods in terms of performance and forgetting DellaLibera2023CLMASRAC ; Mai2020BatchlevelER ; chaudhry2019efficient ; aljundi2019onlinecontinuallearningmaximally .
 However, the selection of the replay buffer typically remains a heuristic, and precise upper bounds for the generalization error are not available.
 Our proposed continual learning scheme builds upon the Pick-to-Learn algorithm (P2L) of Paccagnan et al.
-(2024).
-P2L is a meta algorithm that selects a compression set, which amounts to express the dataset in terms of ‘core’ examples.
-This meta-algorithm was developed specifically with sample compression theory (Floyd & Warmuth, 1995; Campi & Garatti, 2023; Laviolette et al., 2005) in mind.
+paccagnan_pick_learn_2023 .
+P2L is a meta-algorithm that selects a compression set, which amounts to expressing the dataset in terms of ‘core’ examples.
+This meta-algorithm was developed specifically with sample compression theory in mind floyd1995sample ; campi2023compression .
 By finding a small subset of the data such that a predictor learned on this data achieves low error on the whole training set, Pick-to-Learn enables us to compute tight upper bounds for the generalization error of the learned predictor.
-In this paper, we propose ‘Continual Pick-to-Learn‘ (CoP2L), an algorithm that leverages the sample compression theory to intelligently select the training data from earlier tasks to mitigate forgetting.
+This ability of a learning algorithm to simultaneously output a predictor and its risk certificate has been coined as self-certified learning in the literature freund98selfbounding ; perezortiz2021tighter ; marks2025pick2learn-gp .
+This certification is arguably an asset to increase the trustworthiness towards machine learning systems.
+In this paper, we propose Continual Pick-to-Learn (CoP2L), an algorithm that leverages sample compression theory to intelligently select the training data from earlier tasks to mitigate forgetting.
 We derive high-confidence upper bounds on the generalization error for each task simultaneously, estimated directly from the training set.
-We also empirically show that CoP2L mitigates the forgetting significantly, and outperforms the standard replay baseline significantly in the small replay buffer size settings.
-Our contribution overall can be summarized as follows :
-  • We propose the algorithm Continual Pick-to-Learn (CoP2L), that integrates the sample compression theory within the continual learning setup.
+We also empirically show that CoP2L mitigates the forgetting significantly and outperforms the standard replay baseline significantly in the small replay buffer size settings.
+Our contributions can be summarized as follows:
+(1) We propose the algorithm Continual Pick-to-Learn (CoP2L), that integrates the sample compression theory within the continual learning setup.
 To the best of our knowledge, we are the first to integrate the theoretical results from sample compression to continual learning.
-  • CoP2L is able to provide high confidence non-trivial upper bounds for the generalization error by means of the sample compression theory.
+(2) CoP2L is able to provide high confidence non-trivial upper bounds for the generalization error by means of the sample compression theory.
 We experimentally showcase that the bounds are numerically computable, have non-trivial values and follow the general error trends observed on the test set.
-Therefore, they can actually be used as risk certificates on the model behavior on the learned tasks.
-  • We experimentally show that CoP2L significantly mitigates forgetting, while also being able to obtain better performance compared to vanilla replay, in the settings where the replay buffer is small.
+Therefore, they can actually be used as risk certificates on the model behavior on the learned tasks, improving the trustworthiness of the continually learnt model.
+(3) We experimentally show with a wide range of experiments that CoP2L significantly mitigates forgetting, while also being able to obtain better or comparable performance compared with several strong continual learning baselines in most cases.
 1.1 Related Work
-Practical Continual learning
+Continual learning approaches.
 In the continual learning literature, various practical approaches have been developed to mitigate forgetting.
-Broadly, the approaches could be divided into three categories which include regularization approaches (Kirkpatrick et al., 2017; Zenke et al., 2017; Li & Hoiem, 2018; Aljundi et al., 2018), architecture-based approaches (Rusu et al., 2016; Mallya et al., 2018; Aljundi et al., 2017; Mallya & Lazebnik, 2018; Wang et al., 2022) and rehearsal based approaches (Rolnick et al., 2019; Chaudhry et al., 2019; Shin et al., 2017; Rebuffi et al., 2017; Buzzega et al., 2020).
-In this paper, we use experience replay as a baseline as it is shown in the literature to be a very strong continual learning baseline to mitigate forgetting (Libera et al., 2023).
-Theoretical Continual Learning
-There have been several works in theoretical continual learning in the recent years.
-For example, Bennani & Sugiyama (2020); Doan et al.
-(2021) study the generalization error obtained with Neural Tangent Kernel (NTK) models when the model is trained with Orthogonal Gradient Descent (Farajtabar et al., 2020).
-Yin et al.
-(2020) provides generalization bounds for regularization-based continual learning methods.
-Asanuma et al.
-(2021) studies the impact of task similarity in the student-teacher setup.
-Li et al.
-(2022) provides a sample complexity bound in the context of continual representation learning for an adapted PackNet (Mallya & Lazebnik, 2018).
-In Chen et al.
-(2022), the authors provide a theoretical bound on memory using PAC learning theory.
-These works however do not provide bounds that are computable in practice.
-A line of work also investigates the generalization of continual learning models when linear models are employed.
-In Goldfarb & Hand (2023), the authors investigate the effect of overparametrized linear models.
-In Evron et al.
-(2022); Lin et al.
-(2023), the authors analyze the generalization behavior of linear regression models, even though Lin et al.
-(2023) conjectures that their results could be extended to neural networks as well.
-Different from all the listed works above, in this paper, we provide a practical non-trivial generalization bound that is computed on the training set, and predicts the test error for each task separately.
-Furthermore, also unlike the approaches above, our bound is applicable to any neural network architecture and does not have limitations on the continual learning setup that is employed.
-Sample Compression
-The sample compress theory has been shown effective for providing tight generalization bounds (Laviolette et al., 2005; Marchand et al., 2003; Marchand & Shawe-Taylor, 2002; Marchand & Sokolova, 2005; Shah, 2007).
+Broadly, the approaches could be divided into three categories, which include regularization approaches ewc_kirkpatrick ; zenke2017continual ; cl_li ; mas , architecture-based approaches pnn_rusu ; mallya2018piggyback ; expert_aljundi ; packnet_mallya ; l2p_wang and rehearsal based approaches rolnick2019experience ; chaudhry2019efficient ; replay_shin ; icarl_rebuffi ; der .
+Continual learning generalization bounds.
+In recent years, generalization bounds were proposed to study the asymptotic behavior of various continual learning schemes, for example using linear models Evron2022HowCC ; Goldfarb2022AnalysisOC ; pmlr-v202-lin23f , Neural Tangent Kernel (NTK) models trained with Orthogonal Gradient Descent AbbanaBennani2020GeneralisationGF ; Doan2020ATA ; farajtabar2019orthogonalgradientdescentcontinual , regularization-based methods yin2021optimizationgeneralizationregularizationbasedcontinual , or PackNet packnet_mallya inspired continual representation learning Li2022ProvableAE .
+Learning algorithms with convergence guarantees are were also proposed under data assumptions, such that all target tasks lie in a same space obtainable from a (linear or non-linear) combination of some input features cao2022lifelong .
+Different from all the listed works above, in this paper, we provide a generalization bound that is computed from the training set.
+Risk certificates on the test error are provided simultaneously for each task.
+Furthermore, also unlike the approaches above, our non-asymptotic bound is applicable to any neural network architecture, under minimal data assumptions, and does not have limitations on the continual learning setup that is employed.
+Sample compression.
+The sample compression theory has been shown effective for providing tight generalization bounds shah_margin-sparsity_2005 ; marchand2003set ; marchand2002set ; marchand2005learning ; shah_sample_2007 .
 However, most sample compress approaches are limited to low-complexity models.
-A notable exception is the method Pick-to-Learn (P2L), which successfully provides guarantees for deep neural networks (Campi & Garatti, 2023; Paccagnan et al., 2024).
+A notable exception is the method Pick-to-Learn (P2L), which successfully provides guarantees for deep neural networks campi2023compression ; bazinet2024 and Gaussian Processes marks2025pick2learn-gp .
 In this paper, we adapt the key ideas introduced in the sample compression theory to the continual learning case, which enables us to provide an upper bound on the test error from training samples.
 2 Background
-2.1 Continual Learning
-Notation.
+Continual learning.
 Let us consider a series of task distributions ${\mathcal{D}}_{1},{\mathcal{D}}_{2},\ldots,{\mathcal{D}}_{T}$ on an input-output space ${\mathcal{X}}\times{\mathcal{Y}}$, where the amount of tasks $T\in{\mathbb{N}}^{*}$ can be (countably) infinite.
@@
 Nevertheless, at task $t$, we would like the system to perform well on all tasks, including the previous ones $11,\dots,t-$ 1.
-However, simply updating the parameters $\theta_{t-1}$ learned from a previous tasks to optimize the loss ^$\widehat{\mathcal{L}}_{S_{t}}(\theta_{t})$ on a current task $t$ would lead to catastrophic forgetting (French, 1999).
+However, simply updating the parameters $\theta_{t-1}$ learned from a previous tasks to optimize the loss ^$\widehat{\mathcal{L}}_{S_{t}}(\theta_{t})$ on a current task $t$ would lead to catastrophic forgetting forgetting_french .
 Therefore, as indicated in Section 1.1, numerous empirical methods have been developed to mitigate forgetting.
@@
 where the second term is the loss on the replay buffer ${\mathcal{B}}$.
-In our experiments, we use this method as the baseline as it’s shown to be a simple and very effective continual learning strategy (Libera et al., 2023; Mai et al., 2020).
-2.2 Sample Compression Theory
-Given a dataset $S=\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{n}$, a family of learnable parameters $\Theta$ and a learning algorithm $A$ such that $A(S)\in\Theta$, sample compression theory provides generalization bounds for any predictor $A(S)$ on the condition that the predictor can be provably represented as a function of a small subset of the dataset, called the compression set, and an additional source of information, called the message.
-If this is the case, we call $A(S)$ a sample-compressed predictor.
+In our experiments, we use this method as a baseline as it’s shown to be a simple and very effective continual learning strategy DellaLibera2023CLMASRAC ; Mai2020BatchlevelER .
+Sample compression theory.
+Given a training dataset $S=\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{n}\in(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{n}$, a family of learnable parameters $\Theta$ and a learning algorithm $A$ such that $A(S)\in\Theta$, sample compression theory provides generalization bounds for any predictor $f_{\theta}$, with $\theta=A(S)$, on the condition that $f_{\theta}$ can be provably represented as a function of a small subset of the training dataset $S$, called the compression set, and an additional source of information, called the message.
+If this is the case, we call $f_{\theta}$ a sample-compressed predictor.
 The compression set is denoted $S^{\operatorname{\mathbf{i}}}$ and is parameterized by a vector of indices $\operatorname{\mathbf{i}}\subset\{1,\ldots,n\}$.
-Each vector $\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)$, with $\operatorname{\mathscr{P}}(n)$ the set of all the possible subsets of $\{1,\ldots,n\}$, is ordered such that we have
+Each vector $\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)$, with $\operatorname{\mathscr{P}}(n)$ the set of all the possible subsets of $\{1,\ldots,n\}$, is ordered such that
 $\operatorname{\mathbf{i}}=(i_{1},i_{2},\ldots,i_{\operatorname{|\mathbf{i}|}})\,,\mbox{ with }1\leq i_{1}<\ldots<i_{\operatorname{|\mathbf{i}|}}\leq n\,.$ (4)
-The compression set is defined such that
-$S^{\operatorname{\mathbf{i}}}=S^{(i_{1},\ldots,i_{\operatorname{|\mathbf{i}|}})}\!=\{(\boldsymbol{x}_{i_{1}},y_{i_{1}}),\ldots,(\boldsymbol{x}_{i_{\operatorname{|\mathbf{i}|}}},y_{i_{\operatorname{|\mathbf{i}|}}})\}\!\subseteq S.$ (5)
-We define the complement vector $\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}$ such that $\operatorname{\mathbf{i}}\cap\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}=\emptyset$ and $\operatorname{\mathbf{i}}\cup\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}=\{1,\ldots,n\}$.
-Thus, we have the complement set $S_{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}=S\setminus S_{\operatorname{\mathbf{i}}}$ and $|\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}|=n-\operatorname{|\mathbf{i}|}$.
+The compression set is defined such that $S^{\operatorname{\mathbf{i}}}=S^{(i_{1},\ldots,i_{\operatorname{|\mathbf{i}|}})}\!=\{(\boldsymbol{x}_{i_{1}},y_{i_{1}}),\ldots,(\boldsymbol{x}_{i_{\operatorname{|\mathbf{i}|}}},y_{i_{\operatorname{|\mathbf{i}|}}})\}\!\subseteq S.$ We define the complement vector $\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}$ such that $\operatorname{\mathbf{i}}\cap\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}=\emptyset$ and $\operatorname{\mathbf{i}}\cup\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}=\{1,\ldots,n\}$.
+Thus, we have the complement set $S^{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}=S\setminus S^{\operatorname{\mathbf{i}}}$ and $|\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}|=n-\operatorname{|\mathbf{i}|}$.
 In addition to the compression set, a message $\mu$ is sometimes needed to compress the predictor.
@@
 For all $\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)$, we choose $\mathscr{M}(\operatorname{\mathbf{i}})$ a countable subset of $\Sigma^{*}$, which represents all the possible messages that can be chosen for $\operatorname{\mathbf{i}}$.
-Given a learning algorithm $A$ and a dataset $S$, to prove that the predictor $A(S)$ is a sample-compressed one, we need to define two functions: a compression function and a reconstruction function.
-The compression function $\mathscr{C}:\bigcup_{1\leq n\leq\infty}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{n}\to\bigcup_{m\leq n}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{m}\times\bigcup_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\mathscr{M}(\operatorname{\mathbf{i}})$ outputs a compression set and a message that are used to represent $A(S)$.
-Given a dataset $S$, we have $\mathscr{C}(S)=(S^{\operatorname{\mathbf{i}}},\mu)$.
-Then, the reconstruction function $\operatorname{\mathscr{R}}:\bigcup_{m\leq n}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{m}\times\bigcup_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\mathscr{M}(\operatorname{\mathbf{i}})\to\Theta$ is defined such that $A(S)=\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},\mu)$.
-Both the reconstruction function and the compression function must be deterministic and data-independent.
+Given learned parameters $\theta=A(S)$, to prove that the predictor $f_{\theta}$ is a sample-compressed one, we need to define two functions: a compression function and a reconstruction function.
+The compression function $\mathscr{C}:\boldsymbol{\cup}_{1\leq n\leq\infty}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{n}\to\boldsymbol{\cup}_{m\leq n}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{m}\times\boldsymbol{\cup}_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\mathscr{M}(\operatorname{\mathbf{i}})$ outputs a compression set and a message.
+Then, the reconstruction function $\operatorname{\mathscr{R}}:\boldsymbol{\cup}_{m\leq n}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{m}\times\boldsymbol{\cup}_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\mathscr{M}(\operatorname{\mathbf{i}})\to\Theta$ is defined such that $A(S)=\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},\mu)$.
 The forthcoming results rely on a probability distribution over the set of sample-compressed predictors $\overline{\Theta}\subseteq\Theta$.
 For any sample-compressed predictor $\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)$, this data-independent distribution is expressed as $P_{\overline{\Theta}}\quantity(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu))=P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{i}})}(\mu)$, with $P_{\operatorname{\mathscr{P}}(n)}$ a probability distribution over $\operatorname{\mathscr{P}}(n)$ and $P_{\mathscr{M}(\operatorname{\mathbf{i}})}$ a probability distribution over $\mathscr{M}(\operatorname{\mathbf{i}})$.
-Following the work of Marchand & Sokolova (2005), we consider $P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})=\smallmatrixquantity(n\\ \operatorname{|\mathbf{i}|})^{-1}\zeta(\operatorname{|\mathbf{i}|})$, with $\zeta(k){=}\tfrac{6}{\pi^{2}}(k+1)^{-2}$.
-Given a vector of indices $\operatorname{\mathbf{i}}$, we consider $P_{\mathscr{M}(\operatorname{\mathbf{i}})}$ to be a uniform distribution over the messages.
-We now present the sample compression bound of Bazinet et al.
-(2025).
-We use the shorthand notation ^$\operatorname{\widehat{\mathcal{L}}}_{S}^{\,{}_{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}}(\cdot)$ to denote the empirical loss on the training sample that don’t belong to the compression set: $S^{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}=S\setminus S^{\mathbf{i}}$; this is mandatory to obtain an unbiased estimate of the loss, as the reconstructed predictor $\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)$ relies on $S^{\mathbf{i}}$.
-Theorem 2.1.
+Following common practices marchand2005learning ; bazinet2024 , we consider hereunder $P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})={n\choose\operatorname{|\mathbf{i}|}}^{{}_{-1}}\zeta(\operatorname{|\mathbf{i}|})$, with $\zeta(k){=}\tfrac{6}{\pi^{{}_{2}}}(k+1)^{-2}$.
+Also, given a vector of indices $\operatorname{\mathbf{i}}$, we consider $P_{\mathscr{M}(\operatorname{\mathbf{i}})}$ to be a uniform distribution over the messages.
+We now present a recent sample compression generalization guarantee bazinet2024 on which we build our theoretical framework.
+We use the shorthand notation ^$\operatorname{\widehat{\mathcal{L}}}_{S}^{\,{}_{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}}(\cdot)$ to denote the empirical loss on the training sample that do not belong to the compression set: $S^{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}=S\setminus S^{\mathbf{i}}$; this is mandatory to obtain an unbiased estimate of the loss, as the reconstructed predictor $\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)$ relies on $S^{\mathbf{i}}$.
+Theorem 2.1 (bazinet2024 ).
 For any distribution $\operatorname{\mathcal{D}}$ over $\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}$, for any family of set of messages $\{\mathscr{M}(\operatorname{\mathbf{i}})|\operatorname{\mathbf{i}}{\in}\operatorname{\mathscr{P}}(n)\}$, for any deterministic reconstruction function $\operatorname{\mathscr{R}}$, for any loss $\ell:\Theta\times\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}{\to}[0,1]$ and for any $\delta\in(0,1]$, with probability at least $1-\delta$ over the draw of $S\sim\operatorname{\mathcal{D}}^{n}$, we have
 $\forall\mathbf{i}$ $\in\operatorname{\mathscr{P}}(n),\,\mu\in\mathscr{M}(\operatorname{\mathbf{i}}):$
-$\mathrm{kl}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,{}_{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}}\quantity(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)),\mathcal{L}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)))\leq\epsilon(n,\operatorname{\mathbf{i}},\mu,\delta)$
-with $\mathrm{kl}(q,p)=q\ln\frac{q}{p}+(1-q)\ln\frac{1-q}{1-p}$ the binary Kullback-Leibler divergence and
-$\epsilon(n,\operatorname{\mathbf{i}},\mu,\delta)=\frac{1}{n-\operatorname{|\mathbf{i}|}}\quantity[\log\!\matrixquantity(n\\ \operatorname{|\mathbf{i}|})\!\!+\!\log\!\quantity(\frac{2\sqrt{n-\operatorname{|\mathbf{i}|}}}{\zeta(\operatorname{|\mathbf{i}|})P_{\mathscr{M}(\operatorname{\mathbf{i}})}(\mu)\delta})].$
-Given the parameters $\theta$ of a reconstructed predictor, i.e., $\theta{=}\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)$, Theorem 2.1 states that the $\mathrm{kl}$-discrepancy between the true risk ^$\operatorname{\widehat{\mathcal{L}}}_{S}^{\,{}_{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}}(\theta)$ and the empirical risk $\mathcal{L}_{\operatorname{\mathcal{D}}}(\theta)$ is upper bounded by the complexity term $\epsilon(n,\operatorname{\mathbf{i}},\mu,\delta)$.
-From this result, one can compute an upper bound on the true risk $\mathcal{L}_{\operatorname{\mathcal{D}}}(\theta)$ by inverting the Kullback-Leibler divergence.
-Indeed, Theorem 2.1 can be rewritten as
-$\mathcal{L}_{\operatorname{\mathcal{D}}}\quantity(\theta)\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,{}_{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}}\quantity(\theta),\epsilon(n,\operatorname{\mathbf{i}},\mu,\delta))$ (6)
-with
-$\mathrm{kl}^{-1}(q,\varepsilon)\ =\ \operatorname*{arg\,sup}_{0\leq p\leq 1}\{\mathrm{kl}(q,p)\leq\varepsilon\}.$ (7)
+$\mathcal{L}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu))\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,{}_{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}}\quantity(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)),\frac{1}{n-\operatorname{|\mathbf{i}|}}\quantity[\log\!\matrixquantity(n\\ \operatorname{|\mathbf{i}|})\!\!+\!\log\!\quantity(\frac{2\sqrt{n-\operatorname{|\mathbf{i}|}}}{\zeta(\operatorname{|\mathbf{i}|})P_{\mathscr{M}(\operatorname{\mathbf{i}})}(\mu)\delta})])\,,$
+with $\mathrm{kl}(q,p)=q\ln\frac{q}{p}+(1-q)\ln\frac{1-q}{1-p}$ the binary Kullback-Leibler divergence and its inverse
+$\mathrm{kl}^{-1}(q,\varepsilon)=\mbox{\rm argsup}_{p\in[0,1]}\left\{\mathrm{kl}(q,p)\leq\varepsilon\right\}.$ (5)
 On the one hand, given a fixed compression set size $\operatorname{|\mathbf{i}|}$, the bound decreases when the training set size $n$ increase.
 On the other hand, given a fixed $n$, the bound increases with $\operatorname{|\mathbf{i}|}$.
-2.3 Pick-to-Learn
+Pick-to-Learn.
 To obtain sample compression guarantees for a class of predictors, it is necessary to prove that the learned predictor only depends on a small subset of the data and (optionally) a message.
-Some predictors, such as the SVM (Boser et al., 1992) and the perceptron (Rosenblatt, 1958; Moran et al., 2020), have straightforward compression scheme, as they only depend on a subset of the data after training.
-Some algorithms, such as the set covering machine (SCM) (Marchand & Shawe-Taylor, 2002; Marchand et al., 2003; Marchand & Sokolova, 2005) and decision trees (Shah, 2007), necessitate hand-crafted compression schemes involving a message.
+Some predictors, such as the SVM boser_svm and the perceptron rosenblatt1958perceptron ; moran2020perceptron , have straightforward compression scheme, as they only depend on a subset of the data after training.
+Some algorithms, such as the set covering machine (SCM) marchand2002set ; marchand2003set ; marchand2005learning and decision trees shah_sample_2007 , necessitate hand-crafted compression schemes involving a message.
 Up until recently, there was no sample compression scheme for deep neural networks.
-The meta-algorithm Pick-To-Learn (P2L) was proposed by Paccagnan et al.
-(2024) as a compression scheme for any class of predictors, with a specific focus to deep neural networks.
-The meta-algorithm modifies the training loop of the predictor, by iteratively choosing datapoints over which the model is updated.
-Starting with initial parameters $\theta_{\rm init}$, P2L evaluates the predictor on the whole training dataset, adds the datapoints with the highest losses to the compression set and then updates the predictor using the compression set.
-The meta-algorithm stops once the losses of training examples not in the compression set are lower than a given stopping criteria ($-\ln(0.5)$ for the cross-entropy loss, which is equivalent to achieving zero errors on the complement set $S^{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}$).
-We provide the pseudo-code of Pick-To-Learn in Algorithm 1.
-Algorithm 1 Original Pick-To-Learn (P2L)
+Algorithm 1 Original Pick-To-Learn (P2L) paccagnan_pick_learn_2023
 0:  $S=\{({\mathbf{x}}_{i},y_{i})\}_{i=1}^{n}$Training set}
-0:  $\theta_{\rm init}$Initialization parameters of a given model $f_{\theta}$}
+0:  $\theta_{\rm init}$Initial parameters of the model $f_{\theta}$}
 0:  $\gamma$Stopping criteria}
@@
 8:  end while
-9:  return $\theta_{k},C_{k}$.
-10:  {Learned parameters and compression set}
+9:  return $\theta_{k}$Final parameters}
+The meta-algorithm Pick-To-Learn (P2L) was proposed by Paccagnan and Campi paccagnan_pick_learn_2023 as a compression scheme for any class of predictors, with a specific focus to deep neural networks.
+The meta-algorithm acts as an aditional training loop, by iteratively selecting datapoints over which the model is updated.
+Starting with initial parameters $\theta_{\rm init}$, P2L evaluates the predictor on the whole training dataset, adds the datapoints with the highest losses to the compression set and then updates the predictor using the compression set.
+The meta-algorithm stops once the losses of training examples not in the compression set are lower than a given stopping criteria ($-\ln(0.5)$ for the cross-entropy loss, which is equivalent to achieving zero errors on the complement set $S^{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}$).
+We provide the pseudo-code of the original P2L in Algorithm 1.
+Recent works introduce improvements to the original scheme, notably early stopping strategies marks2025pick2learn-gp ; bazinet2024 and extension to the Bayesian setting marks2025pick2learn-gp .
 3 Sample Compression for Continual Learning
-3.1 A New Training Scheme
 A striking realization coming from the P2L algorithm is that only a small fraction of the training set – the compression set – needs to be provided to the learner in order to achieve good generalization.
@@
 This observation motivates our strategy for managing the replay buffer of our new algorithm Continual Pick-To-Learn (CoP2L): we subsample the complement set to create the replay buffer, in order to maintain a low risk on these examples while learning subsequent tasks.
-When needed, CoP2L can choose to add a datapoint from the buffer, thus mitigating the forgetting by adding a well-chosen datapoint.
+When needed, CoP2L can choose to add a datapoint from the buffer to the compression set, thus mitigating the forgetting by focusing on a well-chosen datapoint.
 Algorithm 2 Modified Pick-To-Learn (mP2L)
-0:  $\theta_{\rm init}$Initialization parameters of a given model $f_{\theta}$}
+0:  $\theta_{\rm init}$Initial parameters of the model $f_{\theta}$}
 0:  $S=\{({\mathbf{x}}_{i},y_{i},w_{i})\}_{i=1}^{n}$Training set (with weights)}
@@
 10:  if $k<K^{\star}$ then
-11:     $k\leftarrow\operatorname*{\mathrm{argmin}}_{0\leq k^{\prime}\leq k}\Psi(S,\theta_{k^{\prime}},C_{k^{\prime}}\cap S)$
+11:    $k\leftarrow\operatorname*{\mathrm{argmin}}_{0\leq k^{\prime}\leq k}\Psi(S,\theta_{k^{\prime}},C_{k^{\prime}}\cap S)$Eq. (6)}
 12:  end if
-13:  return $\theta_{k},C_{k}$.
-14:  {Learned parameters and compression set}
+13:  return $\theta_{k},C_{k}$. {Final parameters and compression set}
 Algorithm 3 Continual Pick-To-Learn (CoP2L)
+0:  $\theta_{0}${Initial parameters of the model}
 0:  $S_{1},S_{2},\ldots,S_{T}$Training sets}
-0:  $\theta_{0}$Initialization parameters of a given model}
-0:  $\gamma$P2L’s stopping criteria}
+0:  $\gamma$mP2L’s stopping criteria}
 0:  $m$Buffer’s max sampling size}
@@
 5:    $\theta_{t},C^{\star}\leftarrow\mbox{mP2L}(\theta_{t-1},\hat{S_{t}},B^{\star},\gamma,\infty)$
-6:     $B_{i}\leftarrow\mbox{sample}(B_{i},\lfloor\frac{m}{t}\rfloor)\quad\forall i=1,\ldots,t-1$
+6:    $B_{i}\leftarrow\mbox{sample}(B_{i},\lfloor\frac{m}{t}\rfloor)\ \forall i{=}1,\ldots,t{-}1$
 7:     $B_{t}\leftarrow\mbox{sample}(\hat{S}_{t}\setminus C^{\star},\lfloor\frac{m}{t}\rfloor)$
@@
 9:  end for
-10:  return $\theta_{T}$. {Learned parameters}
+10:  return $\theta_{T}$. {Final parameters}
 Each new task is learned by a modified version of the P2L meta-algorithm (Algorithm 2, entitled mP2L), as explained further down.
-That is, for each new task, the proposed continual learning algorithm CoP2L calls mP2L and updates the buffer using the datapoints that were not chosen in the compression set.
-Similarly to the replay buffer method, at the end of each task $t$, the buffer contains $\lfloor\tfrac{m}{t}\rfloor$ datapoints of each previous task, with $m$ the maximum size of the buffer.
-We present CoP2L in Algorithm 3.
+That is, for each new task, the proposed continual learning algorithm (Algorithm 3, entitled CoP2L) calls mP2L and updates the buffer using the datapoints that were not chosen in the compression set.
+Similarly to the replay buffer method, at the end of each task $t$, the buffer contains $\lfloor\tfrac{m}{t}\rfloor$ datapoints of each previous task, with $m$ the maximum buffer size.
 We modified Pick-To-Learn in two significant ways.
-First, we introduced weights to the loss functions to tackle the imbalance problem between the current task and the previous tasks.
+(i) First, we introduced weights to the loss functions to tackle the imbalance problem between the current task and the previous tasks.
 When working with a replay buffer, the class imbalance problem is taken care of by training on a number of datapoints from previous tasks.
@@
 This ensures that the model has achieved perfect accuracy on the current task and all the datapoints from the previous tasks that are found in the buffer.
-The second modification leading to mP2L is the choice of the returned predictor.
+(ii) The second modification is the choice of the returned predictor by mP2L.
 The original P2L algorithm trains the model until it first achieves zero errors on the complement set.
-In presence of a noisy dataset, this might lead to an overfitting behavior.
-Instead, to perform early stopping, mP2L relies on the trade-off encoded in Theorem 2.1 between the accuracy on the complement set and the complexity of the model.
+In presence of a noisy dataset, this might lead to an overfitting behavior, P2L adding most training samples into its compression set, which in turn gives a large generalization bound value.
+Instead, we perform early stopping based on the bound value (as proposed by bazinet2024 ).
+That is, mP2L relies on the trade-off encoded in Theorem 2.1 between the accuracy on the complement set and the complexity of the model.
 More precisely, it returns the model’s parameters $\theta$ that minimizes
@@
 The latter is obtained from the procedure sample$(S,m)$ (see Algorithm 3), which represents the random sampling of $m$ instances of $S$ without replacement.
-From a theoretical standpoint, in contrast to the original P2L, both mP2L and CoP2L cannot be reconstructed straightforwardly without a message for two reasons: (1) the bound function $\Psi(S,\theta,C)$ cannot be computed on the whole dataset, when only the compression set is given as input; (2) the sampling procedure sample$(S,m)$ cannot be reproduced in the reconstruction step.
-In the following section, we define the compression and reconstruction functions of CoP2L, which will give rise to sample compression bounds as presented in Theorem 3.1.
-3.2 Compression and Reconstruction Scheme
-To obtain generalization bounds for CoP2L, we need to prove that it can be compressed.
-To do so, we provide the compression and reconstruction functions of CoP2L.
-The compression function is presented in Algorithm 4 (provided in supplementary material) and the reconstruction function is presented in Algorithm 5.
-Compression function.
-In line with the prevalent literature, we presented the sample compression framework (Section 2.2) based on a compression function providing a compression set and optionally a message.
+Compression and reconstruction scheme.
+From a theoretical standpoint, in contrast to the original P2L, both mP2L and CoP2L cannot be reconstructed straightforwardly without a message for two reasons: (1) when only the compression set is given as input, the bound function $\Psi(S,\theta,C)$ used as stopping criteria in mP2L cannot be computed on the whole dataset; (2) the sampling procedure sample$(S,m)$ in CoP2L cannot be reproduced in the reconstruction step.
+In Appendix A, we provide the compression and reconstruction functions of CoP2L, which give rise to sample compression bounds as presented in Theorem 3.1.
+Noteworthy, after learning on $T$ tasks, the compression function provides two sample compression sets $S^{\operatorname{\mathbf{i}}}$ and $S^{\operatorname{\mathbf{j}}}$, along with a message pair $(\mu_{1},\mu_{2})$ chosen among the set of all possible messages, denoted as $\mathscr{M}(\operatorname{\mathbf{j}},T)=\{2,\ldots,T\}^{|\operatorname{\mathbf{j}}|}\times\left[\boldsymbol{\times}_{t=1}^{T}\{1,\ldots,n_{t}+|{\mathcal{B}}|\}\right].$ As explained further in appendix, the use of two compression sets, along with the message $\mu_{1}=[\mu_{1}^{i}]_{i=1}^{T}$, serves to the proper treatement of the buffer $B^{\star}$, while the message $\mu_{2}=[\mu_{2}^{i}]_{i=1}^{T}$ contains the number of iterations $K^{\star}$ to perform for each call to mP2L.
+Generalization bound.
+Given the compression and reconstruction function detailed in Apprendix A, Theorem 2.1 can be used to obtain generalization bounds on the last learned task, which is done by mP2L to compute the bound $\Psi$.
+However, this bound holds for only one distribution of data.
+Therefore, it cannot be applied to bound the risks of the learned predictor on previously tasks seen by CoP2L.
+The next Theorem 3.1 holds for any previously learned tasks.
+For a dataset $S_{t}\sim\operatorname{\mathcal{D}}_{t}$, we denote the loss on the complement set of the joint set $(S_{t}^{\operatorname{\mathbf{i}}}\cup S_{t}^{\operatorname{\mathbf{j}}})^{\rm\bf c}$ as ^$\widehat{{\mathcal{L}}}_{S_{t}}^{\ \operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\!(\theta)$.
+Moreover, we denote the reconstruction function of CoP2L as $\operatorname{\mathscr{R}}_{1:T}(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)=\operatorname{\mathscr{R}}\big{(}S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu\mid S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}\big{)}.$ This formulation is important for the following theorem, as the reconstruction function for task $t$ is conditioned on all datasets $S_{1}$ to $S_{T}$, with the exception of $S_{t}$.
+Theorem 3.1.
+For any set of independent distributions $\left\{{\mathcal{D}}_{t}\right\}_{t=1}^{T}$ over ${\mathcal{X}}\times{\mathcal{Y}}$, with $\operatorname{\mathscr{R}}_{1:T}$ the reconstruction function of CoP2L (Algorithm 5), with the family of set of messages $\{\mathscr{M}(\operatorname{\mathbf{j}},T)|\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)\}$ defined for CoP2L, for any loss $\ell:\Theta\times{\mathcal{X}}\times{\mathcal{Y}}\to[0,1]$, and for any $\delta\in(0,1]$, with probability at least $1-\delta$ over the draw of $S_{t}\sim{\mathcal{D}}_{t},t=1,\ldots,T$, we have
+$\forall t\in\{1,\ldots,T\},{\mathbf{i}}\in\operatorname{\mathscr{P}}(n_{t}),\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n_{t}),\mu=(\mu_{1},[\mu_{2}^{i}]_{i=1}^{T})\in\mathscr{M}(\operatorname{\mathbf{j}},T):$
+$\mathrm{kl}\quantity(\widehat{{\mathcal{L}}}_{S_{t}}^{\ \operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\!\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)),\ {\mathcal{L}}_{\operatorname{\mathcal{D}}_{t}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)))$
+$\leq\tfrac{1}{n_{t}-|{\mathbf{i}}|-|\mathbf{j}|}\big{[}\log{\textstyle{n_{t}\choose|{\mathbf{i}}|}}+\log{\textstyle{n_{t}-\operatorname{|\mathbf{i}|}\choose|\mathbf{j}|}}+\textstyle\sum_{i=1}^{T}\log\tfrac{1}{\zeta(\mu_{2}^{i})}+|\mathbf{j}|\log(T-1)+\log\tfrac{T}{\zeta(|{\mathbf{i}}|)\zeta(|\operatorname{\mathbf{j}}|)\delta}\big{]}\,.$
+The proof is given in Appendix C.
+It first relies on a tighter version of Theorem 2.1, where the $2\sqrt{n-\operatorname{|\mathbf{i}|}}$ term is removed and a second compression set is added.
+Then, this new result is adapted to the continual learning setting.
+For all tasks $1$ through $T$, this new theorem encodes a tradeoff similar to Theorem 2.1 between the predictor accuracy and its complexity.
+Here, the predictor complexity is evaluated using the size of the compression sets $\operatorname{\mathbf{i}}$ and $\operatorname{\mathbf{j}}$ and the probability of choosing a message.
+The probability of the messages $\mu_{1}$ and $\mu_{2}$ are functions of the size of $\operatorname{\mathbf{j}}$ and the number of tasks.
+Indeed, the probability of $\mu_{1}$ decays when $|\mathbf{j}|$ and $T$ grow larger.
+Moreover, the probability of $\mu_{2}$ depends on the size of the compression set outputted by mP2L at each task $T$.
+Thus, the larger the compression sets, the smaller the probability of both messages.
+In conclusion, the bound encodes the ability of a model to minimizes the loss on the dataset whilst still keeping $\operatorname{\mathbf{i}}$ and $\operatorname{\mathbf{j}}$ the small.
+4 Experiments
+Datasets and continual learning settings.
+We carry-out experiments on MNIST lecun2010mnist , Fashion-MNIST fashionmnist , EMNIST (characters) cohen2017emnistextensionmnisthandwritten , PermutedMNIST goodfellow2013empirical , Rotated MNIST, CIFAR10 cifar10cifar100 , CIFAR100 cifar10cifar100 and TinyImageNet tiny-imagenet .
+We report results that include Class-Incremental (CI), Task-Incremental (TI), and Domain-Incremental (DI) continual learning settings.
+The dataset stats regarding how many classes/data items are available per task is given in Appendix H.
+Baselines.
+To assess the effectiveness of our proposed method CoP2L, we benchmark against several strong and widely adopted baselines across Class-Incremental, Task-Incremental, and Domain-Incremental continual learning settings.
+For each continual learning set-up we include baselines that yield the best performance in each category according to the Avalanche benchmark lomonaco2021avalanche .
+For the Class-Incremental setting, we use the baselines Replay rolnick2019experience , iCaRL rebuffi2017icarl , and GDumb prabhu2020gdumb .
+Replay in particular is considered a strong baseline DellaLibera2023CLMASRAC due to its simplicity and effectiveness.
+In the Task-Incremental setting, we compare against Replay rolnick2019experience , LaMAML gupta2020look and Learning without Forgetting (LwF) li2017learning .
+These methods are well suited for TI scenarios where task identity is known at test time.
+In the Domain-Incremental setting, we evaluate against Replay rolnick2019experience , Less Forgetful Learning (LFL) jung2018less , Elastic Weight Consolidation (EWC) kirkpatrick2017overcoming , and Synaptic Intelligence (SI) zenke2017continual .
+Implementation.
+All our experiments are conducted using the Avalanche continual learning framework avalanchelib , which provides a unified interface for datasets, benchmarks, and baseline implementations.
+We also use the official ContinualAI benchmark repository [^1] https://github.com/ContinualAI/continual-learning-baselines for consistent implementation and fair comparisons.
+The code for the experimental results is available in the supplemental material.
+Model architecture and training details.
+We use the SimpleMLP architecture and customized SimpleCNN architectures (SimpleMLP and SimpleCNN from the Avalanche toolkit) for the MNIST, FMNIST, and EMNIST experiments.
+All models are trained using the SGD optimizer.
+The details of the models architectures and the hyperparameters used can be found in Appendix G.
+For experiments on more complex datasets such as CIFAR10, CIFAR100, and TinyImageNet, we use the pretrained feature extractor Dino caron2021emerging using ViT dosovitskiy2020image and ResNet50 he2016deep backbones, followed by a trainable MLP head.
+This choice is motivated by the insight that CoP2L can further benefit from stronger feature representations, especially in settings with limited data per class or high dataset complexity.
+By leveraging Dino’s obtained embeddings, we enhance the quality of the input representations, which we empirically observe to yield good performance.
+Results with standard deviations across seeds and total training times are reported in Appendix F.
+We used a single NVIDIA 4090 GPU for our experiments.
+Accuracy and Forgetting.  We denote the accuracy obtained on a task $t$, after training the model for $T$ tasks with $\text{Accuracy}_{T,t}$. The accuracy we report is defined as,
+. $\textit{Average Accuracy}_{T}=\frac{1}{T}\sum_{t=1}^{T}\text{Accuracy}_{T,t}$, where $t$ is for the task index. .
+We also report the average forgetting at task $T$ which is defined as
+$\textit{Average Forgetting}_{T}=\frac{1}{(T-1)}\sum_{t=1}^{T-1}[\text{Accuracy}_{t,t}-\text{Accuracy}_{T,t}],$ which is calculated when $T>1$.
+4.1 Main Results
+In Table 1, we present the average accuracy and average forgetting values defined in the section above, obtained on MNIST, FMNIST and EMNIST dataset, both with an MLP and a CNN architecture, where the networks are trained from scratch.
+In Appendix F, we also provide the standard deviations as well.
+We observe that with the MLP architecture CoP2L is able to outperform the baselines.
+Furthermore, we also observe that CoP2L is able to obtain competitive results when employed with a CNN architecture as well.
+In Appendix E, we also provide an analysis of the performance of CoP2L for varying buffer sizes.
+In Table 2, we extended our results to cover CIFAR10, CIFAR100, and TinyImageNet datasets, for both Class-Incremental and Task-Incremental settings.
+We also provide results for Domain-Incremental setting with Permuted-MNIST goodfellow2013empirical and Rotated-MNIST datasets.
+Table 1: Summary of average accuracies and forgetting values obtained on MNIST, FMNIST, EMNIST (5 tasks, Class-Incremental), and on Permuted-MNIST and Rotated-MNIST (10 tasks, Domain-Incremental).
+Method / Dataset MNIST FMNIST EMNIST
+Acc / For Acc / For Acc / For
+CoP2L (MLP) 95.1 / 3.4 84.2 / 12.1 78.0 / 16.7
+Replay (MLP) 94.5 / 6.1 84.0 / 15.9 77.2 / 22.1
+GDumb (MLP) 91.7 / 3.8 81.8 / 10.4 62.6 / 16.0
+CoP2L (CNN) 96.6 / 2.2 87.4 / 11.1 87.9 / 8.4
+Replay (CNN) 96.9 / 3.3 87.6 / 13.3 87.7 / 12.3
+GDumb (CNN) 94.4 / 3.0 85.0 / 8.8 85.1 / 7.5
+Method / Dataset PermMNIST RotMNIST
+Acc / For Acc / For
+CoP2L (MLP) 92.3 / 6.0 88.3 / 9.8
+Replay (MLP) 91.5 / 6.4 88.2 / 10.0
+LFL (MLP) 83.5 / 9.3 51.6 / 45.7
+EWC (MLP) 83.7 / 5.2 56.5 / 35.4
+SI (MLP) 86.7 / 10.7 57.7 / 43.9
+Table 2: Summary of average accuracies and forgetting values obtained on various continual learning tasks on CIFAR10, CIFAR100 and TinyImageNet datasets.
+CI denotes Class-Incremental, TI denotes Task-Incremental continual learning settings.
+CIFAR10 (5 tasks), CI CIFAR100 (10 tasks), CI CIFAR100 (20 tasks), CI TinyImNet (40 tasks), CI
+Method Acc ($\uparrow$) Forg ($\downarrow$) Acc ($\uparrow$) Forg ($\downarrow$) Acc ($\uparrow$) Forg ($\downarrow$) Acc ($\uparrow$) Forg ($\downarrow$)
+CoP2L-Dino (ViT) 94.45 3.36 73.83 16.76 72.50 16.96 51.13 36.81
+Replay-Dino (ViT) 94.01 6.01 68.51 30.44 69.23 29.60 47.66 48.15
+iCaRL-Dino (ViT) 93.24 1.55 61.50 14.53 64.11 15.25 20.78 52.95
+GDumb-Dino (ViT) 94.15 3.39 72.73 11.70 72.05 11.38 54.53 14.79
+CoP2L-Dino (RN50) 81.13 5.55 49.18 21.99 49.98 24.32 32.74 36.40
+Replay-Dino (RN50) 81.83 17.09 49.40 42.07 49.91 44.41 28.43 63.26
+iCaRL-Dino (RN50) 74.56 1.54 45.97 8.71 46.87 6.32 33.13 10.95
+GDumb-Dino (RN50) 80.70 8.95 43.83 20.31 43.78 19.03 24.04 21.82
+CIFAR10 (5 tasks), TI CIFAR100 (10 tasks), TI CIFAR100 (20 tasks), TI TinyImNet (40 tasks), TI
+CoP2L-Dino (ViT) 99.01 -0.03 95.15 0.92 97.10 0.71 94.28 1.22
+Replay-Dino (ViT) 99.24 -0.03 96.59 0.23 98.06 0.41 95.82 0.36
+LaMAML-Dino (ViT) 99.18 0.06 95.77 0.40 97.70 0.33 94.90 1.15
+LwF-Dino (ViT) 97.53 2.08 95.26 1.06 95.73 2.15 85.79 9.09
+CoP2L-Dino (RN50) 96.17 0.79 86.58 2.26 90.88 2.37 88.75 2.13
+Replay-Dino (RN50) 96.96 0.42 87.59 0.64 93.13 0.16 90.53 0.74
+LaMAML-Dino (RN50) 96.71 0.66 83.13 6.81 90.57 3.35 87.25 4.55
+LwF-Dino (RN50) 95.90 0.15 79.29 0.67 87.94 0.09 85.15 0.90
+Figure 1: Example demonstration of emprical upper bounds on test errors, on Class-Incremental (top-left CIFAR10, top-right CIFAR100) and Task-Incremental (bottom-left CIFAR10, bottom-right CIFAR100) datasets.
+Empirical study of the CoP2L generalization bound.
+In Figure 1, we showcase the empirical upper bounds on test errors.
+The details on the computation of the bound can be found in the Appendix  A.
+We observe that the bound values are non-trivial and generally follow similar trends as the actual error we report on the test set, while the estimate for the bound is calculated exclusively on the training set.
+In Figure 1, we provide plots for the bounds for the models trained with a buffer size of 2000, for Class-Incremental and Task-Incremental settings.
+CoP2L achieves particularly tight generalization bounds in the Task-Incremental setting, where forgetting is inherently reduced due to access to task identities and the use of a separate classifier head for each task.
+These conditions allow the model to better compress task-specific knowledge, leading to both stronger theoretical guarantees and improved empirical performance.
+We also observe that with the ViT backbone, the bounds are significantly tighter than those obtained with the ResNet50 backbone.
+This is likely because the ViT backbone provides more structured and information-rich representations, which are easier to compress effectively, thereby enhancing the performance of CoP2L.
+We also give the results for all buffer sizes for the three MNIST datasets, and also bound estimates for the CIFAR10, CIFAR100, and ImageNet datasets (TinyImageNet tiny-imagenet , ImageNette imagenette and ImageWoof Howard_Imagewoof_2019 ) as well as bounds for the Domain-Incremental setting in Appendix D.
+4.2 Analysis of CoP2L
+In this section we provide an analysis of CoP2L with an ablation study on loss terms, a sensitivity analysis with respect to task ordering, and buffer size.
+Due to space limitations the figures are provided in the appendix, however the main results are mentioned here in the text.
+Ablation on CoP2L loss terms  We have conducted an ablation study to understand the effect of additional weighting that counteracts class imbalance that arises in the continual learning setup when new tasks are introduced (Standard replay counteracts this by extending the batch size to have equal class representation).
+In the modified P2L (Algorithm 2), we incorporate an additional weighting term over classes to adapt P2L for class-balanced continual learning.
+In order to assess whether this idea is indeed helpful, we have run an experiment on the MNIST dataset with and without the additional weighting idea.
+In Appendix E, in Figure 10, we show this experiment.
+For both accuracy (left panel) and forgetting (right panel), the inclusion of the additional loss weighting term significantly improves the results.
+[…]
+Analysis on buffer size.
+In Appendix E, in Figure 12, we compare the accuracy and forgetting obtained with different replay buffer sizes of (1000, 2000, 3000, 4000, 5000) samples for CoP2L and replay.
+This represents roughly 8% to 40% of the per-task dataset size, for MNIST and Fashion-MNIST, and 10% to 52% for EMNIST.
+For each replay buffer size, we report the variation of accuracy obtained at the end of training averaged over different seeds, and test accuracy for different tasks.
+For the MNIST dataset, on average CoP2L significantly outperforms replay (as evidenced by the variation over seeds).
+For the Fashion-MNIST and EMNIST datasets, when using small buffer sizes, the average accuracy obtained with our CoP2L is significantly better.
+For larger buffer sizes, in general CoP2L performs comparably or slightly better.
+We also note for all three datasets, CoP2L significantly outperforms replay in terms of forgetting.
+We attribute this due to the fact that CoP2L is able to intelligently select examples from the replay buffer.
+Namely, CoP2L is able to focus on examples that are most representative of the earlier tasks, which we observe to help significantly in mitigating the deterioration over earlier tasks as evidenced by the results presented in the right column in Figure 12.
+5 Limitations and Broader Impact Statement
+We observe with our experiments that the performance of CoP2L is dependent to the ability of the feature extractor (backbone).
+In case a weak backbone is employed, CoP2L could necessitate a larger compression set (e.g.
+more data points per task).
+We have also seen that in those cases CoP2L might take longer to converge.
+We noted in our experiments that using deep pretrained networks such as ViT results in good performance, by increasing the compression ability.
+More broadly, we do not believe there is any particular negative social impact to note for our proposed methodology.
+On the contrary, we believe that our algorithm could be instrumental for improving trustworthiness of continual learning / lifelong learning in practice.
+6 Conclusions
+In this paper, we have proposed CoP2L, an algorithm rooted in the sample compression theory for self-certified continual learning.
+To the best of our knowledge, this is the first attempt in the literature to employ sample compression theory within the continual learning context.
+We provided sample compression bounds for CoP2L and verified empirically that they are non-vacuous and informative.
+We furthermore showed that on several challenging datasets, that our approach yields comparable or better results compared several strong baselines, while also providing learning guarantees, for trustworthy continual learning.
+As the combination of Pick-to-Learn with experience replay leads to a successful learning scheme, we foresee that combining Pick-to-Learn with other continual learning approaches could also lead to new compelling ways of obtaining self-certified predictors.
+Acknowledgements
+Mathieu Bazinet is supported by a FRQNT B2X scholarship (343192, doi: https://doi.org/10.69777/343192).
+Pascal Germain, and Cem Subakan are supported by Discovery Grants RGPIN-2020-07223 and RGPIN 2023-05759, respectively.
+This research was also enabled in part by support provided by Calcul Québec (https://www.calculquebec.ca/) and the Digital Research Alliance of Canada (https://alliancecan.ca/en).
+Appendix A Compression and Reconstruction Functions
+In line with the prevalent literature, we introduced the sample compression framework (Section 2) based on a compression function providing a compression set and optionally a message.
 In the case of CoP2L, the reconstruction scheme relies on two compression sets $(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}})$ and a message pair $(\mu_{1},\mu_{2})$, which we explain later in this section.
-[^1] The idea of using multiple compression sets appeared in Marchand & Shawe-Taylor (2002) and Marchand et al.
-(2003), but in a setting without a message Two challenges prevent CoP2L from being used as its own compression and reconstruction function as the original P2L.
-The first challenge comes from the sampling of the buffer ${\mathcal{B}}$.
+[^2] The idea of using multiple compression sets appeared in [40] and [39], but in a setting without a message.
+Algorithm 4 Compression function of Continual Pick-To-Learn (CoP2L)
+0:  $S_{1},S_{2},\ldots,S_{T}$Training sets}
+0:  $\theta_{0}$Initialization parameters of a given model}
+0:  $\gamma$P2L’s stopping criteria}
+0:  $m$Buffer’s max sampling size}
+0:  $\omega$Weight for buffer tasks}
+1:  $B_{t}\leftarrow\emptyset,S^{\operatorname{\mathbf{i}}}_{t}\leftarrow\emptyset,S^{\operatorname{\mathbf{j}}}_{t}\leftarrow\emptyset,\mu_{1,t}\leftarrow\emptyset\quad\forall t=1,\ldots,T$
+2:  $B^{\star}\leftarrow\emptyset,\mu_{2}\leftarrow\emptyset$
+3:  for $t\in\{1,\ldots,T\}$ do
+4:    $\hat{S_{t}}\leftarrow\{({\mathbf{x}},y,1)\}_{({\mathbf{x}},y)\in S_{t}}$
+5:    $\theta_{t},C^{\star}\leftarrow\mbox{mP2L}(\theta_{t-1},\hat{S_{t}},B^{\star},\gamma,\infty)$
+6:    $\mu_{2}^{t}\leftarrow|C^{\star}|$
+7:    $S^{\operatorname{\mathbf{i}}}_{i}\leftarrow S^{\operatorname{\mathbf{i}}}_{i}\cup(C^{\star}\cap S_{i})\quad\forall i=1,\ldots,t$
+8:    $B_{i}\leftarrow\mbox{sample}(B_{i},\lfloor\frac{m}{t}\rfloor)\quad\forall i=1,\ldots,t-1$
+9:    $B_{t}\leftarrow\mbox{sample}(S_{t}\setminus C^{\star},\lfloor\frac{m}{t}\rfloor)$
+10:    for $i\in\{1,\ldots,t-1\}$ do
+11:       for $({\mathbf{x}}_{i,j},y_{i,j},\cdot)\in B^{\star}$ and $({\mathbf{x}}_{i,j},y_{i,j},\cdot)\notin B_{i}$ do
+12:          $S^{\operatorname{\mathbf{j}}}_{i}\leftarrow S^{\operatorname{\mathbf{j}}}_{i}\cup\{({\mathbf{x}}_{i,j},y_{i,j})\}$
+13:          $S^{\operatorname{\mathbf{i}}}_{i}\leftarrow S^{\operatorname{\mathbf{i}}}_{i}\setminus\{({\mathbf{x}}_{i,j},y_{i,j})\}$
+14:          $\mu_{1,i}^{j}\leftarrow\{t\}$
+15:       end for
+16:    end for
+17:    $B^{\star}\leftarrow\bigcup_{i=1}^{t}\{({\mathbf{x}},y,\omega)\}_{({\mathbf{x}},y,\cdot)\in B_{i}}$
+18:  end for
+19:  return $\theta_{T},\{S^{\operatorname{\mathbf{i}}}_{t}\}_{t=1}^{T},\{S^{\operatorname{\mathbf{j}}}_{t}\}_{t=1}^{T},\{\mu_{1,t}\}_{t=1}^{T},\mu_{2}$.
+{Learned parameters, compression sets and messages sets}
+A.1 Compression Function
+Two challenges prevent CoP2L (as presented by Algorithm 3) from being used as its own compression and reconstruction function.
+In this subsection, we expose each of these challenges while explaining how we the proposed compression algorithm (Algorithm 4) allows to circumvent them.
+The first challenge comes from the sampling of the buffer $B^{\star}$ (Lines 6 to 8 of Algorithm 3).
 When training using CoP2L, datapoints from the buffer can be chosen to be part of the compression set.
@@
 Therefore, the reconstruction function needs to know when to remove the datapoint.
-The message thus provide the task after which the datapoint was removed during CoP2L execution.
+The message $\mu_{1}$ returned by Algorithm 4 thus provides the task index after which the datapoint was removed during CoP2L execution.
 As the reconstruction function only requires a message for datapoints who were removed from the buffer after being added to the compression set $S^{\operatorname{\mathbf{i}}}$, a second compression set $S^{\operatorname{\mathbf{j}}}$ is dedicated to datapoints that necessitate a message.
 The message $\mu_{1}$ given is chosen among $\{2,\ldots,T\}$ for each datapoint belonging to $S^{\operatorname{\mathbf{j}}}$.
-The second one is the use of the bound as stopping criterion in mP2L.
+The second challenge is the use of the bound as stopping criterion in mP2L.
 Recall that the reconstruction function does not have access to the whole dataset.
@@
 After learning on $T$ tasks, the compression function provides the sample compression set $S^{\operatorname{\mathbf{i}}}$ and  $S^{\operatorname{\mathbf{j}}}$, along with a message pair $(\mu_{1},\mu_{2})$ chosen among the set of all possible messages, denoted as
-$\mathscr{M}(\operatorname{\mathbf{j}},T)\,=\,\{2,\ldots,T\}^{|\operatorname{\mathbf{j}}|}\times\left[\bigtimes\limits_{t=1}^{T}\{1,\ldots,n_{t}+|{\mathcal{B}}|\}\right].$ (9)
-Reconstruction function.
-Under the assumption that the input parameters $\theta_{0},\gamma,m,\omega$ are the same for the compression function and for CoP2L, the reconstruction function must output the exact same predictor as CoP2L.
+$\mathscr{M}(\operatorname{\mathbf{j}},T)\,=\,\{2,\ldots,T\}^{|\operatorname{\mathbf{j}}|}\times\left[\textstyle\bigtimes_{t=1}^{T}\{1,\ldots,n_{t}+|{\mathcal{B}}|\}\right].$ (7)
+A.2 Reconstruction Function
+Under the assumption that the input parameters $\theta_{0},\gamma,m,\omega$ are the same for the compression function (Algorithm 4) and for CoP2L (Algorithm 3), the composition of the compression and the reconstruction function must output the exact same predictor as CoP2L.
 However, it only has access to the compression sets and the messages.
@@
 If so, we remove it from the buffer.
-The second challenge is adressed in line 4 of Algorithm 5 by giving the message $\mu_{2}$ to mP2L as stopping criterion.
-Algorithm 4 Compression function of CoP2L
+The second challenge is addressed in line 4 of Algorithm 5 by giving the message $\mu_{2}$ to mP2L as stopping criterion.
+Algorithm 5 Reconstruction function of Continual Pick-To-Learn (CoP2L)
+0:  $S^{\operatorname{\mathbf{i}}}_{1},S^{\operatorname{\mathbf{i}}}_{2},\ldots,S^{\operatorname{\mathbf{i}}}_{T}$Compression sets}
+0:  $S^{\operatorname{\mathbf{j}}}_{1},S^{\operatorname{\mathbf{j}}}_{2},\ldots,S^{\operatorname{\mathbf{j}}}_{T}$Compression sets}
+0:  $\mu_{1,1},\mu_{1,2},\ldots,\mu_{1,T}${Message sets}
+0:  $\mu_{2}${Message set}
 0:  $\theta_{0}$Initialization parameters of a given model}
@@
 0:  $\omega$Weight for buffer tasks}
-0:  $S_{1},S_{2},\ldots,S_{T}$Training sets}
-1:  $B_{i}\leftarrow\emptyset,C_{i}\leftarrow\emptyset,C^{\mu}_{i}\leftarrow\emptyset,\mu_{1,i}\leftarrow\emptyset\quad\forall i=1,\ldots,T$
-2:  $B^{\star}\leftarrow\emptyset,\mu_{2}\leftarrow\emptyset$
-3:  for $t\in\{1,\ldots,T\}$ do
-4:     $\hat{S_{t}}\leftarrow\{({\mathbf{x}},y,1)\}_{({\mathbf{x}},y)\in S_{t}}$
-5:     $\theta_{t},C^{\star}\leftarrow\mbox{mP2L}(\theta_{t-1},\hat{S_{t}},B^{\star},\gamma,\infty)$
-6:     $\mu_{2}^{t}\leftarrow|C^{\star}|$
-7:     $C_{i}\leftarrow C_{i}\cup(C^{\star}\cap S_{i})\quad\forall i=1,\ldots,t$
-8:     $B_{i}\leftarrow\mbox{sample}(B_{i},\lfloor\frac{m}{t}\rfloor)\quad\forall i=1,\ldots,t-1$
-9:     $B_{t}\leftarrow\mbox{sample}(S_{t}\setminus C^{\star},\lfloor\frac{m}{t}\rfloor)$
-10:     for $i\in\{1,\ldots,t-1\}$ do
-11:        for $({\mathbf{x}}_{i,j},y_{i,j},\cdot)\in B^{\star}$ and $({\mathbf{x}}_{i,j},y_{i,j},\cdot)\notin B_{i}$ do
-12:           $C^{\mu}_{i}\leftarrow C^{\mu}_{i}\cup\{({\mathbf{x}}_{i,j},y_{i,j})\}$
-13:           $C_{i}\leftarrow C_{i}\setminus\{({\mathbf{x}}_{i,j},y_{i,j})\}$
-14:           $\mu_{1,i}^{j}\leftarrow\{t\}$
-15:        end for
-16:     end for
-17:     $B^{\star}\leftarrow\bigcup_{i=1}^{t}\{({\mathbf{x}},y,\omega)\}_{({\mathbf{x}},y,\cdot)\in B_{i}}$
-18:  end for
-19:  return $\theta_{T},\{C_{t}\}_{t=1}^{T},\{C^{\mu}_{t}\}_{t=1}^{T},\{M_{t}\}_{t=1}^{T},N$. {Learned parameters, compression sets and messages sets}
-Algorithm 5 Reconstruction function of CoP2L
-0:  $\theta_{0}$Initialization parameters of a given model}
-0:  $\gamma$P2L’s stopping criteria}
-0:  $m$Buffer’s max sampling size}
-0:  $\omega$Weight for buffer tasks}
-0:  $C_{1},C_{2},\ldots,C_{T},\ C^{\mu}_{1},C^{\mu}_{2},\ldots,C^{\mu}_{T}$Compression sets}
-0:  $\mu_{1,1},\mu_{1,2},\ldots,\mu_{1,T},\,\mu_{2}${Message sets}
 1:  $B^{\star}\leftarrow\emptyset$
 2:  for $t\in\{1,\ldots,T\}$ do
-3:     $\hat{S_{t}}\leftarrow\{({\mathbf{x}},y,1)\}_{({\mathbf{x}},y)\in C_{t}}$
+3:    $\hat{S_{t}}\leftarrow\{({\mathbf{x}},y,1)\}_{({\mathbf{x}},y)\in S^{\operatorname{\mathbf{i}}}_{t}}$
 4:    $\theta_{t},C^{\star}\leftarrow\mbox{mP2L}(\theta_{t-1},\hat{S_{t}},B^{\star},\gamma,\mu_{2}^{t})$
-5:     $B_{t}\leftarrow(C_{t}\setminus C^{\star})\cup C^{\mu}_{t}$
+5:    $B_{t}\leftarrow(S^{\operatorname{\mathbf{i}}}_{t}\setminus C^{\star})\cup S^{\operatorname{\mathbf{j}}}_{t}$
 6:    for $i\in\{1,\ldots,t-1\}$ do
-7:        for $(\boldsymbol{x}_{i,j},y_{i,j},\omega_{i})\in C^{\mu}_{i}$ do
+7:       for $(\boldsymbol{x}_{i,j},y_{i,j},\omega_{i})\in S^{\operatorname{\mathbf{j}}}_{i}$ do
 8:          if $\mu_{1,i}^{j}=t$ then
@@
 15:  return $\theta_{T}$. {Learned parameters}
-3.3 Generalization Bound
-The last subsection provided the compression and reconstruction functions of CoP2L.
-Thus, we know CoP2L provides a sample-compressed predictor.
-Theorem 2.1 can be used to obtain generalization bounds on the last task, which we do in mP2L when we compute the bound $\Psi$, but it holds for only one distribution of data.
-Therefore, Theorem 2.1 cannot be applied to bound all tasks learned by our continual learning scheme.
-Le us now consider a reconstruction function that takes as input two compression sets $\operatorname{\mathbf{i}},\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)$ and a message $\mu=(\mu_{1},\mu_{2})\in\mathscr{M}(\operatorname{\mathbf{j}},T)$.
-For any distribution $\operatorname{\mathcal{D}}_{t}$ and any dataset $S_{t}\sim\operatorname{\mathcal{D}}_{t}$, we denote the loss on both compression sets as ^$\widehat{{\mathcal{L}}}_{S_{t}}^{\ \operatorname{\mathbf{i}}}(\theta)$ and ^$\widehat{{\mathcal{L}}}_{S_{t}}^{\ \operatorname{\mathbf{j}}}(\theta)$.
-The loss on the joint compression sets is denoted ^$\widehat{{\mathcal{L}}}_{S_{t}}^{\ \operatorname{\mathbf{i}}\cup\operatorname{\mathbf{j}}}(\theta)$ and the loss on the complement set of the joint set $\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}$ is denoted ^$\widehat{{\mathcal{L}}}_{S_{t}}^{\ \operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\!(\theta)$.
-Moreover, we denote the reconstruction function of CoP2L as $\operatorname{\mathscr{R}}_{1:T}(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)=\operatorname{\mathscr{R}}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu;S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}).$ This formulation is important for the following theorem, as the reconstruction function should be fixed for all datasets $S_{1}$ to $S_{T}$, with the exception of $S_{t}$.
-Theorem 3.1.
-For any set of independent distributions $\left\{{\mathcal{D}}_{t}\right\}_{t=1}^{T}$ over ${\mathcal{X}}\times{\mathcal{Y}}$, with $\operatorname{\mathscr{R}}_{1:T}$ the reconstruction function of CoP2L (Algorithm 5), with the family of set of messages $\{\mathscr{M}(\operatorname{\mathbf{j}},T)|\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)\}$ defined for CoP2L, for any loss $\ell:\Theta\times{\mathcal{X}}\times{\mathcal{Y}}\to[0,1]$, and for any $\delta\in(0,1]$, with probability at least $1-\delta$ over the draw of $S_{t}\sim{\mathcal{D}}_{t},t=1,\ldots,T$, we have
-$\forall t\in\{1,\ldots,T\},{\mathbf{i}},\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n_{t}),\mu\in\mathscr{M}(\operatorname{\mathbf{j}},T):$
-$\mathrm{kl}\quantity(\widehat{{\mathcal{L}}}_{S_{t}}^{\ \operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\!\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)),\ {\mathcal{L}}_{\operatorname{\mathcal{D}}_{t}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)))$
-$\leq\frac{1}{n_{t}-|{\mathbf{i}}|-|\mathbf{j}|}\bigg{[}\log\matrixquantity(n_{t}\\ |{\mathbf{i}}|)+\log\matrixquantity(n_{t}-\operatorname{|\mathbf{i}|}\\ |\mathbf{j}|)$
-$~{}~{}~{}~{}~{}~{}~{}~{}~{}+\sum_{i=1}^{T}\log\frac{1}{\zeta(\mu_{2}^{i})}+|\mathbf{j}|\log(T-1)+\log\frac{T}{\zeta(|{\mathbf{i}}|)\zeta(|\operatorname{\mathbf{j}}|)\delta}\bigg{]}$
-with $\mu=(\mu_{1},\mu_{2})$.
-The proof is given in Appendix B.
-It first relies on a tighter version of Theorem 2.1, where the $2\sqrt{n-\operatorname{|\mathbf{i}|}}$ term is removed and a second compression set is added.
-Then, this new result is adapted to the continual learning setting.
-For all tasks $1$ through $T$, this new theorem encodes a tradeoff similar to Theorem 2.1.
-A more accurate predictor will achieve a tighter bound.
-However, a simpler predictor will achieve a tighter bound.
-In the sample compression theory, we define the complexity of the model using the size of the compression sets $\operatorname{\mathbf{i}}$ and $\operatorname{\mathbf{j}}$ and the probability of choosing a message.
-The probability of the messages $\mu_{1}$ and $\mu_{2}$ are functions of the size of $\operatorname{\mathbf{j}}$ and the number of tasks.
-Indeed, the probability of $\mu_{1}$ is independent of its content but decays when $|\operatorname{\mathbf{j}}|$ and $T$ grow larger.
-Moreover, the probability of $\mu_{2}$ depends on the size of the compression set outputted by mP2L at each task $T$.
-Thus, the larger the compression sets are, the smaller the probability of both messages will be and the complexity of the model grows when the compression sets grow.
-In conclusion, the model that will achieve the best possible bound is a model that minimizes the loss on the dataset whilst still keeping $\operatorname{\mathbf{i}}$ and $\operatorname{\mathbf{j}}$ the smallest possible.
-4 Experiments
-Figure 1: Accuracy (above) and forgetting (below) over replay buffer size for (left column) MNIST dataset, (middle column) Fashion-MNIST dataset, (right column) EMNIST dataset.
-The orange triangles show the average performance for the corresponding box plot.
-Figure 2: CoP2L bounds with respect to tasks on the MNIST and Fashion-MNIST datasets
-Datasets We carry-out experiments on MNIST (LeCun et al., 2010), Fashion-MNIST (Xiao et al., 2017), and EMNIST (Cohen et al., 2017) datasets on class-incremental learning.
-We divide the tasks into increments of 2 classes.
-This yields 5 tasks for MNIST and Fashion-MNIST datasets, and 13 tasks for the EMNIST dataset.
-Implementation Details We have used the Avalanche continual learning framework for our experimental setup (Carta et al., 2023).
-Avalanche is a popular toolkit for continual learning.
-We have used their class-incremental MNIST and Fashion-MNIST dataset implementations, as well as the replay baseline in order to ensure compatibility with the results in the literature.
-We implement a block version of mP2L, which adds $k$ examples to the compression set at a time, as described in Algorithm 2 of Paccagnan et al.
-(2024).
-For the experiments, we use $k=8$.
-We provide the code for our experimental results in an anonymous repository [^2] https://anonymous.4open.science/r/sample_compression_continual_learning-00C7..
-Model architecture and training details We use a simple multilayer perceptron (MLP) architecture for all experiments (SimpleMLP class from the Avalanche toolkit), and train using the SGD optimizer, with a learning rate of 0.001 for CoP2L and 0.01 for Replay.
-For all experiments with the Replay method, we train the model 20 epochs on each tasks.
-In contrast, CoP2L follows a different training strategy, continuing until it reaches zero errors on the complement set.
-We add a weight of $\omega=15$ in the cross-entropy loss when training on previous tasks.
-Computation of the bound To compute the bound, we need to keep track of the size of the compression sets and of the messages.
-We thus implement the compression function as described in Algorithm 4.
-However, we need to adjust the code to account for mP2L adding $k$ datapoints to the compression set at each iterations.
-When computing the message $\mu_{2}$, at iteration $t$, $\mu_{2}^{t}$ is the number of iterations of the mP2L algorithm.
+A.3 Computation of the Bound
+To compute the bound of Theorem 3.1, we need three things : the size of the first compression set $\operatorname{\mathbf{i}}$, the size of the second compression set $\operatorname{\mathbf{j}}$ and the message $\mu$.
+By simply running our CoP2L, we are able to retrieve $|\operatorname{\mathbf{i}}|+|\operatorname{\mathbf{j}}|$, but not the size of each compression set individually.
+Moreover, we are completely unable to retrieve the messages.
+Thus, instead of implementing Algorithm 3, we actually implemented the compression function as described in Algorithm 4.
+Moreover, we use a block version of P2L, as defined by Algorithm 2 of [23].
+We thus add $k$ datapoints at once to the compression set at Line 6 of Algorithm 2 (the used values of $k$ for each experiments is given in Appendix G.
+Thus, the value of $\mu_{2}^{t}$ (Line 6 of Algorithm 4) becomes the number of iterations of the mP2L algorithm at for task $t$.
 Knowing that mP2L outputs a compression set $C^{\star}_{t}$, and adds $k$ datapoints to the compression set at each iteration, the number of iterations is $\mu_{2}^{t}=\frac{|C^{\star}_{t}|}{k}$.
-Finally, when computing the bound, we use the code provided by Viallard et al.
-(2021) to invert the $\mathrm{kl}$ divergence.
-4.1 Metric Comparison over Buffer Size
+Finally, when computing the bound, we use the code provided by [65] to invert the $\mathrm{kl}$ divergence.
+Appendix B Tighter sample compression $\mathrm{kl}$ bound
+In this section, we prove a tighter $\mathrm{kl}$ bound that is not derived from Theorem 2.1 but proved in a very similar way.
+To do so, we start by providing the Chernoff test-set bound [66] for losses in $[0,1]$ [67].
+Theorem B.1 ([67]).
+Let $X_{1},\ldots,X_{n}$ be i.i.d. random variables with $X_{i}\in[0,1]$ and $\operatorname*{\mathbb{E}}[X_{i}]=p$. Then, for any $\delta\in(0,1]$, with probability at least $1-\delta$
+$p\leq\mathrm{kl}^{-1}\quantity(\tfrac{1}{n}\textstyle\sum_{i=1}^{n}X_{i},\tfrac{1}{n}\log\tfrac{1}{\delta}).$
+We now prove a tighter $\mathrm{kl}$ sample compression bound.
+As our algorithm CoP2L needs two compression sets, we consider a second compression set $\operatorname{\mathbf{j}}$.
+We define a conditional probability distribution $P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})$ that incorporates the knowledge that a vector $\operatorname{\mathbf{i}}$ was already drawn from $\operatorname{\mathscr{P}}(n)$ and that $\operatorname{\mathbf{i}}\cap\operatorname{\mathbf{j}}=\emptyset$.
+Thus, we have $\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})\leq 1.$ If the choice of $\operatorname{\mathbf{j}}$ isn’t conditional to the choice of $\operatorname{\mathbf{i}}$, we simply have $P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})=P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}})$.
+Theorem B.2.
+For any distribution $\operatorname{\mathcal{D}}$ over $\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}$, for any family of set of messages $\{\mathscr{M}(\operatorname{\mathbf{j}})\,|\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)\}$, for any deterministic reconstruction function $\operatorname{\mathscr{R}}$ , for any loss $\ell:\Theta\times\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}\to[0,1]$ and for any $\delta\in(0,1]$, with probability at least $1-\delta$ over the draw of $S\sim\operatorname{\mathcal{D}}^{n}$, we have
+$\forall\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n),\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n),\mu\in\mathscr{M}(\operatorname{\mathbf{j}}):$
+$\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta})$
+Proof.
+We prove the complement of the expression in Theorem B.2. Denote the upper bound
+$U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)=\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta})$
+$\operatorname*{\mathbb{P}}_{S\sim\operatorname{\mathcal{D}}^{n}}\quantity(\exists\operatorname{\mathbf{i}},\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n),\mu\in\mathscr{M}(\operatorname{\mathbf{j}}):\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))>U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta))$
+$\leq\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\operatorname*{\mathbb{P}}_{S\sim\operatorname{\mathcal{D}}^{n}}\quantity(\exists\mu\in\mathscr{M}(\operatorname{\mathbf{j}}):\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))>U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta))$ (8)
+$\leq\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\sum_{\mu\in\mathscr{M}(\operatorname{\mathbf{j}})}\operatorname*{\mathbb{P}}_{S\sim\operatorname{\mathcal{D}}^{n}}\quantity(\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))>U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta))$ (9)
+$\leq\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\sum_{\mu\in\mathscr{M}(\operatorname{\mathbf{j}})}\operatorname*{\mathbb{E}}_{S^{\operatorname{\mathbf{i}}}\sim\operatorname{\mathcal{D}}^{\operatorname{|\mathbf{i}|}}}\operatorname*{\mathbb{E}}_{S^{\operatorname{\mathbf{j}}}\sim\operatorname{\mathcal{D}}^{|\operatorname{\mathbf{j}}|}}\operatorname*{\mathbb{P}}_{S^{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\sim\operatorname{\mathcal{D}}^{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}}\quantity(\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))>U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta))$ (10)
+$\leq\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\sum_{\mu\in\mathscr{M}(\operatorname{\mathbf{j}})}\operatorname*{\mathbb{E}}_{S^{\operatorname{\mathbf{i}}}\sim\operatorname{\mathcal{D}}^{\operatorname{|\mathbf{i}|}}}\operatorname*{\mathbb{E}}_{S^{\operatorname{\mathbf{j}}}\sim\operatorname{\mathcal{D}}^{|\operatorname{\mathbf{j}}|}}P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta$ (11)
+$\leq\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\sum_{\mu\in\mathscr{M}(\operatorname{\mathbf{j}})}P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta$
+$\leq\delta.$
+We use the union bound in Equation 8 and Equation 9.
+In Equation 10, we use the i.i.d.
+assumption.
+In Equation 11 we use the Chernoff Test-set bound of Theorem B.1 with $p=\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))=\operatorname*{\mathbb{E}}_{(\boldsymbol{x},y)\sim\operatorname{\mathcal{D}}}\ell(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu),\boldsymbol{x},y)$ and $X_{i}=\ell(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu),\boldsymbol{x}_{i},y_{i})\forall i\in\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}$.
+Finally, the last line is derived from the definition of $P_{\operatorname{\mathscr{P}}(n)}$ and $P_{\mathscr{M}(\operatorname{\mathbf{j}})}$.
+∎
+Appendix C Proof of Theorem 3.1
+See 3.1
+Proof.
+Let us choose $t\in[1,T]$.
+Let us sample $S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}$.
+We denote the reconstruction function of CoP2L as $\operatorname{\mathscr{R}}_{1:T}(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)=\operatorname{\mathscr{R}}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu;S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}).$
+Let
+$U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)=\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n_{t}-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta}).$
+Then, we have :
+$\operatorname*{\mathbb{P}}_{S^{t}}\quantity(\forall\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n_{t}),\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n_{t}),\mu\in\mathscr{M}(\operatorname{\mathbf{j}}):\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta))$
+As all the datasets (except $S^{t}$) are sampled beforehand, we can define $\operatorname{\mathscr{R}}_{1:T}(\cdot)=\operatorname{\mathscr{R}}\quantity(\cdot;S^{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T})$ before drawing $S^{t}$.
+Thus, the reconstruction function is only a function of the dataset $S^{t}$, which is the setting of the result of Theorem B.2.
+We can then lower bound this probability by $1-\delta$.
+We finish the proof by applying the bound to all datasets. We have
+$\operatorname*{\mathbb{P}}_{S_{1},\ldots,S_{T}}\Biggl{(}\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)\Biggr{)}$
+$=$ $\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{T}}\operatorname{\mathbb{I}}\Bigg{(}\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)\Biggr{)}$
+$=$ $\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}}\operatorname*{\mathbb{E}}_{S_{t}}\operatorname{\mathbb{I}}\Bigg{(}\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)\Biggr{)}$
+$=$ $\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}}\operatorname*{\mathbb{P}}_{S_{t}}\Bigg{(}\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)\Biggr{)}$
+$\geq$ $\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}}1-\delta$
+$\geq$ $1-\delta.$
+We use a union bound argument to add the $\forall t\in[1,T]$, leading to the term $\frac{\delta}{T}$.
+As this theorem holds specifically for CoP2L, we specify the distributions $P_{\operatorname{\mathscr{P}}(n_{t})}$ and $P_{\mathscr{M}(\operatorname{\mathbf{j}})}$.
+Following the work of [39], which used sample compression bounds with three compression sets, with $\zeta(k)=\frac{6}{\pi}(k+1)^{-1}$, we choose
+$P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{i}})=\matrixquantity(n_{t}\\ \operatorname{|\mathbf{i}|})^{-1}\zeta(\operatorname{|\mathbf{i}|})\quad\mbox{and }\quad P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})=\matrixquantity(n_{t}-\operatorname{|\mathbf{i}|}\\ |\operatorname{\mathbf{j}}|)^{-1}\zeta(|\operatorname{\mathbf{j}}|)\,.$
+Finally, we split the message $\mu$ into two messages $\mu_{1}$ and $\mu_{2}$.
+The first message is defined using the alphabet $\Sigma_{1}=\{2,\ldots,T\}$.
+This message indicates the task number where a datapoint is sampled out of the buffer.
+Thus, the size of $\Sigma_{1}$ is $T-1$ and the probability of a symbol is $\frac{1}{T-1}$.
+For any vector $\operatorname{\mathbf{j}}$, we have a sequence of length $|\operatorname{\mathbf{j}}|$, and thus the probability of choosing each sequence is $(\frac{1}{T-1})^{|\operatorname{\mathbf{j}}|}$.
+The second message $\mu_{2}$ is defined using the alphabet $\Sigma_{2}=\{1,\ldots,n_{t}\}$.
+We choose to use $\zeta(k)=\frac{6}{\pi^{2}}(k+1)^{-2}$ as the probability distribution over each character.
+We know that for any $N\geq 0$, $\sum_{k=1}^{N}\zeta(k)\leq 1$.
+Thus, we have the probability of a sequence $\mu_{2}=\mu_{2}^{1}\ldots\mu_{2}^{T}$ is $\prod_{i=1}^{T}\zeta(\mu_{2}^{i})$.
+We define $\mathscr{M}(\operatorname{\mathbf{j}},T)$ the set of messages such that the sequence of symbols from $\Sigma_{1}$ is of length $|\operatorname{\mathbf{j}}|$ and the sequence of symbols from $\Sigma_{2}$ is of length $T$ (Equation 7).
+∎
+Appendix D Bound Plots
+D.1 Bound plots for different buffer sizes
+In this section, we present the computed bound values over task risks on MNIST, Fashion-MNIST and EMNIST datasets (letters).
+In Figure 2, we present results for different replay buffer sizes (1000, 2000, 3000, 4000, 5000), obtained using the MLP architecture.
+We also show the results obtained with the CNN architecture in Figure 3.
+We report the estimated deviation of the bound values over different seeds using the shaded regions which show the $\pm$1 standard deviation, and also the whiskers (plotted using the seaborn toolkit’s boxplot function) which show the spread of the values excluding the outliers.
+This plotting style for showing the deviation across seeds applies to all bound plots presented in this paper.
+Figure 2: CoP2L Bounds with respect to tasks on the MNIST, Fashion-MNIST and EMNIST datasets using an MLP architecture
+Figure 3: CoP2L Bounds with respect to tasks on the MNIST, Fashion-MNIST and EMNIST datasets using a CNN architecture
+We observe from Figures 2 and 3 that the MLP architecture and the CNN architecture give non-trivial bounds that follow the trends exhibited by the test error curve.
+D.2 Bound plots for CIFAR10, CIFAR100, TinyImageNet and ImageNet (subsets) datasets
+In this section, we present the plots for the bound values obtained for the additional datasets, namely CIFAR10, CIFAR100, and Imagenet subsets.
+The Imagenet subsets include the Tinyimagenet [51], ImageWoof [64], and ImageNette [63] datasets.
+We provide the results on ImageWoof and ImageNette to showcase that the estimated bounds get tighter when more data is avaible per each task.
+Figure 4: CoP2L bounds with respect to tasks on the CIFAR10 (5 tasks) dataset in Class-Incremental (left) and Task-Incremental (right) settings
+Figure 5: CoP2L bounds with respect to tasks on the CIFAR100 (10 tasks) dataset in Class-Incremental (left) and Task-Incremental (right) settings
+Figure 6: CoP2L bounds with respect to tasks on the CIFAR100 (20 tasks) dataset in Class-Incremental (left) and Task-Incremental (right) settings
+Figure 7: CoP2L bounds with respect to tasks on the TinyImageNet (40 tasks) dataset in Class-Incremental (left) and Task-Incremental (right) settings
+Figure 8: CoP2L bounds with respect to tasks on the ImageNette (5 tasks) dataset (left) and ImageWoof dataset (right) in Task-Incremental setting
+Figure 9: CoP2L bounds with respect to tasks on the PermutedMNIST (10 tasks) dataset (left) and on the RotatedMNIST dataset in Domain-Incremental setting
+Appendix E Analysis of CoP2L
+In this section we provide an analysis of CoP2L with an ablation study on loss terms, a sensitivity analysis with respect to task ordering, and buffer size.
+The text matches the one provided in the main paper, but we also include the aforementioned figures.
+Ablation on CoP2L loss terms.
+We have conducted an ablation study to understand the effect of additional weighting that counteracts class imbalance that arises in the continual learning setup when new tasks are introduced (Standard replay counteracts this by extending the batch size to have equal class representation).
+In the modified P2L (Algorithm 2), we incorporate an additional weighting term over classes to adapt P2L for class-balanced continual learning.
+In order to assess whether this idea is indeed helpful, we have run an experiment on the MNIST dataset with and without the additional weighting idea.
+In Figure 10, we show the results of this experiment.
+For both accuracy (left panel) and forgetting (right panel), we observe that the inclusion of the additional loss weighting term significantly improves the results.
+[…]
+Figure 10: Ablation study on MNIST and Fashion-MNIST on the additional weighting to ensure class balance we propose in Algorithm 2 on the CoP2L loss function.
+Accuracy (left) and forgetting (right) over replay buffer size for the MNIST dataset
+Figure 11: Studying the variation of cumulative accuracy with respect to task ordering for CoP2L and Replay on MNIST dataset (left) and Fashion-MNIST dataset (right).
+We use 5 different task order, and show the variability with respect to each task.
+E.1 Analysis with respect to buffersize
 In Figure 12, we compare the accuracy and forgetting obtained with different replay buffer sizes of (1000, 2000, 3000, 4000, 5000) samples for CoP2L and replay.
@@
 For each replay buffer size, we report the variation of accuracy obtained at the end of training averaged over different seeds, and test accuracy for different tasks.
-Note that we denote the accuracy obtained on a task $t$, after training the model for $T$ tasks with $\text{Accuracy}_{T,t}$.
-The accuracy we report is defined as
-$\text{Average Accuracy}_{T}=\frac{1}{s\cdot T}\sum_{i=1}^{s}\sum_{t=1}^{T}\text{Accuracy}_{T,t}^{i},$ (10)
-where $i$ index is for different seeds, and $t$ is for the task index.
-In our experiments, we used $s=5$ seeds to measure the variability with respect to weight initialization.
-We also report the average forgetting at task $T$ which is defined as follows,
-$\text{Average Forgetting}_{T}=$ (11)
-$\frac{1}{(T-1)\cdot s}\sum_{t=1}^{T-1}\sum_{i=1}^{s}$ $\text{Accuracy}_{t,t}^{i}-\text{Accuracy}_{T,t}^{i},$
-and can be calculated when $T>1$. Note that in Figure 1, we report the average accuracy obtained after the final task, such that $T=5$ for MNIST and Fashion-MNIST and $T=13$ for EMNIST.
+Note that in Figure 12, we report the average accuracy obtained after the final task, such that $T=5$ for MNIST and Fashion-MNIST and $T=13$ for EMNIST.
 We observe that for the MNIST dataset, on average CoP2L significantly outperforms replay (as evidenced by the variation over seeds).
@@
 Namely, CoP2L is able to focus on examples that are most representative of the earlier tasks, which we observe to help significantly in mitigating the deterioration over earlier tasks as evidenced by the results presented in the right column in Figure 12.
-4.2 Empirical Study of the CoP2L generalization bound
-In Figure 2, we showcase the generalization bound of CoP2L.
-We observe that the bound values for the generalization error always remain larger than the actual error incurred after each task, and we observe that the reported bound values are non-trivial.
-We also observe that the bound values generally follow similar trends as the actual error we report on the test set, while the estimate for the bound is calculated exclusively on the training set.
-In Figure 2, we provide the bounds for the models trained with a buffersize of 1000, but we give the results for all buffersizes, and all three datasets in Appendix C.
-Figure 3: Ablation study on MNIST and Fashion-MNIST on the additional weighting to ensure class balance we propose in Algorithm 2 on the CoP2L loss function.
-Accuracy (left) and forgetting (right) over replay buffer size for the MNIST dataset
-Figure 4: Studying the variation of cumulative accuracy with respect to task ordering for CoP2L and Replay on MNIST dataset (left) and Fashion-MNIST dataset (right).
-We use 5 different task order, and show the variability with respect to each task.
-4.3 Ablation on CoP2L loss terms
-We have conducted an ablation study to understand the effect of additional weighting that counteracts class imbalance that arises in the continual learning setup when new tasks are introduced (Standard replay counteracts this by extending the batch size to have equal class representation).
-In the modified P2L algorithm (2), we incorporate an additional weighting term over classes to adapt P2L for class-balanced continual learning.
-In order to assess whether this idea is indeed helpful, we have run an experiment on the MNIST dataset with and without the additional weighting idea.
-In Figure 3, we show this experiment.
-We observe that for both accuracy (left panel) and forgetting (right panel), the inclusion of the additional loss weighting term significantly improves the results.
-4.4 Sensitivity analysis with respect to the task ordering
-In order to also assess the sensitivity of CoP2L with respect to ordering of the tasks, on MNIST and Fashion-MNIST experiments we have conducted an experiment where we randomly shuffle the task order.
-For both dataset, we have tested our algorithm with 5 different random task ordering.
-For each task, we report the cumulative accuracy
+Figure 12: Accuracy (left) and forgetting (right) over replay buffer size for (top row) MNIST dataset, (middle row) Fashion-MNIST dataset, (bottom row) EMNIST dataset.
+The orange triangles show the average performance for the corresponding box plot.
+Appendix F All results with standard deviation and training time measurement
+In this section we provide the average accuracy values obtained on MNIST, FMNIST, EMNIST, CIFAR10, CIFAR100, TinyImagenet, PermutedMNIST and RotatedMNIST datasets with the standard deviation estimates.
+Note that Table 2 of the main paper provides the results obtained with a single seed, whereas the following tables provide the average results over five seeds, and the corresponding standard deviations.
+Note that we use different random number generator seeds to account for variability in random number generations (e.g.
+model parameter initializations) – but train / test splits are kept fixed with respect to seeds across different experiments.
+Note that, for CIFAR10, CIFAR100, TinyImagenet, ImageWoof, ImageNette, PermutedMNIST and RotatedMNIST experiments we also report the total time it takes to complete each experiment (for all tasks).
+CIFAR10 (5 tasks) Class-Incremental
+Methods Stream Accuracy (%) Stream Forgetting (%) Total time (s)
+CoP2L-Dino (ViT) $94.45\pm 0.14$ $2.10\pm 0.28$ $118.07\pm 2.59$
+Replay-Dino (ViT) $94.00\pm 0.19$ $6.11\pm 0.25$ $486.25\pm 43.07$
+iCaRL-Dino (ViT) $88.60\pm 10.04$ $3.13\pm 4.20$ $884.59\pm 1.39$
+GDumb-Dino (ViT) $94.16\pm 0.24$ $3.16\pm 0.39$ $47.72\pm 0.12$
+CoP2L-Dino (ResNet50) $80.98\pm 0.17$ $5.84\pm 0.37$ $527.95\pm 8.43$
+Replay-Dino (ResNet50) $81.38\pm 0.39$ $17.72\pm 0.51$ $287.38\pm 25.06$
+iCaRL-Dino (ResNet50) $74.49\pm 0.22$ $1.79\pm 0.43$ $579.17\pm 2.39$
+GDumb-Dino (ResNet50) $80.84\pm 0.14$ $9.35\pm 0.30$ $31.98\pm 0.12$
+CIFAR100 (10 tasks) Class-Incremental
+Methods Stream Accuracy (%) Stream Forgetting (%) Total time (s)
+CoP2L-Dino (ViT) $73.49\pm 0.27$ $17.06\pm 0.29$ $522.18\pm 3.90$
+Replay-Dino (ViT) $68.71\pm 0.25$ $30.02\pm 0.31$ $453.04\pm 11.93$
+iCaRL-Dino (ViT) $65.54\pm 2.23$ $13.64\pm 0.64$ $2491.81\pm 2889.82$
+GDumb-Dino (ViT) $72.51\pm 0.26$ $11.53\pm 0.65$ $95.43\pm 0.34$
+CoP2L-Dino (ResNet50) $49.53\pm 0.24$ $22.31\pm 0.20$ $1724.15\pm 23.50$
+Replay-Dino (ResNet50) $49.41\pm 0.38$ $42.12\pm 0.54$ $295.80\pm 15.88$
+iCaRL-Dino (ResNet50) $46.06\pm 0.40$ $8.50\pm 0.28$ $676.76\pm 56.25$
+GDumb-Dino (ResNet50) $43.37\pm 0.24$ $20.23\pm 0.42$ $63.58\pm 0.36$
+CIFAR100 (20 tasks) Class-Incremental
+Methods Stream Accuracy (%) Stream Forgetting (%) Total time (s)
+CoP2L-Dino (ViT) $70.56\pm 0.99$ $21.15\pm 2.10$ $820.66\pm 176.27$
+Replay-Dino (ViT) $69.21\pm 0.26$ $29.52\pm 0.25$ $457.46\pm 2.07$
+iCaRL-Dino (ViT) $62.36\pm 1.46$ $15.88\pm 0.84$ $1359.13\pm 4.58$
+GDumb-Dino (ViT) $72.35\pm 0.30$ $11.77\pm 0.40$ $190.80\pm 0.42$
+CoP2L-Dino (ResNet50) $46.74\pm 1.63$ $22.54\pm 0.92$ $1771.12\pm 757.50$
+Replay-Dino (ResNet50) $49.44\pm 0.30$ $44.96\pm 0.36$ $293.25\pm 1.77$
+iCaRL-Dino (ResNet50) $46.47\pm 0.33$ $6.74\pm 0.31$ $858.11\pm 120.41$
+GDumb-Dino (ResNet50) $43.60\pm 0.38$ $19.93\pm 0.56$ $126.74\pm 0.46$
+TinyImageNet (40 tasks) Class-Incremental
+Methods Stream Accuracy (%) Stream Forgetting (%) Total time (s)
+CoP2L-Dino (ViT) $51.59\pm 0.24$ $36.32\pm 0.29$ $1451.93\pm 4.71$
+Replay-Dino (ViT) $47.49\pm 0.47$ $48.39\pm 0.49$ $1128.97\pm 16.95$
+iCaRL-Dino (ViT) $21.06\pm 1.11$ $34.49\pm 9.37$ $3919.23\pm 2170.08$
+GDumb-Dino (ViT) $53.93\pm 0.34$ $15.11\pm 0.19$ $471.23\pm 17.85$
+CoP2L-Dino (ResNet50) $33.30\pm 0.32$ $35.93\pm 0.27$ $3671.69\pm 29.19$
+Replay-Dino (ResNet50) $28.38\pm 0.42$ $63.36\pm 0.44$ $882.12\pm 180.18$
+iCaRL-Dino (ResNet50) $33.89\pm 0.47$ $10.42\pm 0.44$ $1905.26\pm 295.18$
+GDumb-Dino (ResNet50) $24.29\pm 0.58$ $21.29\pm 0.56$ $332.22\pm 0.98$
+CIFAR10 (5 tasks) Task-Incremental
+Methods Stream Accuracy (%) Stream Forgetting (%) Total time (s)
+CoP2L-Dino (ViT) $99.04\pm 0.05$ $-0.00\pm 0.10$ $64.51\pm 0.68$
+Replay-Dino (ViT) $99.29\pm 0.04$ $-0.01\pm 0.03$ $217.83\pm 0.76$
+LaMAML-Dino (ViT) $99.25\pm 0.06$ $0.05\pm 0.07$ $344.65\pm 1.46$
+LwF-Dino (ViT) $97.89\pm 0.41$ $1.69\pm 0.48$ $165.41\pm 0.75$
+CoP2L-Dino (ResNet50) $96.22\pm 0.18$ $0.56\pm 0.28$ $172.51\pm 4.42$
+Replay-Dino (ResNet50) $96.88\pm 0.06$ $0.42\pm 0.13$ $144.57\pm 0.51$
+LaMAML-Dino (ResNet50) $96.65\pm 0.21$ $0.73\pm 0.21$ $310.51\pm 1.64$
+LwF-Dino (ResNet50) $95.80\pm 0.15$ $0.33\pm 0.20$ $97.09\pm 0.21$
+CIFAR100 (10 tasks) Task-Incremental
+Methods Stream Accuracy (%) Stream Forgetting (%) Total time (s)
+CoP2L-Dino (ViT) $95.03\pm 0.08$ $0.90\pm 0.11$ $290.95\pm 30.98$
+Replay-Dino (ViT) $96.43\pm 0.14$ $0.34\pm 0.17$ $235.81\pm 1.10$
+LaMAML-Dino (ViT) $95.58\pm 0.16$ $0.58\pm 0.12$ $397.78\pm 3.40$
+LwF-Dino (ViT) $95.09\pm 0.10$ $0.98\pm 0.10$ $173.57\pm 0.63$
+CoP2L-Dino (ResNet50) $86.78\pm 0.16$ $2.01\pm 0.17$ $435.10\pm 7.74$
+Replay-Dino (ResNet50) $87.61\pm 0.05$ $0.86\pm 0.16$ $157.05\pm 0.14$
+LaMAML-Dino (ResNet50) $83.57\pm 0.43$ $6.18\pm 0.53$ $357.24\pm 4.58$
+LwF-Dino (ResNet50) $79.39\pm 0.13$ $0.47\pm 0.11$ $101.27\pm 0.48$
+CIFAR100 (20 tasks) Task-Incremental
+Methods Stream Accuracy (%) Stream Forgetting (%) Total time (s)
+CoP2L-Dino (ViT) $97.18\pm 0.13$ $0.56\pm 0.18$ $226.65\pm 21.43$
+Replay-Dino (ViT) $98.09\pm 0.09$ $0.25\pm 0.12$ $250.78\pm 1.45$
+LaMAML-Dino (ViT) $97.84\pm 0.10$ $0.26\pm 0.14$ $448.43\pm 2.81$
+LwF-Dino (ViT) $95.74\pm 0.19$ $2.15\pm 0.21$ $180.36\pm 0.98$
+CoP2L-Dino (ResNet50) $90.89\pm 0.22$ $2.27\pm 0.28$ $278.67\pm 6.20$
+Replay-Dino (ResNet50) $92.89\pm 0.18$ $0.47\pm 0.24$ $167.65\pm 1.26$
+LaMAML-Dino (ResNet50) $89.89\pm 0.75$ $4.12\pm 0.78$ $407.84\pm 7.38$
+LwF-Dino (ResNet50) $87.73\pm 0.18$ $0.33\pm 0.23$ $105.46\pm 0.49$
+TinyImageNet (40 tasks) Task-Incremental
+Methods Stream Accuracy (%) Stream Forgetting (%) Total time (s)
+CoP2L-Dino (ViT) $93.81\pm 0.26$ $1.50\pm 0.28$ $1039.14\pm 146.21$
+Replay-Dino (ViT) $95.66\pm 0.08$ $0.50\pm 0.09$ $632.18\pm 2.97$
+LaMAML-Dino (ViT) $94.82\pm 0.23$ $0.79\pm 0.19$ $1527.24\pm 891.43$
+LwF-Dino (ViT) $85.18\pm 0.76$ $9.72\pm 0.71$ $430.63\pm 2.14$
+CoP2L-Dino (ResNet50) $88.57\pm 0.17$ $2.46\pm 0.19$ $890.94\pm 3.44$
+Replay-Dino (ResNet50) $90.03\pm 0.26$ $1.26\pm 0.27$ $514.42\pm 102.82$
+LaMAML-Dino (ResNet50) $87.20\pm 0.34$ $4.53\pm 0.28$ $1268.50\pm 551.93$
+LwF-Dino (ResNet50) $85.03\pm 0.17$ $0.89\pm 0.21$ $284.79\pm 5.67$
+ImageNette (5 tasks) Task-Incremental
+Methods Stream Accuracy (%) Stream Forgetting (%) Total time (s)
+CoP2L-Dino (ViT) $98.22\pm 0.19$ $-0.43\pm 0.12$ $118.71\pm 24.35$
+ImageWoof (5 tasks) Task-Incremental
+Methods Stream Accuracy (%) Stream Forgetting (%) Total time (s)
+CoP2L-Dino (ViT) $93.34\pm 0.30$ $-0.65\pm 0.40$ $246.77\pm 13.43$
+Permuted-MNIST (10 tasks) Domain-Incremental
+Methods Stream Accuracy (%) Stream Forgetting (%) Total time (s)
+CoP2L $91.61\pm 0.39$ $6.69\pm 0.40$ $1381.06\pm 1232.99$
+Replay $91.28\pm 0.18$ $6.75\pm 0.21$ $1128.99\pm 49.82$
+LFL $82.76\pm 1.23$ $10.15\pm 1.36$ $92.86\pm 0.24$
+EWC $82.53\pm 0.82$ $6.50\pm 0.88$ $340.16\pm 3.81$
+SI $85.85\pm 0.61$ $11.69\pm 0.67$ $319.95\pm 1.38$
+Rotated-MNIST (10 tasks) Domain-Incremental
+Methods Stream Accuracy (%) Stream Forgetting (%) Total time (s)
+CoP2L $89.40\pm 1.00$ $8.71\pm 1.04$ $4970.86\pm 1452.34$
+Replay $88.75\pm 0.67$ $9.53\pm 0.72$ $3894.62\pm 75.11$
+LFL $53.09\pm 7.06$ $44.18\pm 7.93$ $387.61\pm 160.54$
+EWC $55.59\pm 5.57$ $36.58\pm 6.27$ $1133.19\pm 18.10$
+SI $59.90\pm 6.48$ $41.38\pm 7.18$ $1043.86\pm 7.28$
+MNIST (5 tasks) Class-Incremental
+Methods Stream Accuracy (%) Stream Forgetting (%)
+CoP2L (MLP) $95.12\pm 0.18$ $3.35\pm 0.24$
+Replay (MLP) $94.48\pm 0.28$ $6.08\pm 0.32$
+GDumb (MLP) $91.70\pm 0.31$ $3.84\pm 0.29$
+CoP2L (CNN) $96.61\pm 0.07$ $2.16\pm 0.28$
+Replay (CNN) $96.93\pm 0.29$ $3.31\pm 0.38$
+GDumb (CNN) $94.41\pm 0.13$ $2.97\pm 0.46$
+Fashion-MNIST (5 tasks) Class-Incremental
+Methods Stream Accuracy (%) Stream Forgetting (%)
+CoP2L (MLP) $84.23\pm 0.27$ $12.10\pm 0.66$
+Replay (MLP) $83.99\pm 0.52$ $15.92\pm 0.42$
+GDumb (MLP) $81.76\pm 0.36$ $10.44\pm 0.50$
+CoP2L (CNN) $87.35\pm 0.15$ $11.12\pm 0.36$
+Replay (CNN) $87.59\pm 0.11$ $13.32\pm 0.29$
+GDumb (CNN) $84.98\pm 0.21$ $8.78\pm 0.41$
+EMNIST (13 tasks) Class-Incremental
+Methods Stream Accuracy (%) Stream Forgetting (%)
+CoP2L (MLP) $78.03\pm 0.17$ $16.67\pm 0.31$
+Replay (MLP) $77.22\pm 0.44$ $22.08\pm 0.50$
+GDumb (MLP) $62.64\pm 0.20$ $15.96\pm 0.58$
+CoP2L (CNN) $87.90\pm 0.41$ $8.38\pm 0.54$
+Replay (CNN) $87.70\pm 0.50$ $12.29\pm 0.49$
+GDumb (CNN) $85.05\pm 0.27$ $7.50\pm 0.61$
+Appendix G Training, architecture details and hyperparameter settings for the experiments
 […]
-where $\text{Accuracy}_{T,t}^{o}$ denotes the accuracy obtained at task $t$ after finishing training until task $T$, and the $o$ index denotes the task order.
-We have used $O=5$ different task orderings.
-We present the results of this experiment in Figure 4.
-We observe that on both MNIST and Fashion-MNIST datasets CoP2L is able outperform with more tasks.
-We also observe that the variability of the results decrase more significantly for CoP2L compared to replay.
-5 Conclusions
-In this paper, we have proposed CoP2L, an algorithm rooted in the sample compression theory.
-To the best of our knowledge, this is the first attempt in the literature to employ sample compression theory within the continual learning context.
-We provided sample compression bounds for CoP2L and verified empirically that they were non-vacuous.
-We furthermore showed that on three different datasets, our approach is able to outperform experience replay, which is known to be a strong continual learning baseline.
-Acknowledgements
-We also wish to thank Benjamin Leblanc for his help proof-reading the manuscript.
-Mathieu Bazinet is supported by a FRQNT B2X scholarship (343192).
-Pascal Germain is supported by the NSERC Discovery grant RGPIN-2020-07223.
-Cem Subakan is supported by NSERC Discovery grant RGPIN-2023-05759.
-This research was also enabled in part by support provided by the Digital Research Alliance of Canada.
-Appendix A Tighter sample compression $\mathrm{kl}$ bound
-In this section, we prove a tighter $\mathrm{kl}$ bound that is not derived from Theorem 2.1 but proved in a very similar way.
-To do so, we start by providing the Chernoff test-set bound (Langford, 2005) for losses in $[0,1]$ (Foong et al., 2022).
-Theorem A.1 ((Foong et al., 2022)).
-Let $X_{1},\ldots,X_{n}$ be i.i.d. random variables with $X_{i}\in[0,1]$ and $\operatorname*{\mathbb{E}}[X_{i}]=p$. Then, for any $\delta\in(0,1]$, with probability at least $1-\delta$
-$p\leq\mathrm{kl}^{-1}\quantity(\frac{1}{n}\sum_{i=1}^{n}X_{i},\frac{1}{n}\log\frac{1}{\delta}).$
-We now prove a tighter $\mathrm{kl}$ sample compression bound.
-As our algorithm CoP2L needs two compression sets, we consider a second compression set $\operatorname{\mathbf{j}}$.
-We define a conditional probability distribution $P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})$ that incorporates the knowledge that a vector $\operatorname{\mathbf{i}}$ was already drawn from $\operatorname{\mathscr{P}}(n)$ and that $\operatorname{\mathbf{i}}\cap\operatorname{\mathbf{j}}=\emptyset$.
-Thus, we have $\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})\leq 1.$ If the choice of $\operatorname{\mathbf{j}}$ isn’t conditional to the choice of $\operatorname{\mathbf{i}}$, we simply have $P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})=P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}})$.
-Theorem A.2.
-For any distribution $\operatorname{\mathcal{D}}$ over $\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}$, for any family of set of messages $\{\mathscr{M}(\operatorname{\mathbf{j}})\,|\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)\}$, for any deterministic reconstruction function $\operatorname{\mathscr{R}}$ , for any loss $\ell:\Theta\times\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}\to[0,1]$ and for any $\delta\in(0,1]$, with probability at least $1-\delta$ over the draw of $S\sim\operatorname{\mathcal{D}}^{n}$, we have
-$\forall\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n),\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n),\mu\in\mathscr{M}(\operatorname{\mathbf{j}}):$
-$\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta})$
-Proof.
-We prove the complement of the expression in Theorem A.2.
-We choose $\delta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}=P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta$.
-$\operatorname*{\mathbb{P}}_{S\sim\operatorname{\mathcal{D}}^{n}}\quantity(\exists\operatorname{\mathbf{i}},\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n),\mu\in\mathscr{M}(\operatorname{\mathbf{j}}):\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))>\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{\delta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}}))$
-$\leq\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\operatorname*{\mathbb{P}}_{S\sim\operatorname{\mathcal{D}}^{n}}\quantity(\exists\mu\in\mathscr{M}(\operatorname{\mathbf{j}}):\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))>\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{\delta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}}))$ (13)
-$\leq\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\sum_{\mu\in\mathscr{M}(\operatorname{\mathbf{j}})}\operatorname*{\mathbb{P}}_{S\sim\operatorname{\mathcal{D}}^{n}}\quantity(\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))>\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{\delta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}}))$ (14)
-$\leq\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\sum_{\mu\in\mathscr{M}(\operatorname{\mathbf{j}})}\operatorname*{\mathbb{E}}_{S^{\operatorname{\mathbf{i}}}\sim\operatorname{\mathcal{D}}^{\operatorname{|\mathbf{i}|}}}\operatorname*{\mathbb{E}}_{S^{\operatorname{\mathbf{j}}}\sim\operatorname{\mathcal{D}}^{|\operatorname{\mathbf{j}}|}}\operatorname*{\mathbb{P}}_{S^{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\sim\operatorname{\mathcal{D}}^{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}}\quantity(\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))>\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{\delta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}}))$ (15)
-$\leq\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\sum_{\mu\in\mathscr{M}(\operatorname{\mathbf{j}})}\operatorname*{\mathbb{E}}_{S^{\operatorname{\mathbf{i}}}\sim\operatorname{\mathcal{D}}^{\operatorname{|\mathbf{i}|}}}\operatorname*{\mathbb{E}}_{S^{\operatorname{\mathbf{j}}}\sim\operatorname{\mathcal{D}}^{|\operatorname{\mathbf{j}}|}}\delta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}$ (16)
-$\leq\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\sum_{\mu\in\mathscr{M}(\operatorname{\mathbf{j}})}\delta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}$
-$\leq\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\sum_{\mu\in\mathscr{M}(\operatorname{\mathbf{j}})}P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta$
-$\leq\delta.$
-We use the union bound in Equation 13 and Equation 14.
-In Equation 15, we use the i.i.d.
-assumption.
-In Equation 16 we use the Chernoff Test-set bound of Theorem A.1 with $p=\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))=\operatorname*{\mathbb{E}}_{(\boldsymbol{x},y)\sim\operatorname{\mathcal{D}}}\ell(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu),\boldsymbol{x},y)$ and $X_{i}=\ell(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu),\boldsymbol{x}_{i},y_{i})\forall i\in\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}$.
-Finally, the last line is derived from the definition of $P_{\operatorname{\mathscr{P}}(n)}$ and $P_{\mathscr{M}(\operatorname{\mathbf{j}})}$.
-∎
-Appendix B Proof of Theorem 3.1
-We now prove Theorem 3.1.
-See 3.1
-Proof.
-Let us choose $t\in[1,T]$.
-Let us sample $S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}$.
-We denote the reconstruction function of CoP2L as $\operatorname{\mathscr{R}}_{1:T}(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)=\operatorname{\mathscr{R}}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu;S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}).$
-Then, we have :
-$\operatorname*{\mathbb{P}}_{S^{t}}\Bigg{(}\begin{split}&\forall\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n_{t}),\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n_{t}),\mu\in\mathscr{M}(\operatorname{\mathbf{j}}):\\ &\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n_{t}-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta})\end{split}\Biggr{)}$
-As all the datasets (except $S^{t}$) are sampled beforehand, we can define $\operatorname{\mathscr{R}}_{1:T}(\cdot)=\operatorname{\mathscr{R}}\quantity(\cdot;S^{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T})$ before drawing $S^{t}$.
-Thus, the reconstruction function is only a function of the dataset $S^{t}$, which is the setting of the result of Theorem A.2.
-We can then lower bound this probability by $1-\delta$.
-We finish the proof by applying the bound to all datasets.
-With
-$\epsilon(n_{k},\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu,\delta)=\frac{1}{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta},$
-we have
-$\operatorname*{\mathbb{P}}_{S_{1},\ldots,S_{T}}\Bigg{(}$ $\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)),\epsilon(n_{t},\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu,\delta))\Biggr{)}$
-$=\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{T}}\operatorname{\mathbb{I}}\Bigg{(}$ $\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)),\epsilon(n_{t},\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu,\delta))\Biggr{)}$
-$=\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}}\operatorname*{\mathbb{E}}_{S_{t}}\operatorname{\mathbb{I}}\Bigg{(}$ $\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)),\epsilon(n_{t},\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu,\delta))\Biggr{)}$
-$=\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}}\operatorname*{\mathbb{P}}_{S_{t}}\Bigg{(}$ $\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)),\epsilon(n_{t},\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu,\delta))\Biggr{)}$
-$\geq\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}}1-\delta$
-$\geq 1-\delta$
-We use a union bound argument to add the $\forall k\in[1,T]$, leading to the term $\frac{\delta}{T}$.
-As this theorem holds specifically for CoP2L, we specify the distributions $P_{\operatorname{\mathscr{P}}(n_{t})}$ and $P_{\mathscr{M}(\operatorname{\mathbf{j}})}$.
-Following the work of Marchand et al.
-(2003), which used sample compression bounds with three compression sets, with $\zeta(k)=\frac{6}{\pi}(k+1)^{-1}$, we choose $P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{i}})=\matrixquantity(n_{t}\\ \operatorname{|\mathbf{i}|})^{-1}\zeta(\operatorname{|\mathbf{i}|})$ and $P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})=\matrixquantity(n_{t}-\operatorname{|\mathbf{i}|}\\ |\operatorname{\mathbf{j}}|)^{-1}\zeta(|\operatorname{\mathbf{j}}|)$.
-Finally, we split the message $\mu$ into two messages $\mu_{1}$ and $\mu_{2}$.
-The first message is defined using the alphabet $\Sigma_{1}=\{2,\ldots,T\}$.
-This message indicates the task number where a datapoint is sampled out of the buffer.
-Thus, the size of $\Sigma_{1}$ is $T-1$ and the probability of a symbol is $\frac{1}{T-1}$.
-For any vector $\operatorname{\mathbf{j}}$, we have a sequence of length $|\operatorname{\mathbf{j}}|$, and thus the probability of choosing each sequence is $\quantity(\frac{1}{T-1})^{|\operatorname{\mathbf{j}}|}$.
-The second message $\mu_{2}$ is defined using the alphabet $\Sigma_{2}=\{1,n_{t}\}$.
-We choose to use $\zeta(k)=\frac{6}{\pi^{2}}(k+1)^{-2}$ as the probability distribution over each character.
-We know that for any $N\geq 0$, $\sum_{k=1}^{N}\zeta(k)\leq 1$.
-Thus, we have the probability of a sequence $\mu_{2}=\mu_{2}^{1}\ldots\nu_{2}^{T}$ is $\prod_{i=1}^{T}\zeta(\mu_{2}^{i})$.
-We define $\mathscr{M}(\operatorname{\mathbf{j}},T)$ the set of messages such that the sequence of symbols from $\Sigma_{1}$ is of length $|\operatorname{\mathbf{j}}|$ and the sequence of symbols from $\Sigma_{2}$ is of length $T$.
-∎
-Appendix C Bound plots with different buffer sizes
-In this section, we introduce the empirical bound estimates over tasks on MNIST, Fashion-MNIST and EMNIST datasets.
-In figure 5, we present results for different replay buffer sizes (1000, 2000, 3000, 4000, 5000).
-Figure 5: CoP2L Bounds with respect to tasks on the MNIST, Fashion-MNIST and EMNIST datasets
+In experiments using a Dino backbone (both ResNet50 and ViT backbones), we upsampled the images to 64x64, since Dino was pretrained on 224x224 images.
+We observed better performance from the pretrained feature extractor when the input image size is closer to the original training image size of Dino.
+We present below the two custom CNN architectures derive from the SimpleCNN architecture (from the Avalanche toolkit).
+All other architectures used for the experiments are available in the Avalanche toolkit.
+The specific CNN architecture used for MNIST experiments is defined as follows:
+MnistCnn(
+  (cnn_layers): Sequential(
+    (0): Conv2d(1, 32, kernel_size=(3, 3), stride=(1, 1))
+    (1): ReLU()
+    (2): Dropout2d(p=0.2, inplace=False)
+    (3): Conv2d(32, 64, kernel_size=(3, 3), stride=(1, 1))
+    (4): ReLU()
+    (5): Dropout2d(p=0.2, inplace=False)
+    (6): MaxPool2d(kernel_size=2, stride=2, padding=0, dilation=1, ceil_mode=False)
+  (linear_layers): Sequential(
+    (0): Linear(in_features=9216, out_features=128, bias=True)
+    (1): ReLU()
+    (2): Dropout(p=0.2, inplace=False)
+    (3): Linear(in_features=128, out_features=10, bias=True)
+The specific CNN architecture used for FMNIST and EMNIST experiments is defined as follows:
+EMNISTFashionMnistCnn(
+  (cnn_layers): Sequential(
+    (0): Conv2d(1, 32, kernel_size=(3, 3), stride=(1, 1))
+    (1): ReLU()
+    (2): BatchNorm2d(32, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
+    (3): Dropout2d(p=0.3, inplace=False)
+    (4): Conv2d(32, 64, kernel_size=(3, 3), stride=(1, 1))
+    (5): ReLU()
+    (6): BatchNorm2d(64, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
+    (7): MaxPool2d(kernel_size=2, stride=2, padding=0, dilation=1, ceil_mode=False)
+    (8): Dropout2d(p=0.3, inplace=False)
+    (9): Conv2d(64, 128, kernel_size=(3, 3), stride=(1, 1))
+    (10): ReLU()
+    (11): BatchNorm2d(128, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
+    (12): MaxPool2d(kernel_size=2, stride=2, padding=0, dilation=1, ceil_mode=False)
+    (13): Dropout2d(p=0.3, inplace=False)
+  (linear_layers): Sequential(
+    (0): Linear(in_features=3200, out_features=256, bias=True)
+    (1): ReLU()
+    (2): Dropout(p=0.3, inplace=False)
+    (3): Linear(in_features=256, out_features=10, bias=True)
+Method ${k}$ ${\omega}$ Epochs Batch size Learning rate
+CoP2L 8 15.0 10 256 0.001
+Replay NA NA 20 128 0.01
+GDumb NA NA 20 128 0.01
+Table 3: Hyperparameter setup for MNIST, FMNIST and EMNIST using a MLP architecture. NA signifies Not Applicable.
+Method ${k}$ ${\omega}$ Epochs Batch size Learning rate
+CoP2L (MNIST & FMNIST) 48 25.0 25 256 0.0009
+CoP2L (EMNIST) 105 20.0 15 130 0.0006
+Replay NA NA 20 128 0.01
+GDumb NA NA 20 128 0.01
+Table 4: Hyperparameter setup for MNIST, FMNIST and EMNIST using a CNN architecture. NA signifies Not Applicable.
+Method ${k}$ ${\omega}$ Epochs Batch size Learning rate
+CoP2L-Dino (ViT) 4 25.0 2 256 0.001
+Replay-Dino (ViT) NA NA 20 128 0.01
+iCaRL-Dino (ViT) NA NA 10 128 0.01
+GDumb-Dino (ViT) NA NA 20 128 0.01
+CoP2L-Dino (RN50) 16 25.0 2 256 0.001
+Replay-Dino (RN50) NA NA 20 128 0.01
+iCaRL-Dino (RN50) NA NA 30 128 0.01
+GDumb-Dino (RN50) NA NA 20 128 0.01
+Table 5: Hyperparameter setup for Class-Incremental setup. NA signifies Not Applicable.
+Method ${k}$ ${\omega}$ Epochs Batch size Learning rate
+CoP2L-Dino (ViT) 4 1.0 2 256 0.001
+Replay-Dino (ViT) NA NA 10 128 0.01
+LaMAML-Dino (ViT) NA NA 10 10 0.01
+LwF-Dino (ViT) NA NA 10 200 0.001
+CoP2L-Dino (RN50) 16 1.0 2 256 0.001
+Replay-Dino (RN50) NA NA 10 128 0.01
+LaMAML-Dino (RN50) NA NA 10 10 0.01
+LwF-Dino (RN50) NA NA 10 200 0.001
+Table 6: Hyperparameter setup for Task-Incremental setup. NA signifies Not Applicable.
+Method ${k}$ ${\omega}$ Epochs Batch size Learning rate
+CoP2L 16 15.0 1 256 0.001
+Replay NA NA 20 128 0.01
+LFL NA NA 3 128 0.01
+EWC NA NA 10 256 0.001
+SI NA NA 10 256 0.001
+Table 7: Hyperparameter setup for Domain-Incremental setup. NA signifies Not Applicable.
+Appendix H Dataset Details
+H.1 Dataset summary
+Datasets used for class incremental learning experiments:
+  • CIFAR10, 5 tasks of 2 classes each (total of 10 classes)
+  • CIFAR100 (10 tasks), 10 tasks of 10 classes each (total of 100 classes)
+  • CIFAR100 (20 tasks), 20 tasks of 5 classes each (total of 100 classes)
+  • TinyImageNet (40 tasks), 40 tasks of 5 classes each (total of 200 classes)
+  • ImageNette (5 tasks), 5 tasks of 2 classes each (total of 10 classes)
+  • ImageWoof (5 tasks), 5 tasks of 2 classes each (total of 10 classes)

### v4

@@
 Abstract
-Continual learning algorithms aim to learn from a sequence of tasks, making the training distribution non-stationary.
-The majority of existing continual learning approaches in the literature rely on heuristics and do not provide learning guarantees.
-In this paper, we present a new method called Continual Pick-to-Learn (CoP2L), which is able to retain the most representative samples for each task in an efficient way.
-CoP2L combines the Pick-to-Learn algorithm (rooted in the sample compression theory) and the experience replay continual learning scheme.
-This allows us to provide non-vacuous upper bounds on the generalization loss of the learned predictors, numerically computable after each task.
-We empirically evaluate our approach on several standard continual learning benchmarks across Class-Incremental, Task-Incremental, and Domain-Incremental settings.
-Our results show that CoP2L is highly competitive across all setups, often outperforming existing baselines, and significantly mitigating catastrophic forgetting compared to vanilla experience replay in the Class-Incremental setting.
-It is possible to leverage the bounds provided by CoP2L in practical scenarios to certify the predictor reliability on previously learned tasks, in order to improve the trustworthiness of the continual learning algorithm.
+Continual learning algorithms aim to learn from a sequence of tasks.
+In order to avoid catastrophic forgetting, most existing approaches rely on heuristics and do not provide computable learning guarantees.
+In this paper, we introduce Continual Pick-to-Learn (CoP2L), a method grounded in sample compression theory that retains representative samples for each task in a principled and efficient way.
+This allows us to derive non-vacuous, numerically computable upper bounds on the generalization loss of the learned predictors after each task.
+We evaluate CoP2L on standard continual learning benchmarks under Class-Incremental and Task-Incremental settings, showing that it effectively mitigates catastrophic forgetting.
+It turns out that CoP2L is empirically competitive with baseline methods while certifying predictor reliability in continual learning with a non-vacuous bound.
 1 INTRODUCTION
 A common assumption in traditional machine learning is that the underlying data distribution does not evolve with time.
-In Continual Learning de2021continual ; wang2023comprehensive , the goal is to develop machine learning algorithms that are able to learn under settings where this assumption is replaced by a set-up where the model is trained on an evolving training data distribution, in such a way that samples are revealed one task at a time.
-However, when neural networks are trained on evolving data distributions, networks tend to forget the training from earlier tasks.
-This fact is known as the catastrophic forgetting McCloskey1989CatastrophicII ; forgetting_french ; goodfellow2013empirical .
-In order to cope with forgetting, various methodologies have been developed, such as regularization based approaches ewc_kirkpatrick ; zenke2017continual ; cl_li ; mas , architectural approaches pnn_rusu ; mallya2018piggyback ; expert_aljundi ; packnet_mallya ; l2p_wang , or rehearsal based approaches rolnick2019experience ; chaudhry2019efficient ; replay_shin ; icarl_rebuffi ; der .
-It has been shown on various applications that rehearsal-based approaches typically perform very competitively compared to the other methods in terms of performance and forgetting DellaLibera2023CLMASRAC ; Mai2020BatchlevelER ; chaudhry2019efficient ; aljundi2019onlinecontinuallearningmaximally .
-However, the selection of the replay buffer typically remains a heuristic, and precise upper bounds for the generalization error are not available.
-Our proposed continual learning scheme builds upon the Pick-to-Learn algorithm (P2L) of Paccagnan et al.
-paccagnan_pick_learn_2023 .
-P2L is a meta-algorithm that selects a compression set, which amounts to expressing the dataset in terms of ‘core’ examples.
-This meta-algorithm was developed specifically with sample compression theory in mind floyd1995sample ; campi2023compression .
-By finding a small subset of the data such that a predictor learned on this data achieves low error on the whole training set, Pick-to-Learn enables us to compute tight upper bounds for the generalization error of the learned predictor.
-This ability of a learning algorithm to simultaneously output a predictor and its risk certificate has been coined as self-certified learning in the literature freund98selfbounding ; perezortiz2021tighter ; marks2025pick2learn-gp .
-This certification is arguably an asset to increase the trustworthiness towards machine learning systems.
-In this paper, we propose Continual Pick-to-Learn (CoP2L), an algorithm that leverages sample compression theory to intelligently select the training data from earlier tasks to mitigate forgetting.
+In Continual Learning (De Lange et al., 2021; Wang et al., 2024), the goal is to develop machine learning algorithms that are able to learn under settings where this assumption is replaced by a setup where the model is trained on an evolving training data distribution, in such a way that samples are revealed one task at a time.
+However, neural networks trained on evolving data distributions tend to forget earlier tasks, a phenomenon known as catastrophic forgetting (McCloskey and Cohen, 1989; French, 1999; Goodfellow et al., 2014).
+In order to cope with forgetting, various methodologies have been developed, such as regularization-based approaches (Kirkpatrick et al., 2017; Zenke et al., 2017; Li and Hoiem, 2017; Aljundi et al., 2018), architectural-based approaches (Rusu et al., 2016; Mallya et al., 2018; Aljundi et al., 2017; Mallya and Lazebnik, 2018; Wang et al., 2022; He et al., 2025; Gao et al., 2023; Zhou et al., 2025, 2024) or rehearsal-based approaches (Rolnick et al., 2019; Chaudhry et al., 2019; Shin et al., 2017; Rebuffi et al., 2017a; Buzzega et al., 2020).
+(a) Class-incremental setting
+(b) Task-incremental setting
+Figure 1: Numerical values of the proposed generalization bounds for continual learning over 10 tasks on CIFAR100, using ViT and ResNet50 backbones. The bounds hold simultaneously for all tasks.
+We develop a novel sample compression theory for continual learning.
+Being agnostic to the architecture, our theoretical results are well-suited for rehearsal-based approaches, which are commonly used in the continual learning literature for practical applications (Libera et al., 2023; Mai et al., 2020; Chaudhry et al., 2019; Aljundi et al., 2019).
+We then propose a theoretically motivated algorithm that uses the replay buffer only when necessary to mitigate the forgetting.
+These new results pave the way for continual learning approaches based on theoretical guarantees.
+More precisely, our proposed continual learning scheme builds upon the Pick-to-Learn meta-algorithm (P2L) of Paccagnan et al.
+(2024).
+This meta-algorithm was developed specifically with sample compression theory in mind (Floyd and Warmuth, 1995; Campi and Garatti, 2023).
+By finding a small subset of the data such that a predictor learned on this subset achieves low error on the whole training set, P2L enables computing tight upper bounds for the generalization error of the learned predictor.
+This ability of a learning algorithm to simultaneously output a predictor and its risk certificate has been coined as self-certified learning (Freund, 1998; Pérez-Ortiz et al., 2021; Marks and Paccagnan, 2025).
+This certification is arguably an asset to increase the trustworthiness of machine learning systems.
+The Continual Pick-to-Learn (CoP2L) [^1] Our code is available at https://anonymous.4open.science/r/CoP2L_paper_code-0058/ algorithm that we propose leverages sample compression theory to intelligently select the training data from earlier tasks to mitigate forgetting.
 We derive high-confidence upper bounds on the generalization error for each task simultaneously, estimated directly from the training set.
-We also empirically show that CoP2L mitigates the forgetting significantly and outperforms the standard replay baseline significantly in the small replay buffer size settings.
+We also empirically show that CoP2L substantially mitigates forgetting and is comparable with several standard baselines.
 Our contributions can be summarized as follows:
 (1) We propose the algorithm Continual Pick-to-Learn (CoP2L), that integrates the sample compression theory within the continual learning setup.
-To the best of our knowledge, we are the first to integrate the theoretical results from sample compression to continual learning.
+To the best of our knowledge, we are the first to integrate the theoretical results from sample compression to rehearsal-based continual learning.
 (2) CoP2L is able to provide high confidence non-trivial upper bounds for the generalization error by means of the sample compression theory.
 We experimentally showcase that the bounds are numerically computable, have non-trivial values and follow the general error trends observed on the test set.
-Therefore, they can actually be used as risk certificates on the model behavior on the learned tasks, improving the trustworthiness of the continually learnt model.
-(3) We experimentally show with a wide range of experiments that CoP2L significantly mitigates forgetting, while also being able to obtain better or comparable performance compared with several strong continual learning baselines in most cases.
+Therefore, they can actually be used as risk certificates for the model’s behaviour on the learned tasks, improving the trustworthiness of the continually learnt model.
+(3) We experimentally show with a wide range of experiments that CoP2L significantly mitigates forgetting, while also being able to obtain comparable performance compared with several strong continual learning baselines.
 1.1 Related Work
@@
 In the continual learning literature, various practical approaches have been developed to mitigate forgetting.
-Broadly, the approaches could be divided into three categories, which include regularization approaches ewc_kirkpatrick ; zenke2017continual ; cl_li ; mas , architecture-based approaches pnn_rusu ; mallya2018piggyback ; expert_aljundi ; packnet_mallya ; l2p_wang and rehearsal based approaches rolnick2019experience ; chaudhry2019efficient ; replay_shin ; icarl_rebuffi ; der .
+Broadly, the approaches could be divided into three categories: 1) regularization-based approaches (Kirkpatrick et al., 2017; Zenke et al., 2017; Li and Hoiem, 2017; Aljundi et al., 2018), 2) architecture-based approaches (Rusu et al., 2016; Mallya et al., 2018; Aljundi et al., 2017; Mallya and Lazebnik, 2018; Wang et al., 2022; He et al., 2025; Gao et al., 2023; Zhou et al., 2025, 2024) and 3) rehearsal-based approaches (Rolnick et al., 2019; Chaudhry et al., 2019; Shin et al., 2017; Rebuffi et al., 2017a; Buzzega et al., 2020).
+Most rehearsal-based approaches use a replay buffer that contains a small subset of the dataset over which the model continues to train, even when learning over new tasks.
+The buffer helps to mitigate the forgetting of a model on previous tasks.
+However, the effect of the buffer depends on the quality of the data chosen in the buffer.
+A naive approach is to sample randomly the datapoints (Mai et al., 2020; Libera et al., 2023).
+Non-naive methods include methods such as iCarl (Rebuffi et al., 2017a), Dark Experience Replay (Buzzega et al., 2020), Gradient Episodic Memory (Lopez-Paz and Ranzato, 2017), Example Forgetting (Benkő, 2024), Global Pseudo-Task simulation (Liu et al., 2022) and Coreset-based methods (e.g., Hao et al., 2023; Tong et al., 2025; Borsos et al., 2020; Yoon et al., 2022).
 Continual learning generalization bounds.
-In recent years, generalization bounds were proposed to study the asymptotic behavior of various continual learning schemes, for example using linear models Evron2022HowCC ; Goldfarb2022AnalysisOC ; pmlr-v202-lin23f , Neural Tangent Kernel (NTK) models trained with Orthogonal Gradient Descent AbbanaBennani2020GeneralisationGF ; Doan2020ATA ; farajtabar2019orthogonalgradientdescentcontinual , regularization-based methods yin2021optimizationgeneralizationregularizationbasedcontinual , or PackNet packnet_mallya inspired continual representation learning Li2022ProvableAE .
-Learning algorithms with convergence guarantees are were also proposed under data assumptions, such that all target tasks lie in a same space obtainable from a (linear or non-linear) combination of some input features cao2022lifelong .
-Different from all the listed works above, in this paper, we provide a generalization bound that is computed from the training set.
-Risk certificates on the test error are provided simultaneously for each task.
-Furthermore, also unlike the approaches above, our non-asymptotic bound is applicable to any neural network architecture, under minimal data assumptions, and does not have limitations on the continual learning setup that is employed.
+In recent years, generalization bounds were proposed to study the asymptotic behavior of various specific models, e.g., for linear models (Evron et al., 2022; Goldfarb and Hand, 2023; Lin et al., 2023; Banayeeanzade et al., 2025); for Neural Tangent Kernel models trained with Orthogonal Gradient Descent (Bennani and Sugiyama, 2020; Doan et al., 2021; Farajtabar et al., 2020); for regularization-based methods (Yin et al., 2020); or for a PackNet-inspired (Mallya and Lazebnik, 2018) model for continual representation learning (Li et al., 2022).
+Learning algorithms with sample complexity guarantees were also proposed under data assumptions, such that all target tasks lie in a same space obtainable from a (linear or non-linear) combination of some input features (Cao et al., 2022).
+Finally, Friedman and Meir (2025) presented new PAC-Bayes bounds for the backward transfer loss of stochastic models trained without a replay buffer.
+In contrast with the listed works above, in this paper, we simultaneously provide risk certificates on the true risk of each task, our bound is non-asymptotic and can be computed from the training set.
+Moreover, our bound is applicable to any neural network architecture, under minimal data assumptions and does not have limitations on the continual learning setup that is employed.
 Sample compression.
-The sample compression theory has been shown effective for providing tight generalization bounds shah_margin-sparsity_2005 ; marchand2003set ; marchand2002set ; marchand2005learning ; shah_sample_2007 .
-However, most sample compress approaches are limited to low-complexity models.
-A notable exception is the method Pick-to-Learn (P2L), which successfully provides guarantees for deep neural networks campi2023compression ; bazinet2024 and Gaussian Processes marks2025pick2learn-gp .
-In this paper, we adapt the key ideas introduced in the sample compression theory to the continual learning case, which enables us to provide an upper bound on the test error from training samples.
+The sample compression theory has been shown to be effective for providing tight generalization bounds (Marchand and Shawe-Taylor, 2002; Marchand and Sokolova, 2005; Laviolette et al., 2005; Campi and Garatti, 2023).
+However, most sample compression approaches are limited to low-complexity models.
+A notable exception is the method Pick-to-Learn (P2L), presented by Paccagnan et al.
+(2024), which successfully provides guarantees for deep neural networks (Bazinet et al., 2025) and Gaussian Processes (Marks and Paccagnan, 2025).
+In this paper, we adapt the key ideas introduced in the sample compression theory to the continual learning case, which enables us to provide an upper bound on the true risk from training samples.
 2 BACKGROUND
 Continual learning.
-Let us consider a series of task distributions ${\mathcal{D}}_{1},{\mathcal{D}}_{2},\ldots,{\mathcal{D}}_{T}$ on an input-output space ${\mathcal{X}}\times{\mathcal{Y}}$, where the amount of tasks $T\in{\mathbb{N}}^{*}$ can be (countably) infinite.
+Let us consider a meta-distribution $\operatorname{\mathfrak{D}}$ and a sequence of i.i.d.
+(independent and identically distributed) task distributions ${\mathcal{D}}_{1},{\mathcal{D}}_{2},\ldots,{\mathcal{D}}_{T}\sim\operatorname{\mathfrak{D}}$ on an input-output space ${\mathcal{X}}\times{\mathcal{Y}}$.
 We are given a predictive model $f_{\theta}:{\mathcal{X}}\to{\mathcal{Y}}$ (e.g., a neural network architecture), with learnable parameters $\theta\in\Theta$.
@@
 We denote the training set at task $t$ as $S_{t}=\{({\mathbf{x}}_{t,i},y_{t,i})\}_{i=1}^{n_{t}}$, with ${\mathbf{x}}_{t,i}\in{\mathcal{X}}$ and $y_{t,i}\in{\mathcal{Y}}_{t}\subseteq{\mathcal{Y}}$; that is, the same input space $\mathcal{X}$} is shared among all tasks, and the output space $\mathcal{Y}$} may be the union of several task specific output spaces ${\mathcal{Y}}_{t}$.
-Given a loss function $\ell:\Theta\times{\mathcal{X}}\times{\mathcal{Y}}\to{\mathbb{R}}^{+}$, we want the last updated predictor $f_{\theta_{T}}$ to maintain a low generalization loss on all observed tasks $t\in\{1,\ldots,T\}$:
+Given a loss function $\ell:\Theta\times{\mathcal{X}}\times{\mathcal{Y}}\to[0,1]$, we want the last updated predictor $f_{\theta_{T}}$ to maintain a low generalization loss on all observed tasks $t\in\{1,\ldots,T\}$:
 ${\mathcal{L}}_{\operatorname{\mathcal{D}}_{t}}(\theta_{T})\ \coloneqq\operatorname*{\mathbf{E}}_{({\mathbf{x}},y)\sim{\mathcal{D}}_{t}}\ell(\theta_{T},{\mathbf{x}},y)\,.$ (1)
-The empirical loss counterparts on observed training samples are given by
-^$\widehat{\mathcal{L}}_{S_{t}}(\theta_{T})\ \coloneqq\ \frac{1}{n_{t}}\sum_{i=1}^{n_{t}}\ell(\theta_{T},{\mathbf{x}}_{t,i},y_{t,i})\,.$ (2)
-The challenge of continual learning lies in the fact that the learner observes the task datasets $S_{1},S_{2},\ldots,S_{t},\ldots,S_{T}$ sequentially, and we assume that the system cannot keep all observed data in memory.
-Nevertheless, at task $t$, we would like the system to perform well on all tasks, including the previous ones $11,\dots,t-$ 1.
-However, simply updating the parameters $\theta_{t-1}$ learned from a previous tasks to optimize the loss ^$\widehat{\mathcal{L}}_{S_{t}}(\theta_{t})$ on a current task $t$ would lead to catastrophic forgetting forgetting_french .
+The empirical loss counterpart on observed training samples is given by
+$\widehat{\mathcal{L}}_{S_{t}}(\theta_{T})\ \coloneqq\ \frac{1}{n_{t}}\sum_{i=1}^{n_{t}}\ell(\theta_{T},{\mathbf{x}}_{t,i},y_{t,i})\,.$ (2)
+The challenge of continual learning lies in the fact that the learner observes the task datasets $S_{1},S_{2},\ldots,S_{t},\ldots,S_{T}$ sequentially and that we assume that the system cannot keep all observed data in memory.
+Nevertheless, at task $t$, we would like the system to perform well on all tasks, including the previous ones $1,\dots,t-1$.
+However, simply updating the parameters $\theta_{t-1}$ learned from a previous tasks to optimize the loss $\widehat{\mathcal{L}}_{S_{t}}(\theta_{t})$ on a current task $t$ would lead to catastrophic forgetting (French, 1999).
 Therefore, as indicated in Section 1.1, numerous empirical methods have been developed to mitigate forgetting.
@@
 where the second term is the loss on the replay buffer $\mathcal{B}$}.
-In our experiments, we use this method as a baseline as it’s shown to be a simple and very effective continual learning strategy DellaLibera2023CLMASRAC ; Mai2020BatchlevelER .
 Sample compression theory.
-Given a training dataset $S=\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{n}\in(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{n}$, a family of learnable parameters $\Theta$ and a learning algorithm $A$ such that $A(S)\in\Theta$, sample compression theory provides generalization bounds for any predictor $f_{\theta}$, with $\theta=A(S)$, on the condition that $f_{\theta}$ can be provably represented as a function of a small subset of the training dataset $S$, called the compression set, and an additional source of information, called the message.
+Let us consider a training dataset $S=\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{n}\in(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{n}$ sampled i.i.d.
+from an unknown distribution $\operatorname{\mathcal{D}}$, a family of learnable parameters $\Theta$ and a learning algorithm $A$ such that $A(S)\in\Theta$.
+In this setting, sample compression theory provides generalization bounds for any predictor $f_{\theta}$, with $\theta=A(S)$, on the condition that $f_{\theta}$ can be provably represented as a function of a small subset of the training dataset $S$, called the compression set, and an additional source of information, called the message.
 If this is the case, we call $f_{\theta}$ a sample-compressed predictor.
-The compression set is denoted $S^{\operatorname{\mathbf{i}}}$ and is parameterized by a vector of indices $\operatorname{\mathbf{i}}\subset\{1,\ldots,n\}$.
-Each vector $\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)$, with $\operatorname{\mathscr{P}}(n)$ the set of all the possible subsets of $\{1,\ldots,n\}$, is ordered such that
-$\operatorname{\mathbf{i}}=(i_{1},i_{2},\ldots,i_{\operatorname{|\mathbf{i}|}})\,,\mbox{ with }1\leq i_{1}<\ldots<i_{\operatorname{|\mathbf{i}|}}\leq n\,.$ (4)
-The compression set is defined such that $S^{\operatorname{\mathbf{i}}}=S^{(i_{1},\ldots,i_{\operatorname{|\mathbf{i}|}})}\!=\{(\boldsymbol{x}_{i_{1}},y_{i_{1}}),\ldots,(\boldsymbol{x}_{i_{\operatorname{|\mathbf{i}|}}},y_{i_{\operatorname{|\mathbf{i}|}}})\}\!\subseteq S.$ We define the complement vector $\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}$ such that $\operatorname{\mathbf{i}}\cap\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}=\emptyset$ and $\operatorname{\mathbf{i}}\cup\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}=\{1,\ldots,n\}$.
-Thus, we have the complement set $S^{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}=S\setminus S^{\operatorname{\mathbf{i}}}$ and $|\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}|=n-\operatorname{|\mathbf{i}|}$.
-In addition to the compression set, a message $\mu$ is sometimes needed to compress the predictor.
-Although generally defined as a binary sequence, the message set can be defined by any set of countable sequences of symbols.
-Let $\Sigma=\{\sigma_{1},\sigma_{2},\ldots,\sigma_{N}\}$ be the alphabet used to construct the messages and $\Sigma^{*}$ be the set of all possible sequences, of length 00 to $\infty$, constructed using the alphabet $\Sigma$.
+The compression set is denoted $S^{\operatorname{\mathbf{i}}}$ and is parameterized by a strictly increasing sequence of indices $\operatorname{\mathbf{i}}{\,\in\,}\operatorname{\mathscr{P}}(n)$, with $\operatorname{\mathscr{P}}(n)$ the powerset of $\{1,\ldots,n\}$.
+The compression set is defined such that $S^{\operatorname{\mathbf{i}}}=S^{(i_{1},\ldots,i_{\operatorname{|\mathbf{i}|}})}\!=\{(\boldsymbol{x}_{i_{1}},y_{i_{1}}),\ldots,(\boldsymbol{x}_{i_{\operatorname{|\mathbf{i}|}}},y_{i_{\operatorname{|\mathbf{i}|}}})\}\!\subseteq S.$ We define the complement sequence $\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}$ such that $\operatorname{\mathbf{i}}\cap\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}=\emptyset$ and $\operatorname{\mathbf{i}}\cup\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}=\{1,\ldots,n\}$.
+Thus, we have the complement set $S^{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}=S\setminus S^{\operatorname{\mathbf{i}}}$ and $\operatorname{|\mathbf{i}^{{\rm\bf c}}|}=n-\operatorname{|\mathbf{i}|}$.
+In addition to the compression set, a message $\mu$ could be needed to compress the predictor.
+Although generally defined as a binary sequence, the message set can be any set of countable sequences of symbols.
+Let $\Sigma=\{\sigma_{1},\sigma_{2},\ldots,\sigma_{N}\}$ be the alphabet used to construct the messages and $\Sigma^{*}$ be the set of all possible sequences, of length 0 to $\infty$, constructed using the alphabet $\Sigma$.
 For all $\operatorname{\mathbf{i}}\,{\in}\,\operatorname{\mathscr{P}}(n)$, we choose $\mathscr{M}(\operatorname{\mathbf{i}})$ a countable subset of $\Sigma^{*}$, which represents all the possible messages that can be chosen for $\operatorname{\mathbf{i}}$.
 Given learned parameters $\theta=A(S)$, to prove that the predictor $f_{\theta}$ is a sample-compressed one, we need to define two functions: a compression function and a reconstruction function.
-The compression function $\mathscr{C}:\boldsymbol{\cup}_{1\leq n\leq\infty}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{n}\to\boldsymbol{\cup}_{m\leq n}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{m}\times\boldsymbol{\cup}_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\mathscr{M}(\operatorname{\mathbf{i}})$ outputs a compression set and a message.
+The compression function $\mathscr{C}:\boldsymbol{\cup}_{1\leq n\leq\infty}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{n}\to\boldsymbol{\cup}_{m\leq n}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{m}\times\boldsymbol{\cup}_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\mathscr{M}(\operatorname{\mathbf{i}})$ outputs a compression set $S^{\operatorname{\mathbf{i}}}$ and a message $\mu$.
 Then, the reconstruction function $\operatorname{\mathscr{R}}:\boldsymbol{\cup}_{m\leq n}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{m}\times\boldsymbol{\cup}_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\mathscr{M}(\operatorname{\mathbf{i}})\to\Theta$ is defined such that $A(S)=\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},\mu)$.
@@
 For any sample-compressed predictor $\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)$, this data-independent distribution is expressed as $P_{\overline{\Theta}}\quantity(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)){=}P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{i}})}(\mu)$, with $P_{\operatorname{\mathscr{P}}(n)}$ a probability distribution over $\operatorname{\mathscr{P}}(n)$ and $P_{\mathscr{M}(\operatorname{\mathbf{i}})}$ a probability distribution over $\mathscr{M}(\operatorname{\mathbf{i}})$.
-Following common practices marchand2005learning ; bazinet2024 , we consider hereunder $P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})={n\choose\operatorname{|\mathbf{i}|}}^{{}_{-1}}\zeta(\operatorname{|\mathbf{i}|})$, with $\zeta(k){=}\tfrac{6}{\pi^{{}_{2}}}(k+1)^{-2}$.
-Also, given a vector of indices $\operatorname{\mathbf{i}}$, we consider $P_{\mathscr{M}(\operatorname{\mathbf{i}})}$ to be a uniform distribution over the messages.
-We now present a recent sample compression generalization guarantee bazinet2024 on which we build our theoretical framework.
-We use the shorthand notation ^$\operatorname{\widehat{\mathcal{L}}}_{S}^{\,{}_{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}}(\cdot)$ to denote the empirical loss on the training sample that do not belong to the compression set: $S^{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}=S\setminus S^{\mathbf{i}}$; this is mandatory to obtain an unbiased estimate of the loss, as the reconstructed predictor $\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)$ relies on $S^{\mathbf{i}}$.
-Theorem 2.1 (bazinet2024 ).
-For any distribution $\operatorname{\mathcal{D}}$ over $\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}$, for any family of set of messages $\{\mathscr{M}(\operatorname{\mathbf{i}})|\operatorname{\mathbf{i}}{\in}\operatorname{\mathscr{P}}(n)\}$, for any deterministic reconstruction function $\operatorname{\mathscr{R}}$, for any loss $\ell:\Theta\times\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}{\to}[0,1]$ and for any $\delta\in(0,1]$, with probability at least $1-\delta$ over the draw of $S\sim\operatorname{\mathcal{D}}^{n}$, we have
+Following common practices (Marchand and Sokolova, 2005; Bazinet et al., 2025), we consider hereunder $P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}}){=}{n\choose\operatorname{|\mathbf{i}|}}^{{}_{-1}}\zeta(\operatorname{|\mathbf{i}|})$, with $\zeta(k){=}\tfrac{6}{\pi^{{}_{2}}}(k{+}1)^{-2}$.
+Also, given a sequence of indices $\operatorname{\mathbf{i}}$, we set $P_{\mathscr{M}(\operatorname{\mathbf{i}})}$ to a uniform distribution over the messages.
+We now present a recent generalization guarantee for sample compression (Bazinet et al., 2025), on which we build our theoretical framework.
+We use the shorthand notation $\operatorname{\widehat{\mathcal{L}}}_{S}^{{\;\;}_{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}}(\cdot)$ to denote the empirical loss on the training sample that do not belong to the compression set: $S^{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}=S\setminus S^{\mathbf{i}}$; it is mandatory to use $\operatorname{\widehat{\mathcal{L}}}_{S}^{{\;\;}_{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}}(\cdot)$ to obtain an unbiased estimate of the loss, as the predictor $\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)$ relies on $S^{\mathbf{i}}$ and the result wouldn’t hold otherwise.
+Theorem 2.1 (Bazinet et al., 2025).
+[…]
 $\forall\mathbf{i}\in\operatorname{\mathscr{P}}(n),\;\mu\in\mathscr{M}(\operatorname{\mathbf{i}}):$
-$\mathcal{L}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu))\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,{}_{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}}\quantity(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)),\frac{1}{n-\operatorname{|\mathbf{i}|}}\quantity[\log\!\matrixquantity(n\\ \operatorname{|\mathbf{i}|})\!\!+\!\log\!\quantity(\frac{2\sqrt{n-\operatorname{|\mathbf{i}|}}}{\zeta(\operatorname{|\mathbf{i}|})P_{\mathscr{M}(\operatorname{\mathbf{i}})}(\mu)\delta})])\,,$
-with $\mathrm{kl}(q,p)=q\ln\frac{q}{p}+(1-q)\ln\frac{1-q}{1-p}$ the binary Kullback-Leibler divergence and its inverse
-$\mathrm{kl}^{-1}(q,\varepsilon)=\mbox{\rm argsup}_{p\in[0,1]}\left\{\mathrm{kl}(q,p)\leq\varepsilon\right\}.$ (5)
-On the one hand, given a fixed compression set size $\operatorname{|\mathbf{i}|}$, the bound decreases when the training set size $n$ increase.
-On the other hand, given a fixed $n$, the bound increases with $\operatorname{|\mathbf{i}|}$.
+$\mathcal{L}_{\operatorname{\mathcal{D}}}\quantity(\theta_{\operatorname{\mathbf{i}},\mu})\leq\mathrm{kl}^{-1}\left(\operatorname{\widehat{\mathcal{L}}}_{S}^{{\;\;}_{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}}\quantity(\theta_{\operatorname{\mathbf{i}},\mu}),\frac{1}{\operatorname{|\mathbf{i}^{{\rm\bf c}}|}}\ln\frac{2\smallmatrixquantity(n\\ \operatorname{|\mathbf{i}|})\sqrt{\operatorname{|\mathbf{i}^{{\rm\bf c}}|}}}{\zeta(\operatorname{|\mathbf{i}|})P_{\mathscr{M}(\operatorname{\mathbf{i}})}(\mu)\delta}\right),$
+with $\theta_{\operatorname{\mathbf{i}},\mu}\,{=}\,\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)$ the reconstructed parameters, $\mathrm{kl}(q,p)=q\ln\frac{q}{p}+(1-q)\ln\frac{1-q}{1-p}$ the binary Kullback-Leibler divergence and its inverse
+$\mathrm{kl}^{-1}(q,\varepsilon)=\operatorname*{arg\,sup}_{p\in[0,1]}\left\{\mathrm{kl}(q,p)\leq\varepsilon\right\}.$ (4)
+On the one hand, given a fixed compression set size $\operatorname{|\mathbf{i}|}$, the bound decreases when the training set size $n$ increases.
+On the other hand, given a fixed dataset size $n$, the bound increases with $\operatorname{|\mathbf{i}|}$.
 Pick-to-Learn.
 To obtain sample compression guarantees for a class of predictors, it is necessary to prove that the learned predictor only depends on a small subset of the data and (optionally) a message.
-Some predictors, such as the SVM boser_svm and the perceptron rosenblatt1958perceptron ; moran2020perceptron , have straightforward compression scheme, as they only depend on a subset of the data after training.
-Some algorithms, such as the set covering machine (SCM) marchand2002set ; marchand2003set ; marchand2005learning and decision trees shah_sample_2007 , necessitate hand-crafted compression schemes involving a message.
+Some predictors, such as the SVM (Boser et al., 1992) and the perceptron (Rosenblatt, 1958; Moran et al., 2020), have a straightforward compression scheme, as they only depend on a subset of the data after training.
+Some learning algorithms, such as the set covering machine (SCM) (Marchand and Shawe-Taylor, 2002; Marchand et al., 2003; Marchand and Sokolova, 2005) and decision trees (Shah, 2007), necessitate hand-crafted compression schemes involving a message.
 Up until recently, there was no sample compression scheme for deep neural networks.
-Algorithm 1 Original Pick-To-Learn (P2L) paccagnan_pick_learn_2023
+Algorithm 1 Pick-To-Learn (P2L)    (Paccagnan et al., 2024)
+0:  $\theta_{\rm init}$ {Initial parameters of the model $f_{\theta}$}
 0:  $S=\{({\mathbf{x}}_{i},y_{i})\}_{i=1}^{n}$ {Training set}
-0:  $\theta_{\rm init}$Initial parameters of the model $f_{\theta}$}
 0:  $\gamma$ {Stopping criteria}
@@
 9:  return $\theta_{k}$ {Final parameters}
-The meta-algorithm Pick-To-Learn (P2L) was proposed by Paccagnan and Campi paccagnan_pick_learn_2023 as a compression scheme for any class of predictors, with a specific focus to deep neural networks.
-The meta-algorithm acts as an aditional training loop, by iteratively selecting datapoints over which the model is updated.
+The meta-algorithm Pick-To-Learn (P2L) was proposed by Paccagnan et al.
+(2024) as a compression scheme for any class of predictors, with a specific focus on deep neural networks.
+The meta-algorithm acts as an additional training loop, by iteratively selecting datapoints over which the model is updated.
 Starting with initial parameters $\theta_{\rm init}$, P2L evaluates the predictor on the whole training dataset, adds the datapoints with the highest losses to the compression set and then updates the predictor using the compression set.
-The meta-algorithm stops once the losses of training examples not in the compression set are lower than a given stopping criteria ($-\ln(0.5)$ for the cross-entropy loss, which is equivalent to achieving zero errors on the complement set $S^{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}$).
-We provide the pseudo-code of the original P2L in Algorithm 1.
-Recent works introduce improvements to the original scheme, notably early stopping strategies marks2025pick2learn-gp ; bazinet2024 and extension to the Bayesian setting marks2025pick2learn-gp .
+The meta-algorithm stops once the losses of training points that do not belong to the compression set are lower than a given stopping criterion $\gamma$.
+We provide the pseudo-code of P2L in Algorithm 1.
+Recent works introduce improvements to the original scheme, notably early stopping strategies (Bazinet et al., 2025; Paccagnan et al., 2025) and extension to the Bayesian setting (Marks and Paccagnan, 2025).
 3 SAMPLE COMPRESSION FOR CONTINUAL LEARNING
@@
 The winning strategy is to select this compression set to ensure a low risk on the training samples not being retained in the compression set (i.e., the complement of the compression set).
-This observation motivates our strategy for managing the replay buffer of our new algorithm Continual Pick-To-Learn (CoP2L): we subsample the complement set to create the replay buffer, in order to maintain a low risk on these examples while learning subsequent tasks.
-When needed, CoP2L can choose to add a datapoint from the buffer to the compression set, thus mitigating the forgetting by focusing on a well-chosen datapoint.
+This observation motivates the strategy employed in our new algorithm Continual Pick-To-Learn (CoP2L), which manages the replay buffer by subsampling only from the complement set instead of the whole training set.
+Then, while learning subsequent tasks, CoP2L maintains a low risk on previous tasks by adding well-chosen datapoints from the buffer to its compression set.
+This strategy inherently both mitigates the forgetting and enables the computation of sample compression bounds.
 Algorithm 2 Modified Pick-To-Learn (mP2L)
@@
 12:  end if
-13:  return $\theta_{k},C_{k}$. {Final parameters and compression set}
+13:  return $\theta_{k},C_{k}$ {Final parameters and compression set}
+In CoP2L, each new task is learned by a modified version of the P2L meta-algorithm (Algorithm 2, entitled mP2L), as explained further down.
+That is, for each new task, the proposed continual learning algorithm (Algorithm 3) calls mP2L and then updates the replay buffer by sampling datapoints that were not chosen in the compression set.
+Similar to the replay buffer method, at the end of each task $t$, the buffer contains $\lfloor\tfrac{m}{t}\rfloor$ datapoints of each previous task, with $m$ the maximum buffer size.
+We modified Pick-To-Learn in two significant ways.
+(i) First, we introduced weights to the loss functions to tackle the imbalance problem between the current task and the previous tasks.
+Traditionally, when working with a replay buffer, the class imbalance problem is taken care of by training on a number of datapoints from previous tasks.
+As our model is only trained on a very limited subset of the data, another strategy must be implemented to mitigate the class imbalance effect.
+Before starting the training on a new task, mP2L sets the weight of the datapoints from the previous task to $\omega\,{>}\,1$ and the weight of the current task to $1$.
+For the stopping criteria of mP2L to be satisfied, the worst loss on the current task must be less than $\gamma$ and the worst loss on the previous tasks must be less than $\tfrac{\gamma}{\omega}$.
+(ii) The original P2L algorithm trains the model until it first achieves zero errors on the complement set.
+Instead, mP2L performs early stopping based on the bound value (as proposed by Bazinet et al., 2025).
+That is, mP2L relies on the trade-off encoded in Theorem˜2.1 between the accuracy on the complement set $C$ and the complexity of the model $f_{\theta}$.
+More precisely, it returns the model’s parameters $\theta$ that minimizes
+$\Psi(S,\theta,C)\!=\!\mathrm{kl}^{-1}\!\left(\!\operatorname{\widehat{\mathcal{L}}}_{S\setminus C}(\theta),\tfrac{1}{|S\setminus C|}\ln\tfrac{2\sqrt{|S\setminus C|}{|S|\choose|C|}}{\zeta(|C|)\delta}\right)\!,\!$ (5)
+with $\mathrm{kl}^{-1}$ given by Eq. (4). Thanks to Theorem˜2.1, we have that $\operatorname{\mathcal{L}}_{\mathcal{D}}(\theta)\leq\Psi(S,\theta,C)$ with high probability.
+Note that for the first observed task, the mP2L procedure starts from randomly initialized parameters $\theta_{0}$ and is executed on the training sample $S_{1}$ (as it is done by the original P2L).
+Then, for every subsequent task $t\in\{2,\ldots,T\}$, the mP2L procedure is initialized to the previously learned parameters $\theta_{t-1}$.
+It then learns from the current task sample $S_{t}$ and a subset of randomly selected instances from previous tasks.
+The latter is obtained from the procedure sample(S,m)(S,m) (see Algorithm 3), which represents the random sampling of $m$ instances of $S$ without replacement.
+[^2] Although we only discuss random sampling, any buffer management technique applied to the complement set would be valid in CoP2L, such as coreset methods.
 Algorithm 3 Continual Pick-To-Learn (CoP2L)
@@
 10:  return $\theta_{T}$. {Final parameters}
-Each new task is learned by a modified version of the P2L meta-algorithm (Algorithm 2, entitled mP2L), as explained further down.
-That is, for each new task, the proposed continual learning algorithm (Algorithm 3, entitled CoP2L) calls mP2L and updates the buffer using the datapoints that were not chosen in the compression set.
-Similarly to the replay buffer method, at the end of each task $t$, the buffer contains $\lfloor\tfrac{m}{t}\rfloor$ datapoints of each previous task, with $m$ the maximum buffer size.
-We modified Pick-To-Learn in two significant ways.
-(i) First, we introduced weights to the loss functions to tackle the imbalance problem between the current task and the previous tasks.
-When working with a replay buffer, the class imbalance problem is taken care of by training on a number of datapoints from previous tasks.
-As our model is only trained on a very limited subset of the data, we need to find another way to mitigate the effect of class imbalance.
-Before starting the training on a new task, we set the weight of the datapoints from the previous task to $\omega$ and the weight of the current task to $1$.
-For the stopping criteria of mP2L to be satisfied, the worst loss on the current task must be less than $\gamma$ and the worst loss on the previous task must be less than $\tfrac{\gamma}{\omega}$.
-This ensures that the model has achieved perfect accuracy on the current task and all the datapoints from the previous tasks that are found in the buffer.
-(ii) The second modification is the choice of the returned predictor by mP2L.
-The original P2L algorithm trains the model until it first achieves zero errors on the complement set.
-In presence of a noisy dataset, this might lead to an overfitting behavior, P2L adding most training samples into its compression set, which in turn gives a large generalization bound value.
-Instead, we perform early stopping based on the bound value (as proposed by bazinet2024 ).
-That is, mP2L relies on the trade-off encoded in Theorem 2.1 between the accuracy on the complement set and the complexity of the model.
-More precisely, it returns the model’s parameters $\theta$ that minimizes
-$\Psi(S,\theta,C)$ $=\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S\setminus C}\quantity(\theta),\ \tfrac{1}{|S\setminus C|}\log\quantity(\tfrac{2\sqrt{|S\setminus C|}{|S|\choose|C|}}{\zeta(|C|)\delta})),$ (6)
-with $\mathrm{kl}^{-1}$ defined by Equation 5. Thanks to Theorem 2.1, we have that $\operatorname{\mathcal{L}}_{\mathcal{D}}(\theta)\leq\Psi(S,\theta,C)$ with probability at least $1-\delta$.
-Note that for the first observed task, the mP2L procedure starts from randomly initialized parameters $\theta_{0}$ and is executed on the training sample $S_{1}$ (as it is done by the original P2L).
-Then, for every subsequent task $t\in\{2,\ldots,T\}$, the mP2L procedure is initialized to the previously learned parameters $\theta_{t-1}$.
-It then learns from the current task sample $S_{t}$ and a subset of randomly selected instances from previous tasks.
-The latter is obtained from the procedure sample$(S,m)$ (see Algorithm 3), which represents the random sampling of $m$ instances of $S$ without replacement.
 Compression and reconstruction scheme.
-From a theoretical standpoint, in contrast to the original P2L, both mP2L and CoP2L cannot be reconstructed straightforwardly without a message for two reasons: (1) when only the compression set is given as input, the bound function $\Psi(S,\theta,C)$ used as stopping criteria in mP2L cannot be computed on the whole dataset; (2) the sampling procedure sample$(S,m)$ in CoP2L cannot be reproduced in the reconstruction step.
-In Appendix A, we provide the compression and reconstruction functions of CoP2L, which give rise to sample compression bounds as presented in Theorem 3.1.
-Noteworthy, after learning on $T$ tasks, the compression function provides two sample compression sets $S^{\operatorname{\mathbf{i}}}$ and $S^{\operatorname{\mathbf{j}}}$, along with a message pair $(\mu_{1},\mu_{2})$ chosen among the set of all possible messages, denoted as $\mathscr{M}(\operatorname{\mathbf{j}},T)=\{2,\ldots,T\}^{|\operatorname{\mathbf{j}}|}\times\left[\boldsymbol{\times}_{t=1}^{T}\{1,\ldots,n_{t}+|{\mathcal{B}}|\}\right].$ As explained further in appendix, the use of two compression sets, along with the message $\mu_{1}=[\mu_{1}^{i}]_{i=1}^{T}$, serves to the proper treatement of the buffer $B^{\star}$, while the message $\mu_{2}=[\mu_{2}^{i}]_{i=1}^{T}$ contains the number of iterations $K^{\star}$ to perform for each call to mP2L.
+From a theoretical standpoint, in contrast to the original P2L, both mP2L and CoP2L cannot be reconstructed straightforwardly without a message for two reasons: (1) when only the compression set is given as input, the bound function $\Psi(S,\theta,C)$ used as stopping criteria in mP2L cannot be computed on the whole dataset; (2) the sampling procedure sample(S,m)(S,m) in CoP2L cannot be reproduced in the reconstruction step.
+In Appendix B, we prove that CoP2L is a sample-compression algorithm by providing its compression and reconstruction functions, which subsequently gives rise to sample-compression bounds as presented in Theorem˜3.1.
+Noteworthy, after learning on $T$ tasks, the compression function provides two compression sets $S^{\operatorname{\mathbf{i}}}$ and $S^{\operatorname{\mathbf{j}}}$, along with a message pair $\mu=(\mu_{1},\mu_{2})$.
+As explained further in the appendix, the use of two compression sets, along with the first part of the message $\mu_{1}=[\mu_{1}^{i}]_{i=1}^{T}$, serves the proper treatment of the weighted buffer $B^{\star}$, while the second part of the message $\mu_{2}=[\mu_{2}^{i}]_{i=1}^{T}$ contains the number of iterations $K^{\star}$ to perform for each call to mP2L.
+The set of all possible message pairs is denoted as $\mathscr{M}_{1:T}(\operatorname{\mathbf{j}})=\{2,\ldots,T\}^{|\operatorname{\mathbf{j}}|}\times\left[\boldsymbol{\times}_{t=1}^{T}\{1,\ldots,n_{t}+|{\mathcal{B}}|\}\right].$
 Generalization bound.
-Given the compression and reconstruction function detailed in Apprendix A, Theorem 2.1 can be used to obtain generalization bounds on the last learned task, which is done by mP2L to compute the bound $\Psi$.
+Given the compression and reconstruction function detailed in Appendix B, Theorem˜2.1 can be used to obtain generalization bounds on the last learned task, which is done by mP2L to compute the bound $\Psi$.
 However, this bound holds for only one distribution of data.
-Therefore, it cannot be applied to bound the risks of the learned predictor on previously tasks seen by CoP2L.
+Therefore, it cannot be applied to bound the risks of the learned predictor on previous tasks seen by CoP2L.
 The next Theorem 3.1 holds for any previously learned tasks.
-For a dataset $S_{t}\sim\operatorname{\mathcal{D}}_{t}$, we denote the loss on the complement set of the joint set $(S_{t}^{\operatorname{\mathbf{i}}}\cup S_{t}^{\operatorname{\mathbf{j}}})^{\rm\bf c}$ as ^$\widehat{{\mathcal{L}}}_{S_{t}}^{\ \operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\!(\theta)$.
-Moreover, we denote the reconstruction function of CoP2L as $\operatorname{\mathscr{R}}_{1:T}(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)=\operatorname{\mathscr{R}}\big{(}S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu\mid S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}\big{)}.$ This formulation is important for the following theorem, as the reconstruction function for task $t$ is conditioned on all datasets $S_{1}$ to $S_{T}$, with the exception of $S_{t}$.
+For a dataset $S_{t}\sim\operatorname{\mathcal{D}}_{t}$, we denote the loss on the complement set of the joint set $S_{t}^{\operatorname{\mathbf{i}}}\cup S_{t}^{\operatorname{\mathbf{j}}}$ as $\widehat{{\mathcal{L}}}_{S_{t}}^{\ \operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\!(\theta)$.
+Moreover, we denote the reconstruction function of CoP2L as $\operatorname{\mathscr{R}}_{1:T}\big(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu\mid S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}\big).$ This formulation is important for the following theorem, as the reconstruction function for task $t$ is conditioned on all datasets $S_{1}$ to $S_{T}$, with the exception of $S_{t}$.
 Theorem 3.1.
-For any set of independent distributions $\left\{{\mathcal{D}}_{t}\right\}_{t=1}^{T}$ over ${\mathcal{X}}\times{\mathcal{Y}}$, with $\operatorname{\mathscr{R}}_{1:T}$ the reconstruction function of CoP2L (Algorithm 5), with the family of set of messages $\{\mathscr{M}(\operatorname{\mathbf{j}},T)|\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)\}$ defined for CoP2L, for any loss $\ell:\Theta\times{\mathcal{X}}\times{\mathcal{Y}}\to[0,1]$, and for any $\delta\in(0,1]$, with probability at least $1-\delta$ over the draw of $S_{t}\sim{\mathcal{D}}_{t},t=1,\ldots,T$, we have
-$\forall t\in\{1,\ldots,T\},{\mathbf{i}}\in\operatorname{\mathscr{P}}(n_{t}),\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n_{t}),\mu=(\mu_{1},[\mu_{2}^{i}]_{i=1}^{T})\in\mathscr{M}(\operatorname{\mathbf{j}},T):$
-$\mathrm{kl}\quantity(\widehat{{\mathcal{L}}}_{S_{t}}^{\ \operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\!\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)),\ {\mathcal{L}}_{\operatorname{\mathcal{D}}_{t}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)))$
-$\leq\tfrac{1}{n_{t}-|{\mathbf{i}}|-|\mathbf{j}|}\big{[}\log{\textstyle{n_{t}\choose|{\mathbf{i}}|}}+\log{\textstyle{n_{t}-\operatorname{|\mathbf{i}|}\choose|\mathbf{j}|}}+\textstyle\sum_{i=1}^{T}\log\tfrac{1}{\zeta(\mu_{2}^{i})}+|\mathbf{j}|\log(T-1)+\log\tfrac{T}{\zeta(|{\mathbf{i}}|)\zeta(|\operatorname{\mathbf{j}}|)\delta}\big{]}\,.$
-The proof is given in Appendix C.
-It first relies on a tighter version of Theorem 2.1, where the $2\sqrt{n-\operatorname{|\mathbf{i}|}}$ term is removed and a second compression set is added.
-Then, this new result is adapted to the continual learning setting.
-For all tasks $1$ through $T$, this new theorem encodes a tradeoff similar to Theorem 2.1 between the predictor accuracy and its complexity.
+For any set of distributions $\left\{{\mathcal{D}}_{t}\right\}_{t=1}^{T}$ over ${\mathcal{X}}\times{\mathcal{Y}}$ sampled from $\operatorname{\mathfrak{D}}$, and for any $\delta\in(0,1]$, with probability at least $1-\delta$ over the draw of $S_{t^{\prime}}\sim{\mathcal{D}}_{t^{\prime}},{t^{\prime}}=1,\ldots,T$, we have
+$\forall t\in\{1,\ldots,T\},\ {\mathbf{i}},\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n_{t}),\ \mu\!=\!(\mu_{1},\mu_{2})\in\mathscr{M}_{1:T}(\operatorname{\mathbf{j}}):$
+${\mathcal{L}}_{\operatorname{\mathcal{D}}_{t}}\quantity(\theta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}^{(t)})\leq\mathrm{kl}^{-1}\left(\widehat{{\mathcal{L}}}_{S_{t}}^{\ \operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\!\quantity(\theta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}^{(t)}),\frac{\epsilon(\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu)}{n_{t}-|{\mathbf{i}}|-|\mathbf{j}|}\right),$
+with $\theta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}^{(t)}=\operatorname{\mathscr{R}}_{1:T}\big(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu\,{\mid}\,S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}\big)$ the reconstructed parameters given by Algorithm˜5 and
+$\epsilon(\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu)=\ln\left[\tfrac{T}{\delta}\textstyle{n_{t}\choose\operatorname{|\mathbf{i}|}}\textstyle{n_{t}-\operatorname{|\mathbf{i}|}\choose|\mathbf{j}|}\tfrac{(T-1)^{|\mathbf{j}|}}{\zeta(|{\mathbf{i}}|)\zeta(|\operatorname{\mathbf{j}}|)}\prod_{i=1}^{T}\tfrac{1}{\zeta(\mu_{2}^{i})}\right]\!.$
+The proof is given in Appendix˜D.
+It first relies on a tighter version of Theorem˜2.1, where the $2\sqrt{n-\operatorname{|\mathbf{i}|}}$ term is removed and a second compression set is considered.
+This new result is tailored to the continual learning setting.
+For all tasks $1$ through $T$, Theorem˜3.1 encodes a tradeoff similar to Theorem˜2.1 between the predictor accuracy and its complexity.
 Here, the predictor complexity is evaluated using the size of the compression sets $\operatorname{\mathbf{i}}$ and $\operatorname{\mathbf{j}}$ and the probability of choosing a message.
@@
 Thus, the larger the compression sets, the smaller the probability of both messages.
-In conclusion, the bound encodes the ability of a model to minimizes the loss on the dataset whilst still keeping $\operatorname{\mathbf{i}}$ and $\operatorname{\mathbf{j}}$ the small.
+In conclusion, the bound encodes the ability of a model to minimize the loss on the dataset whilst still keeping $\operatorname{\mathbf{i}}$ and $\operatorname{\mathbf{j}}$ small.
 4 EXPERIMENTS
+In this section, we present experiments with CoP2L to demonstrate the tightness of our new generalization bounds for continual learning in two different settings.
+Afterward, we compare CoP2L to established baseline approaches to show that although our algorithm was first designed to obtain tight generalization bounds, its performances are still very competitive with established continual learning schemes.
 Datasets and continual learning settings.
-We carry-out experiments on MNIST lecun2010mnist , Fashion-MNIST fashionmnist , EMNIST (characters) cohen2017emnistextensionmnisthandwritten , PermutedMNIST goodfellow2013empirical , Rotated MNIST, CIFAR10 cifar10cifar100 , CIFAR100 cifar10cifar100 and TinyImageNet tiny-imagenet .
-We report results that include Class-Incremental (CI), Task-Incremental (TI), and Domain-Incremental (DI) continual learning settings.
-The dataset stats regarding how many classes/data items are available per task is given in Appendix H.
+We carry out experiments on CIFAR10 (Krizhevsky, 2012), CIFAR100 (Krizhevsky, 2012) and TinyImageNet (mnmoustafa and Ali, 2017).
+We report results in Class-Incremental (CI) and Task-Incremental (TI) continual learning settings, as defined in the Avalanche framework (Carta et al., 2023).
 Baselines.
-To assess the effectiveness of our proposed method CoP2L, we benchmark against several strong and widely adopted baselines across Class-Incremental, Task-Incremental, and Domain-Incremental continual learning settings.
-For each continual learning set-up we include baselines that yield the best performance in each category according to the Avalanche benchmark lomonaco2021avalanche .
-For the Class-Incremental setting, we use the baselines Replay rolnick2019experience , iCaRL rebuffi2017icarl , and GDumb prabhu2020gdumb .
-Replay in particular is considered a strong baseline DellaLibera2023CLMASRAC due to its simplicity and effectiveness.
-In the Task-Incremental setting, we compare against Replay rolnick2019experience , LaMAML gupta2020look and Learning without Forgetting (LwF) li2017learning .
-These methods are well suited for TI scenarios where task identity is known at test time.
-In the Domain-Incremental setting, we evaluate against Replay rolnick2019experience , Less Forgetful Learning (LFL) jung2018less , Elastic Weight Consolidation (EWC) kirkpatrick2017overcoming , and Synaptic Intelligence (SI) zenke2017continual .
-Implementation.
-All our experiments are conducted using the Avalanche continual learning framework avalanchelib , which provides a unified interface for datasets, benchmarks, and baseline implementations.
-We also use the official ContinualAI benchmark repository [^1] https://github.com/ContinualAI/continual-learning-baselines for consistent implementation and fair comparisons.
-The code for the experimental results is available in the supplemental material.
-Model architecture and training details.
-We use the SimpleMLP architecture and customized SimpleCNN architectures (SimpleMLP and SimpleCNN from the Avalanche toolkit) for the MNIST, FMNIST, and EMNIST experiments.
-All models are trained using the SGD optimizer.
-The details of the models architectures and the hyperparameters used can be found in Appendix G.
-For experiments on more complex datasets such as CIFAR10, CIFAR100, and TinyImageNet, we use the pretrained feature extractor Dino caron2021emerging using ViT dosovitskiy2020image and ResNet50 he2016deep backbones, followed by a trainable MLP head.
-This choice is motivated by the insight that CoP2L can further benefit from stronger feature representations, especially in settings with limited data per class or high dataset complexity.
-By leveraging Dino’s obtained embeddings, we enhance the quality of the input representations, which we empirically observe to yield good performance.
-Results with standard deviations across seeds and total training times are reported in Appendix F.
-We used a single NVIDIA 4090 GPU for our experiments.
-Accuracy and Forgetting.  We denote the accuracy obtained on a task $t$, after training the model for $T$ tasks with $\text{Accuracy}_{T,t}$. The accuracy we report is defined as,
-. $\textit{Average Accuracy}_{T}=\frac{1}{T}\sum_{t=1}^{T}\text{Accuracy}_{T,t}$, where $t$ is for the task index. .
-We also report the average forgetting at task $T$ which is defined as
-$\textit{Average Forgetting}_{T}=\frac{1}{(T-1)}\sum_{t=1}^{T-1}[\text{Accuracy}_{t,t}-\text{Accuracy}_{T,t}],$ which is calculated when $T>1$.
-4.1 Main Results
-In Table 1, we present the average accuracy and average forgetting values defined in the section above, obtained on MNIST, FMNIST and EMNIST dataset, both with an MLP and a CNN architecture, where the networks are trained from scratch.
-In Appendix F, we also provide the standard deviations as well.
-We observe that with the MLP architecture CoP2L is able to outperform the baselines.
-Furthermore, we also observe that CoP2L is able to obtain competitive results when employed with a CNN architecture as well.
-In Appendix E, we also provide an analysis of the performance of CoP2L for varying buffer sizes.
-In Table 2, we extended our results to cover CIFAR10, CIFAR100, and TinyImageNet datasets, for both Class-Incremental and Task-Incremental settings.
-We also provide results for Domain-Incremental setting with Permuted-MNIST goodfellow2013empirical and Rotated-MNIST datasets.
-Table 1: Summary of average accuracies and forgetting values obtained on MNIST, FMNIST, EMNIST (5 tasks, Class-Incremental), and on Permuted-MNIST and Rotated-MNIST (10 tasks, Domain-Incremental).
-Method / Dataset MNIST FMNIST EMNIST
-Acc / For Acc / For Acc / For
-CoP2L (MLP) 95.1 / 3.4 84.2 / 12.1 78.0 / 16.7
-Replay (MLP) 94.5 / 6.1 84.0 / 15.9 77.2 / 22.1
-GDumb (MLP) 91.7 / 3.8 81.8 / 10.4 62.6 / 16.0
-CoP2L (CNN) 96.6 / 2.2 87.4 / 11.1 87.9 / 8.4
-Replay (CNN) 96.9 / 3.3 87.6 / 13.3 87.7 / 12.3
-GDumb (CNN) 94.4 / 3.0 85.0 / 8.8 85.1 / 7.5
-Method / Dataset PermMNIST RotMNIST
-Acc / For Acc / For
-CoP2L (MLP) 92.3 / 6.0 88.3 / 9.8
-Replay (MLP) 91.5 / 6.4 88.2 / 10.0
-LFL (MLP) 83.5 / 9.3 51.6 / 45.7
-EWC (MLP) 83.7 / 5.2 56.5 / 35.4
-SI (MLP) 86.7 / 10.7 57.7 / 43.9
+To assess the effectiveness of our proposed method CoP2L, we benchmark against several strong and widely adopted baselines across Class-Incremental and Task-Incremental learning settings.
+For the Class-Incremental setting, we use the baselines Replay (Rolnick et al., 2019), Dark Experience Replay (DER) (Buzzega et al., 2020), iCaRL (Rebuffi et al., 2017b), GDumb (Prabhu et al., 2020), Contrastive Continual Learning via Importance Sampling (CCLIS) (Li et al., 2024) and Coreset Selection via Reducible Loss (CSReL) (Tong et al., 2025)
+In the Task-Incremental setting, we report Replay, DER, LaMAML (Gupta et al., 2020), Learning without Forgetting (LwF) (Li and Hoiem, 2017), CCLIS and CSReL.
+(a) Class-incremental setting
+(b) Task-incremental setting
+Figure 2: Illustration of the behavior of the bound on CIFAR10 with 5 tasks.
+(a) Class-incremental setting
+(b) Task-incremental setting
+Figure 3: Illustration of the behavior of the bound on CIFAR100 with 20 tasks.
 Table 1: Summary of average accuracies and forgetting values obtained on various continual learning tasks on CIFAR10, CIFAR100 and TinyImageNet datasets.
-CI denotes Class-Incremental, TI denotes Task-Incremental continual learning settings.
-CIFAR10 (5 tasks), CI CIFAR100 (10 tasks), CI CIFAR100 (20 tasks), CI TinyImNet (40 tasks), CI
+(a) Class-Incremental experiments
+CIFAR10 (5 tasks) CIFAR100 (10 tasks) CIFAR100 (20 tasks) TinyImNet (40 tasks)
 Method Acc ($\uparrow$) Forg ($\downarrow$) Acc ($\uparrow$) Forg ($\downarrow$) Acc ($\uparrow$) Forg ($\downarrow$) Acc ($\uparrow$) Forg ($\downarrow$)
-CoP2L-Dino (ViT) 94.45 3.36 73.83 16.76 72.50 16.96 51.13 36.81
-Replay-Dino (ViT) 94.01 6.01 68.51 30.44 69.23 29.60 47.66 48.15
-iCaRL-Dino (ViT) 93.24 1.55 61.50 14.53 64.11 15.25 20.78 52.95
-GDumb-Dino (ViT) 94.15 3.39 72.73 11.70 72.05 11.38 54.53 14.79
-CoP2L-Dino (RN50) 81.13 5.55 49.18 21.99 49.98 24.32 32.74 36.40
-Replay-Dino (RN50) 81.83 17.09 49.40 42.07 49.91 44.41 28.43 63.26
-iCaRL-Dino (RN50) 74.56 1.54 45.97 8.71 46.87 6.32 33.13 10.95
-GDumb-Dino (RN50) 80.70 8.95 43.83 20.31 43.78 19.03 24.04 21.82
-CIFAR10 (5 tasks), TI CIFAR100 (10 tasks), TI CIFAR100 (20 tasks), TI TinyImNet (40 tasks), TI
-CoP2L-Dino (ViT) 99.01 -0.03 95.15 0.92 97.10 0.71 94.28 1.22
-Replay-Dino (ViT) 99.24 -0.03 96.59 0.23 98.06 0.41 95.82 0.36
-LaMAML-Dino (ViT) 99.18 0.06 95.77 0.40 97.70 0.33 94.90 1.15
-LwF-Dino (ViT) 97.53 2.08 95.26 1.06 95.73 2.15 85.79 9.09
-CoP2L-Dino (RN50) 96.17 0.79 86.58 2.26 90.88 2.37 88.75 2.13
-Replay-Dino (RN50) 96.96 0.42 87.59 0.64 93.13 0.16 90.53 0.74
-LaMAML-Dino (RN50) 96.71 0.66 83.13 6.81 90.57 3.35 87.25 4.55
-LwF-Dino (RN50) 95.90 0.15 79.29 0.67 87.94 0.09 85.15 0.90
-Figure 1: Example demonstration of emprical upper bounds on test errors, on Class-Incremental (top-left CIFAR10, top-right CIFAR100) and Task-Incremental (bottom-left CIFAR10, bottom-right CIFAR100) datasets.
-Empirical study of the CoP2L generalization bound.
-In Figure 1, we showcase the empirical upper bounds on test errors.
-The details on the computation of the bound can be found in the Appendix  A.
-We observe that the bound values are non-trivial and generally follow similar trends as the actual error we report on the test set, while the estimate for the bound is calculated exclusively on the training set.
-In Figure 1, we provide plots for the bounds for the models trained with a buffer size of 2000, for Class-Incremental and Task-Incremental settings.
+CoP2L (ViT) 94.45 2.10 73.49 17.06 70.56 21.15 51.59 36.32
+Finetuning (ViT) 27.55 89.74 12.87 93.27 8.49 94.64 2.83 95.36
+Replay (ViT) 94.00 6.11 68.71 30.02 69.21 29.52 47.49 48.39
+DER (ViT) 95.03 3.79 72.29 25.11 77.03 19.27 53.92 41.77
+iCaRL (ViT) 88.60 3.13 65.54 13.64 62.36 15.88 21.06 34.49
+GDumb (ViT) 94.16 3.16 72.51 11.53 72.35 11.77 53.93 15.11
+CCLIS (ViT) 93.15 4.44 67.95 23.68 66.81 24.42 48.55 32.26
+CoP2L (RN50) 80.98 5.84 49.53 22.31 46.74 22.54 33.30 35.93
+Finetuning (RN50) 19.63 97.22 9.36 89.93 4.88 92.81 1.80 78.66
+Replay (RN50) 81.38 17.72 49.41 42.12 49.44 44.96 28.38 63.36
+DER (RN50) 82.35 15.31 49.37 44.01 57.53 33.37 28.77 64.17
+iCaRL (RN50) 74.49 1.79 46.06 8.50 46.47 6.74 33.89 10.42
+GDumb (RN50) 80.84 9.35 43.37 20.23 43.60 19.93 24.29 21.29
+CCLIS (RN50) 76.89 12.32 44.22 23.23 39.10 25.78 18.32 40.40
+CSReL (RN50) 34.50 26.44 38.75 42.12 32.57 40.23 12.84 73.69
+(b) Task-Incremental experiments
+CIFAR10 (5 tasks) CIFAR100 (10 tasks) CIFAR100 (20 tasks) TinyImNet (40 tasks)
+Method Acc ($\uparrow$) Forg ($\downarrow$) Acc ($\uparrow$) Forg ($\downarrow$) Acc ($\uparrow$) Forg ($\downarrow$) Acc ($\uparrow$) Forg ($\downarrow$)
+CoP2L (ViT) 99.04 -0.00 95.03 0.90 97.18 0.56 93.81 1.50
+Finetuning (ViT) 99.15 0.23 95.17 1.84 97.57 0.86 93.62 2.61
+Replay (ViT) 99.29 -0.01 96.43 0.34 98.09 0.25 95.66 0.50
+DER (ViT) 99.15 0.21 96.22 0.71 98.08 0.31 95.08 1.13
+LaMAML (ViT) 99.25 0.05 95.58 0.58 97.84 0.26 94.82 0.79
+LwF (ViT) 97.89 1.69 95.09 0.98 95.74 2.15 85.18 9.72
+CCLIS (ViT) 98.81 0.48 90.99 4.49 94.86 2.88 86.82 6.14
+CoP2L (RN50) 96.22 0.56 86.78 2.01 90.89 2.27 88.57 2.46
+Finetuning (RN50) 96.60 1.00 88.23 1.78 93.25 1.27 90.62 1.38
+Replay (RN50) 96.88 0.42 87.61 0.86 92.89 0.47 90.03 1.26
+DER (RN50) 96.57 1.09 88.29 1.72 93.43 1.11 90.66 1.12
+LaMAML (RN50) 96.65 0.73 83.57 6.18 89.89 4.12 87.20 4.53
+LwF (RN50) 95.80 0.33 79.39 0.47 87.73 0.33 85.03 0.89
+CCLIS (RN50) 95.45 0.79 78.55 6.76 83.51 6.60 75.65 14.12
+CSReL (RN50) 65.46 13.66 69.14 15.17 63.52 14.49 50.22 22.34
+All the details about the experiments can be found in Appendix˜E, alongside supplementary Class Incremental experiments on MNIST (LeCun et al., 2010), FashionMNIST (FMNIST) (Xiao et al., 2017) and EMNIST (Cohen et al., 2017).
+Moreover, we present Domain-Incremental (DI) experiments on PermutedMNIST (Goodfellow et al., 2014) and RotatedMNIST (Ben-David et al., 2010).
+We report all the results in Appendix E.3, including the standard deviations and the training times for each method.
+Accuracy and Forgetting.
+We denote $A(t,\theta_{T})$ the accuracy obtained on a task $t$ of a model $f_{\theta_{T}}$ trained on $T$ tasks. We report the average accuracy over $T$ tasks,
+$\overline{A}(\theta_{T})\,{=}\,\frac{1}{T}\sum_{t=1}^{T}A(t,\theta_{T})\,$,
+and the average forgetting at task $T$,
+$\overline{F}(T)=\frac{1}{T-1}\sum_{t=1}^{T-1}[A(t,\theta_{t})-A(t,\theta_{T})]$}\,.
+4.1 Study of our generalization bound
+In Figures 2 and 3, we present our generalization bound on the true risk for each task for CIFAR10 with 5 tasks and CIFAR100 with 20 tasks.
+These bounds are presented for Vision Transformer (ViT) (Dosovitskiy et al., 2021) networks and ResNet50 (RN50) (He et al., 2016) networks trained with a buffer size of 2000 under both Class-Incremental and Task-Incremental settings.
+We present the details for the computation of the bound in Appendix˜B.
+We observe that although the bounds are calculated exclusively on the training set, their values are non-vacuous and follow the test set error trends.
 CoP2L achieves particularly tight generalization bounds in the Task-Incremental setting, where forgetting is inherently reduced due to access to task identities and the use of a separate classifier head for each task.
@@
 This is likely because the ViT backbone provides more structured and information-rich representations, which are easier to compress effectively, thereby enhancing the performance of CoP2L.
-We also give the results for all buffer sizes for the three MNIST datasets, and also bound estimates for the CIFAR10, CIFAR100, and ImageNet datasets (TinyImageNet tiny-imagenet , ImageNette imagenette and ImageWoof Howard_Imagewoof_2019 ) as well as bounds for the Domain-Incremental setting in Appendix D.
-4.2 Analysis of CoP2L
-In this section we provide an analysis of CoP2L with an ablation study on loss terms, a sensitivity analysis with respect to task ordering, and buffer size.
-Due to space limitations the figures are provided in the appendix, however the main results are mentioned here in the text.
-Ablation on CoP2L loss terms  We have conducted an ablation study to understand the effect of additional weighting that counteracts class imbalance that arises in the continual learning setup when new tasks are introduced (Standard replay counteracts this by extending the batch size to have equal class representation).
-In the modified P2L (Algorithm 2), we incorporate an additional weighting term over classes to adapt P2L for class-balanced continual learning.
-In order to assess whether this idea is indeed helpful, we have run an experiment on the MNIST dataset with and without the additional weighting idea.
-In Appendix E, in Figure 10, we show this experiment.
-For both accuracy (left panel) and forgetting (right panel), the inclusion of the additional loss weighting term significantly improves the results.
+Furthermore, the tightness of the bounds improves when the size of the dataset increases, which explains that the bounds on CIFAR10 (5000 samples per class) are noticeably tighter than the bounds on CIFAR100 (about 500 samples per class).
+In addition, the complexity of the incremental scenario plays a role: CIFAR10 with 5 tasks involve learning two classes at a time, whereas CIFAR100 with 10 tasks require learning ten classes simultaneously, further contributing to looser bounds.
+Generalization bounds for multiple combinations of datasets and architectures are reported in Section˜E.4.
+4.2 Study of the algorithm’s performance
+Table˜1 presents results on CIFAR10, CIFAR100 and TinyImageNet datasets, for both Class-Incremental and Task-Incremental settings.
+We observe that in Class-Incremental settings, CoP2L is competitive with most baselines in terms of accuracy while maintaining low forgetting.
+Alongside these baselines, we present results for finetuning a model without a replay buffer, which evidently leads to a drastic drop in performance.
+In Task-Incremental settings, our method remains competitive.
+We note that even simple finetuning performs well in this setting.
+This behavior is expected, since adding a new classification head for each task naturally reduces forgetting.
+All the while being a competitive continual learning approach, CoP2L offers the additional advantage of certifying the generalization loss of the learned predictor.
+Further experiments on Imagenette and Imagewoof (Howard, 2019a) are presented in Appendix E.
+We do not present results for the ViT with CSReL, as the method is designed for small datasets ($\approx$1000 datapoints), and becomes prohibitively more expensive when the dataset and the model grow in size.
+With the ResNet50, running CSReL was 5 to 32 times longer than executing CoP2L, while achieving much worse accuracy.
+Finally, we refer the reader to Appendix F for an extended analysis of CoP2L behavior, including an ablation study on the loss terms, a sensitivity analysis with respect to task ordering, an analysis of the impact of buffer size, an ablation study on the use of early stopping with the bound and an evaluation of memory cost.
+Furthermore, we report experiments that CoP2L is able to mitigate forgetting more than the other methods while still retaining plasticity, and is able to achieve a balanced performance over past tasks.
+5 CONCLUSION
+We proposed CoP2L, an algorithm rooted in the sample compression theory, for self-certified continual learning.
+To the best of our knowledge, this is the first attempt to employ sample compression theory within the continual learning context.
+We provided sample compression bounds for CoP2L and verified empirically that they are non-vacuous and informative.
+We furthermore showed that on several challenging datasets, our approach yields comparable results when compared to several strong baselines, while also providing learning guarantees for trustworthy continual learning.
+As the combination of Pick-to-Learn with experience replay leads to a successful learning scheme, we foresee that combining Pick-to-Learn with other continual learning approaches could lead to new compelling ways of obtaining self-certified predictors.
+Appendix A Table of symbols
+We summarize the notation and their definitions in the following table.
+Table 2: Summary table of the symbols used in this paper.
+Symbol Definition
+$\operatorname{\mathcal{X}}$ A feature space
+$\operatorname{\mathcal{Y}}$ A label space
+$S$ A dataset composed of datapoints in $\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}$
+$\operatorname{\mathcal{D}}$ A distribution over $\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}$
+$\operatorname{\mathfrak{D}}$ A meta-distribution over the space of distributions $\operatorname{\mathcal{D}}$
+$\Theta$ A set of learnable parameters $\theta$
+$f_{\theta}$ A function $f_{\theta}:\operatorname{\mathcal{X}}\to\operatorname{\mathcal{Y}}$ parametrized by $\theta\in\Theta$
+AA An learning algorithm that returns parameters $\theta\in\Theta$
+$\operatorname{\mathbf{i}}$ The indices of the compression set $S^{\operatorname{\mathbf{i}}}$
+$\operatorname{\mathbf{i}}^{{\rm\bf c}}$ The indices of the compression set’s complement $S^{\operatorname{\mathbf{i}}^{{\rm\bf c}}}$
+$\operatorname{|\mathbf{i}|}$ The size of the compression set
+$\operatorname{\mathbf{j}}$ The indices of the second compression set $S^{\operatorname{\mathbf{j}}}$ used by CoP2L
+$\operatorname{\mathscr{P}}(n)$ The powerset of $\{1,\ldots,n\}$
+$\mu$ A message
+$\Sigma$ The alphabet of symbols $\sigma$ used to create the message $\mu$
+$\Sigma^{*}$ The set of all possible sequences of symbols $\sigma\in\Sigma$
+$\mathscr{M}(\operatorname{\mathbf{i}})$ The set of possible messages in $\Sigma^{*}$ given a sequence $\operatorname{\mathbf{i}}$
+$\operatorname{\mathscr{C}}$ The compression function
 […]
-Analysis on buffer size.
-In Appendix E, in Figure 12, we compare the accuracy and forgetting obtained with different replay buffer sizes of (1000, 2000, 3000, 4000, 5000) samples for CoP2L and replay.
-This represents roughly 8% to 40% of the per-task dataset size, for MNIST and Fashion-MNIST, and 10% to 52% for EMNIST.
-For each replay buffer size, we report the variation of accuracy obtained at the end of training averaged over different seeds, and test accuracy for different tasks.
-For the MNIST dataset, on average CoP2L significantly outperforms replay (as evidenced by the variation over seeds).
-For the Fashion-MNIST and EMNIST datasets, when using small buffer sizes, the average accuracy obtained with our CoP2L is significantly better.
-For larger buffer sizes, in general CoP2L performs comparably or slightly better.
-We also note for all three datasets, CoP2L significantly outperforms replay in terms of forgetting.
-We attribute this due to the fact that CoP2L is able to intelligently select examples from the replay buffer.
-Namely, CoP2L is able to focus on examples that are most representative of the earlier tasks, which we observe to help significantly in mitigating the deterioration over earlier tasks as evidenced by the results presented in the right column in Figure 12.
-5 Limitations and Broader Impact Statement
-We observe with our experiments that the performance of CoP2L is dependent to the ability of the feature extractor (backbone).
-In case a weak backbone is employed, CoP2L could necessitate a larger compression set (e.g.
-more data points per task).
-We have also seen that in those cases CoP2L might take longer to converge.
-We noted in our experiments that using deep pretrained networks such as ViT results in good performance, by increasing the compression ability.
-More broadly, we do not believe there is any particular negative social impact to note for our proposed methodology.
-On the contrary, we believe that our algorithm could be instrumental for improving trustworthiness of continual learning / lifelong learning in practice.
-6 Conclusions
-In this paper, we have proposed CoP2L, an algorithm rooted in the sample compression theory for self-certified continual learning.
-To the best of our knowledge, this is the first attempt in the literature to employ sample compression theory within the continual learning context.
-We provided sample compression bounds for CoP2L and verified empirically that they are non-vacuous and informative.
-We furthermore showed that on several challenging datasets, that our approach yields comparable or better results compared several strong baselines, while also providing learning guarantees, for trustworthy continual learning.
-As the combination of Pick-to-Learn with experience replay leads to a successful learning scheme, we foresee that combining Pick-to-Learn with other continual learning approaches could also lead to new compelling ways of obtaining self-certified predictors.
-Acknowledgements
-Mathieu Bazinet is supported by a FRQNT B2X scholarship (343192, doi: https://doi.org/10.69777/343192).
-Pascal Germain, and Cem Subakan are supported by Discovery Grants RGPIN-2020-07223 and RGPIN 2023-05759, respectively.
-This research was also enabled in part by support provided by Calcul Québec (https://www.calculquebec.ca/) and the Digital Research Alliance of Canada (https://alliancecan.ca/en).
-Appendix A Compression and Reconstruction Functions
-In line with the prevalent literature, we introduced the sample compression framework (Section 2) based on a compression function providing a compression set and optionally a message.
+$\Psi$ The sample-compression bound of Equation˜5
+$\ell$ A loss function $\ell:\Theta\times\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}\to[0,1]$
+$\operatorname{\widehat{\mathcal{L}}}_{S}$ The empirical loss over a dataset $S$ given a loss function $\ell$
+$\operatorname{\widehat{\mathcal{L}}}_{S}^{\ \operatorname{\mathbf{i}}}$ The empirical loss over a dataset $S^{\operatorname{\mathbf{i}}}$ given a loss function $\ell$
+$\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}$ The true loss over a distribution $\operatorname{\mathcal{D}}$ given a loss function $\ell$
+Appendix B Compression and Reconstruction Functions
+In line with the prevalent literature, we introduced the sample compression framework based on a compression function providing a compression set and optionally a message.
 In the case of CoP2L, the reconstruction scheme relies on two compression sets $(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}})$ and a message pair $(\mu_{1},\mu_{2})$, which we explain later in this section.
-[^2] The idea of using multiple compression sets appeared in [40] and [39], but in a setting without a message.
+[^3] The idea of using multiple compression sets appeared in Marchand and Shawe-Taylor [2002] and Marchand et al.
+[2003], but in a setting without a message.
 Algorithm 4 Compression function of Continual Pick-To-Learn (CoP2L)
+0:  $\theta_{0}$ {Initialization parameters of the model}
 0:  $S_{1},S_{2},\ldots,S_{T}$ {Training sets}
-0:  $\theta_{0}$Initialization parameters of a given model}
 0:  $\gamma$ {P2L’s stopping criteria}
@@
 19:  return $\theta_{T},\{S^{\operatorname{\mathbf{i}}}_{t}\}_{t=1}^{T},\{S^{\operatorname{\mathbf{j}}}_{t}\}_{t=1}^{T},\{\mu_{1,t}\}_{t=1}^{T},\mu_{2}$.
-{Learned parameters, compression sets and messages sets}
-A.1 Compression Function
-Two challenges prevent CoP2L (as presented by Algorithm 3) from being used as its own compression and reconstruction function.
-In this subsection, we expose each of these challenges while explaining how we the proposed compression algorithm (Algorithm 4) allows to circumvent them.
+{Learned parameters, compression sets and message sets}
+B.1 Compression Function
+Two challenges prevent CoP2L (as presented by Algorithm 3) from being used as its own compression and reconstruction function, as it is done when working with the original P2L algorithm in a standard (non continual learning) setting.
+In this subsection, we expose each of these challenges while explaining how the proposed compression algorithm (Algorithm 4) allows to circumvent them.
 The first challenge comes from the sampling of the buffer $B^{\star}$ (Lines 6 to 8 of Algorithm 3).
@@
 Therefore, the reconstruction function needs to know when to remove the datapoint.
-The message $\mu_{1}$ returned by Algorithm 4 thus provides the task index after which the datapoint was removed during CoP2L execution.
+The message $\mu_{1}$ returned by Algorithm 4 thus provides the task index after which the datapoint was removed during the execution of CoP2L.
 As the reconstruction function only requires a message for datapoints who were removed from the buffer after being added to the compression set $S^{\operatorname{\mathbf{i}}}$, a second compression set $S^{\operatorname{\mathbf{j}}}$ is dedicated to datapoints that necessitate a message.
@@
 Recall that the reconstruction function does not have access to the whole dataset.
-Thus, it cannot recover the bound value computed during the initial training phase (see Line 11 of Algorithm 2) and use it as a stopping criterion.
+Thus, it cannot recover the bound value computed during the initial training phase (see Line 11 of Algorithm˜2) and use it as a stopping criterion.
 To address this inconvenience, the number of iterations to perform is provided to mP2L as a message $\mu_{2}=(\mu_{2}^{1},\ldots,\mu_{2}^{T})$.
@@
 After learning on $T$ tasks, the compression function provides the sample compression set $S^{\operatorname{\mathbf{i}}}$ and  $S^{\operatorname{\mathbf{j}}}$, along with a message pair $(\mu_{1},\mu_{2})$ chosen among the set of all possible messages, denoted as
-$\mathscr{M}(\operatorname{\mathbf{j}},T)\,=\,\{2,\ldots,T\}^{|\operatorname{\mathbf{j}}|}\times\left[\textstyle\bigtimes_{t=1}^{T}\{1,\ldots,n_{t}+|{\mathcal{B}}|\}\right].$ (7)
-A.2 Reconstruction Function
+$\mathscr{M}_{1:T}(\operatorname{\mathbf{j}})\,=\,\{2,\ldots,T\}^{|\operatorname{\mathbf{j}}|}\times\left[\textstyle\bigtimes_{t=1}^{T}\{1,\ldots,n_{t}+|{\mathcal{B}}|\}\right].$ (6)
+B.2 Reconstruction Function
 Under the assumption that the input parameters $\theta_{0},\gamma,m,\omega$ are the same for the compression function (Algorithm 4) and for CoP2L (Algorithm 3), the composition of the compression and the reconstruction function must output the exact same predictor as CoP2L.
 However, it only has access to the compression sets and the messages.
-The first challenge of the compression function is addressed using $\mu_{1}$ from line 6 to 11 of Algorithm 5.
+The first challenge of the compression function is addressed using $\mu_{1}$ from line 6 to 11 of Algorithm˜5.
 For each datasets, we verify if a datapoint was previously excluded from the buffer by the sampling function.
 If so, we remove it from the buffer.
-The second challenge is addressed in line 4 of Algorithm 5 by giving the message $\mu_{2}$ to mP2L as stopping criterion.
+The second challenge is addressed in line 4 of Algorithm˜5 by giving the message $\mu_{2}$ to mP2L as stopping criterion.
 Algorithm 5 Reconstruction function of Continual Pick-To-Learn (CoP2L)
+0:  $\theta_{0}$ {Initialization parameters of the model}
 0:  $S^{\operatorname{\mathbf{i}}}_{1},S^{\operatorname{\mathbf{i}}}_{2},\ldots,S^{\operatorname{\mathbf{i}}}_{T}$ {Compression sets}
@@
 0:  $\mu_{2}${Message set}
-0:  $\theta_{0}$Initialization parameters of a given model}
 0:  $\gamma$ {P2L’s stopping criteria}
@@
 15:  return $\theta_{T}$. {Learned parameters}
-A.3 Computation of the Bound
-To compute the bound of Theorem 3.1, we need three things : the size of the first compression set $\operatorname{\mathbf{i}}$, the size of the second compression set $\operatorname{\mathbf{j}}$ and the message $\mu$.
+B.3 Computation of the Bound
+To compute the bound of Theorem˜3.1, we need three things : the size of the first compression set $\operatorname{\mathbf{i}}$, the size of the second compression set $\operatorname{\mathbf{j}}$ and the message $\mu$.
 By simply running our CoP2L, we are able to retrieve $|\operatorname{\mathbf{i}}|+|\operatorname{\mathbf{j}}|$, but not the size of each compression set individually.
 Moreover, we are completely unable to retrieve the messages.
-Thus, instead of implementing Algorithm 3, we actually implemented the compression function as described in Algorithm 4.
-Moreover, we use a block version of P2L, as defined by Algorithm 2 of [23].
-We thus add $k$ datapoints at once to the compression set at Line 6 of Algorithm 2 (the used values of $k$ for each experiments is given in Appendix G.
-Thus, the value of $\mu_{2}^{t}$ (Line 6 of Algorithm 4) becomes the number of iterations of the mP2L algorithm at for task $t$.
+Thus, instead of implementing Algorithm˜3, we actually implemented the compression function as described in Algorithm˜4.
+Moreover, we use a block version of P2L, as defined by Algorithm 2 of Paccagnan et al.
+[2024].
+We thus add $k$ datapoints at once to the compression set at Line 6 of Algorithm 2 (the used values of $k$ for each experiments is given in Section˜E.2.
+Thus, the value of $\mu_{2}^{t}$ (Line 6 of Algorithm˜4) becomes the number of iterations of the mP2L algorithm at for task $t$.
 Knowing that mP2L outputs a compression set $C^{\star}_{t}$, and adds $k$ datapoints to the compression set at each iteration, the number of iterations is $\mu_{2}^{t}=\frac{|C^{\star}_{t}|}{k}$.
-Finally, when computing the bound, we use the code provided by [65] to invert the $\mathrm{kl}$ divergence.
-Appendix B Tighter sample compression $\mathrm{kl}$ bound
-In this section, we prove a tighter $\mathrm{kl}$ bound that is not derived from Theorem 2.1 but proved in a very similar way.
-To do so, we start by providing the Chernoff test-set bound [66] for losses in $[0,1]$ [67].
-Theorem B.1 ([67]).
+Finally, when computing the bound, we use the code provided by Viallard et al.
+[2021] to invert the $\mathrm{kl}$ divergence.
+Appendix C Tighter sample compression $\mathrm{kl}$ bound
+In this section, we prove a tighter $\mathrm{kl}$ bound that is not derived from Theorem˜2.1 but proved in a very similar way.
+To do so, we start by providing the Chernoff test-set bound for losses in $[0,1]$ [Langford, 2005, Foong et al., 2022].
+Theorem C.1 (Foong et al., 2022).
 Let $X_{1},\ldots,X_{n}$ be i.i.d. random variables with $X_{i}\in[0,1]$ and $\operatorname*{\mathbb{E}}[X_{i}]=p$. Then, for any $\delta\in(0,1]$, with probability at least $1-\delta$
-$p\leq\mathrm{kl}^{-1}\quantity(\tfrac{1}{n}\textstyle\sum_{i=1}^{n}X_{i},\tfrac{1}{n}\log\tfrac{1}{\delta}).$
+[…]
 We now prove a tighter $\mathrm{kl}$ sample compression bound.
 As our algorithm CoP2L needs two compression sets, we consider a second compression set $\operatorname{\mathbf{j}}$.
+To use these two compression sets, we need to first redefine the reconstruction function
+$\operatorname{\mathscr{R}}:\boldsymbol{\cup}_{m\leq n}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{m}\times\boldsymbol{\cup}_{k\leq n}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{k}\times\boldsymbol{\cup}_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\mathscr{M}(\operatorname{\mathbf{j}})\to\Theta.$
 We define a conditional probability distribution $P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})$ that incorporates the knowledge that a vector $\operatorname{\mathbf{i}}$ was already drawn from $\operatorname{\mathscr{P}}(n)$ and that $\operatorname{\mathbf{i}}\cap\operatorname{\mathbf{j}}=\emptyset$.
 Thus, we have $\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})\leq 1.$ If the choice of $\operatorname{\mathbf{j}}$ isn’t conditional to the choice of $\operatorname{\mathbf{i}}$, we simply have $P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})=P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}})$.
-Theorem B.2.
-For any distribution $\operatorname{\mathcal{D}}$ over $\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}$, for any family of set of messages $\{\mathscr{M}(\operatorname{\mathbf{j}})\,|\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)\}$, for any deterministic reconstruction function $\operatorname{\mathscr{R}}$ , for any loss $\ell:\Theta\times\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}\to[0,1]$ and for any $\delta\in(0,1]$, with probability at least $1-\delta$ over the draw of $S\sim\operatorname{\mathcal{D}}^{n}$, we have
+Theorem C.2.
+[…]
 $\forall\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n),\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n),\mu\in\mathscr{M}(\operatorname{\mathbf{j}}):$
-$\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta})$
+$\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\ln\frac{1}{P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta})$
 Proof.
-We prove the complement of the expression in Theorem B.2. Denote the upper bound
-$U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)=\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta})$
+Let us prove the complement of the expression in Theorem˜C.2. Denote the upper bound
+$U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)\ =\ \mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\ln\frac{1}{P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta})\,.$
+We have
 $\operatorname*{\mathbb{P}}_{S\sim\operatorname{\mathcal{D}}^{n}}\quantity(\exists\operatorname{\mathbf{i}},\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n),\mu\in\mathscr{M}(\operatorname{\mathbf{j}}):\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))>U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta))$
@@
 $\leq\delta.$
-We use the union bound in Equation 8 and Equation 9.
-In Equation 10, we use the i.i.d.
-assumption.
-In Equation 11 we use the Chernoff Test-set bound of Theorem B.1 with $p=\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))=\operatorname*{\mathbb{E}}_{(\boldsymbol{x},y)\sim\operatorname{\mathcal{D}}}\ell(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu),\boldsymbol{x},y)$ and $X_{i}=\ell(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu),\boldsymbol{x}_{i},y_{i})\forall i\in\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}$.
-Finally, the last line is derived from the definition of $P_{\operatorname{\mathscr{P}}(n)}$ and $P_{\mathscr{M}(\operatorname{\mathbf{j}})}$.
+Equations˜7 and 8 use the union bound, Equation˜9 uses the i.i.d.
+assumption, and Equation˜10 uses the Chernoff Test-set bound of Theorem˜C.1 with $p=\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))=\operatorname*{\mathbb{E}}_{(\boldsymbol{x},y)\sim\operatorname{\mathcal{D}}}\ell(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu),\boldsymbol{x},y)$ and $X_{i}=\ell(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu),\boldsymbol{x}_{i},y_{i})\forall i\in\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}$.
+Finally, the last inequality is obtained from the definition of $P_{\operatorname{\mathscr{P}}(n)}$ and $P_{\mathscr{M}(\operatorname{\mathbf{j}})}$.
 ∎
-Appendix C Proof of Theorem 3.1
+Appendix D Proof of Theorem 3.1
 See 3.1
 Proof.
-Let us choose $t\in[1,T]$.
-Let us sample $S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}$.
-We denote the reconstruction function of CoP2L as $\operatorname{\mathscr{R}}_{1:T}(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)=\operatorname{\mathscr{R}}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu;S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}).$
+Let us choose $t\in[1,T]$. Let us sample $S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}$.
 Let
-$U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)=\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n_{t}-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta}).$
-Then, we have :
-$\operatorname*{\mathbb{P}}_{S^{t}}\quantity(\forall\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n_{t}),\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n_{t}),\mu\in\mathscr{M}(\operatorname{\mathbf{j}}):\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta))$
+$U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)=\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\theta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}^{(t)}),\frac{1}{n_{t}-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\ln\frac{1}{P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta}).$
+Then, we have
+$\operatorname*{\mathbb{P}}_{S^{t}}\quantity(\forall\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n_{t}),\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n_{t}),\mu\in\mathscr{M}(\operatorname{\mathbf{j}}):\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\theta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}^{(t)})\leq U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)).$
 As all the datasets (except $S^{t}$) are sampled beforehand, we can define $\operatorname{\mathscr{R}}_{1:T}(\cdot)=\operatorname{\mathscr{R}}\quantity(\cdot;S^{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T})$ before drawing $S^{t}$.
-Thus, the reconstruction function is only a function of the dataset $S^{t}$, which is the setting of the result of Theorem B.2.
+Thus, the reconstruction function is only a function of the dataset $S^{t}$, which is the setting of the result of Theorem˜C.2.
 We can then lower bound this probability by $1-\delta$.
 We finish the proof by applying the bound to all datasets. We have
-$\operatorname*{\mathbb{P}}_{S_{1},\ldots,S_{T}}\Biggl{(}\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)\Biggr{)}$
-$=$ $\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{T}}\operatorname{\mathbb{I}}\Bigg{(}\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)\Biggr{)}$
-$=$ $\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}}\operatorname*{\mathbb{E}}_{S_{t}}\operatorname{\mathbb{I}}\Bigg{(}\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)\Biggr{)}$
-$=$ $\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}}\operatorname*{\mathbb{P}}_{S_{t}}\Bigg{(}\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)\Biggr{)}$
+$\operatorname*{\mathbb{P}}_{S_{1},\ldots,S_{T}}\Biggl(\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\theta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}^{(t)})\leq U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)\Biggr)$
+$=$ $\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{T}}\operatorname{\mathbb{I}}\Bigg(\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\theta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}^{(t)})\leq U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)\Biggr)$
+$=$ $\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}}\operatorname*{\mathbb{E}}_{S_{t}}\operatorname{\mathbb{I}}\Bigg(\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\theta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}^{(t)})\leq U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)\Biggr)$
+$=$ $\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}}\operatorname*{\mathbb{P}}_{S_{t}}\Bigg(\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\theta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}^{(t)})\leq U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)\Biggr)$
 $\geq$ $\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}}1-\delta$
@@
 As this theorem holds specifically for CoP2L, we specify the distributions $P_{\operatorname{\mathscr{P}}(n_{t})}$ and $P_{\mathscr{M}(\operatorname{\mathbf{j}})}$.
-Following the work of [39], which used sample compression bounds with three compression sets, with $\zeta(k)=\frac{6}{\pi}(k+1)^{-1}$, we choose
+Following the work of Marchand et al.
+[2003], which used sample compression bounds with three compression sets, with $\zeta(k)=\frac{6}{\pi}(k+1)^{-1}$, we choose
 $P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{i}})=\matrixquantity(n_{t}\\ \operatorname{|\mathbf{i}|})^{-1}\zeta(\operatorname{|\mathbf{i}|})\quad\mbox{ and }\quad P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})=\matrixquantity(n_{t}-\operatorname{|\mathbf{i}|}\\ |\operatorname{\mathbf{j}}|)^{-1}\zeta(|\operatorname{\mathbf{j}}|)\,.$
@@
 This message indicates the task number where a datapoint is sampled out of the buffer.
-Thus, the size of $\Sigma_{1}$ is $T-1$ and the probability of a symbol is $\frac{1}{T-1}$.
+Thus, the size of $\Sigma_{1}$ is T1T-1 and the probability of a symbol is $\frac{1}{T-1}$.
 For any vector $\operatorname{\mathbf{j}}$, we have a sequence of length $|\operatorname{\mathbf{j}}|$, and thus the probability of choosing each sequence is $(\frac{1}{T-1})^{|\operatorname{\mathbf{j}}|}$.
-The second message $\mu_{2}$ is defined using the alphabet $\Sigma_{2}=\{1,\ldots,n_{t}\}$.
+The second message $\mu_{2}$ is defined using the alphabet $\Sigma_{2}=\{1,\ldots,n_{t}+|{\mathcal{B}}|\}$.
 We choose to use $\zeta(k)=\frac{6}{\pi^{2}}(k+1)^{-2}$ as the probability distribution over each character.
@@
 Thus, we have the probability of a sequence $\mu_{2}=\mu_{2}^{1}\ldots\mu_{2}^{T}$ is $\prod_{i=1}^{T}\zeta(\mu_{2}^{i})$.
-We define $\mathscr{M}(\operatorname{\mathbf{j}},T)$ the set of messages such that the sequence of symbols from $\Sigma_{1}$ is of length $|\operatorname{\mathbf{j}}|$ and the sequence of symbols from $\Sigma_{2}$ is of length $T$ (Equation 7).
+We define $\mathscr{M}(\operatorname{\mathbf{j}},T)$ the set of messages such that the sequence of symbols from $\Sigma_{1}$ is of length $|\mathbf{j}|$ and the sequence of symbols from $\Sigma_{2}$ is of length $T$ (Equation 6).
 ∎
-Appendix D Bound Plots
-D.1 Bound plots for different buffer sizes
-In this section, we present the computed bound values over task risks on MNIST, Fashion-MNIST and EMNIST datasets (letters).
-In Figure 2, we present results for different replay buffer sizes (1000, 2000, 3000, 4000, 5000), obtained using the MLP architecture.
-We also show the results obtained with the CNN architecture in Figure 3.
-We report the estimated deviation of the bound values over different seeds using the shaded regions which show the $\pm$1 standard deviation, and also the whiskers (plotted using the seaborn toolkit’s boxplot function) which show the spread of the values excluding the outliers.
-This plotting style for showing the deviation across seeds applies to all bound plots presented in this paper.
-Figure 2: CoP2L Bounds with respect to tasks on the MNIST, Fashion-MNIST and EMNIST datasets using an MLP architecture
-Figure 3: CoP2L Bounds with respect to tasks on the MNIST, Fashion-MNIST and EMNIST datasets using a CNN architecture
-We observe from Figures 2 and 3 that the MLP architecture and the CNN architecture give non-trivial bounds that follow the trends exhibited by the test error curve.
-D.2 Bound plots for CIFAR10, CIFAR100, TinyImageNet and ImageNet (subsets) datasets
-In this section, we present the plots for the bound values obtained for the additional datasets, namely CIFAR10, CIFAR100, and Imagenet subsets.
-The Imagenet subsets include the Tinyimagenet [51], ImageWoof [64], and ImageNette [63] datasets.
-We provide the results on ImageWoof and ImageNette to showcase that the estimated bounds get tighter when more data is avaible per each task.
-Figure 4: CoP2L bounds with respect to tasks on the CIFAR10 (5 tasks) dataset in Class-Incremental (left) and Task-Incremental (right) settings
-Figure 5: CoP2L bounds with respect to tasks on the CIFAR100 (10 tasks) dataset in Class-Incremental (left) and Task-Incremental (right) settings
-Figure 6: CoP2L bounds with respect to tasks on the CIFAR100 (20 tasks) dataset in Class-Incremental (left) and Task-Incremental (right) settings
-Figure 7: CoP2L bounds with respect to tasks on the TinyImageNet (40 tasks) dataset in Class-Incremental (left) and Task-Incremental (right) settings
-Figure 8: CoP2L bounds with respect to tasks on the ImageNette (5 tasks) dataset (left) and ImageWoof dataset (right) in Task-Incremental setting
-Figure 9: CoP2L bounds with respect to tasks on the PermutedMNIST (10 tasks) dataset (left) and on the RotatedMNIST dataset in Domain-Incremental setting
-Appendix E Analysis of CoP2L
-In this section we provide an analysis of CoP2L with an ablation study on loss terms, a sensitivity analysis with respect to task ordering, and buffer size.
-The text matches the one provided in the main paper, but we also include the aforementioned figures.
-Ablation on CoP2L loss terms.
-We have conducted an ablation study to understand the effect of additional weighting that counteracts class imbalance that arises in the continual learning setup when new tasks are introduced (Standard replay counteracts this by extending the batch size to have equal class representation).
-In the modified P2L (Algorithm 2), we incorporate an additional weighting term over classes to adapt P2L for class-balanced continual learning.
-In order to assess whether this idea is indeed helpful, we have run an experiment on the MNIST dataset with and without the additional weighting idea.
-In Figure 10, we show the results of this experiment.
-For both accuracy (left panel) and forgetting (right panel), we observe that the inclusion of the additional loss weighting term significantly improves the results.
-[…]
-Figure 10: Ablation study on MNIST and Fashion-MNIST on the additional weighting to ensure class balance we propose in Algorithm 2 on the CoP2L loss function.
-Accuracy (left) and forgetting (right) over replay buffer size for the MNIST dataset
-Figure 11: Studying the variation of cumulative accuracy with respect to task ordering for CoP2L and Replay on MNIST dataset (left) and Fashion-MNIST dataset (right).
-We use 5 different task order, and show the variability with respect to each task.
-E.1 Analysis with respect to buffersize
-In Figure 12, we compare the accuracy and forgetting obtained with different replay buffer sizes of (1000, 2000, 3000, 4000, 5000) samples for CoP2L and replay.
-This represents roughly 8% to 40% of the per-task dataset size, for MNIST and Fashion-MNIST, and 10% to 52% for EMNIST.
-For each replay buffer size, we report the variation of accuracy obtained at the end of training averaged over different seeds, and test accuracy for different tasks.
-Note that in Figure 12, we report the average accuracy obtained after the final task, such that $T=5$ for MNIST and Fashion-MNIST and $T=13$ for EMNIST.
-We observe that for the MNIST dataset, on average CoP2L significantly outperforms replay (as evidenced by the variation over seeds).
-For the Fashion-MNIST and EMNIST datasets, we observe that for small buffer sizes, the average accuracy is obtained with our CoP2L is significantly better.
-For larger buffer sizes, in general CoP2L performs comparably or slightly better.
-We would like to also note for all three datasets, CoP2L significantly outperforms replay in terms of forgetting.
-We attribute this due to the fact that CoP2L is able to intelligently select examples from the replay buffer.
-Namely, CoP2L is able to focus on examples that are most representative of the earlier tasks, which we observe to help significantly in mitigating the deterioration over earlier tasks as evidenced by the results presented in the right column in Figure 12.
-Figure 12: Accuracy (left) and forgetting (right) over replay buffer size for (top row) MNIST dataset, (middle row) Fashion-MNIST dataset, (bottom row) EMNIST dataset.
-The orange triangles show the average performance for the corresponding box plot.
-Appendix F All results with standard deviation and training time measurement
-In this section we provide the average accuracy values obtained on MNIST, FMNIST, EMNIST, CIFAR10, CIFAR100, TinyImagenet, PermutedMNIST and RotatedMNIST datasets with the standard deviation estimates.
-Note that Table 2 of the main paper provides the results obtained with a single seed, whereas the following tables provide the average results over five seeds, and the corresponding standard deviations.
-Note that we use different random number generator seeds to account for variability in random number generations (e.g.
-model parameter initializations) – but train / test splits are kept fixed with respect to seeds across different experiments.
-Note that, for CIFAR10, CIFAR100, TinyImagenet, ImageWoof, ImageNette, PermutedMNIST and RotatedMNIST experiments we also report the total time it takes to complete each experiment (for all tasks).
-CIFAR10 (5 tasks) Class-Incremental
-Methods Stream Accuracy (%) Stream Forgetting (%) Total time (s)
-CoP2L-Dino (ViT) $94.45\pm 0.14$ $2.10\pm 0.28$ $118.07\pm 2.59$
-Replay-Dino (ViT) $94.00\pm 0.19$ $6.11\pm 0.25$ $486.25\pm 43.07$
-iCaRL-Dino (ViT) $88.60\pm 10.04$ $3.13\pm 4.20$ $884.59\pm 1.39$
-GDumb-Dino (ViT) $94.16\pm 0.24$ $3.16\pm 0.39$ $47.72\pm 0.12$
-CoP2L-Dino (ResNet50) $80.98\pm 0.17$ $5.84\pm 0.37$ $527.95\pm 8.43$
-Replay-Dino (ResNet50) $81.38\pm 0.39$ $17.72\pm 0.51$ $287.38\pm 25.06$
-iCaRL-Dino (ResNet50) $74.49\pm 0.22$ $1.79\pm 0.43$ $579.17\pm 2.39$
-GDumb-Dino (ResNet50) $80.84\pm 0.14$ $9.35\pm 0.30$ $31.98\pm 0.12$
-CIFAR100 (10 tasks) Class-Incremental
-Methods Stream Accuracy (%) Stream Forgetting (%) Total time (s)
-CoP2L-Dino (ViT) $73.49\pm 0.27$ $17.06\pm 0.29$ $522.18\pm 3.90$
-Replay-Dino (ViT) $68.71\pm 0.25$ $30.02\pm 0.31$ $453.04\pm 11.93$
-iCaRL-Dino (ViT) $65.54\pm 2.23$ $13.64\pm 0.64$ $2491.81\pm 2889.82$
-GDumb-Dino (ViT) $72.51\pm 0.26$ $11.53\pm 0.65$ $95.43\pm 0.34$
-CoP2L-Dino (ResNet50) $49.53\pm 0.24$ $22.31\pm 0.20$ $1724.15\pm 23.50$
-Replay-Dino (ResNet50) $49.41\pm 0.38$ $42.12\pm 0.54$ $295.80\pm 15.88$
-iCaRL-Dino (ResNet50) $46.06\pm 0.40$ $8.50\pm 0.28$ $676.76\pm 56.25$
-GDumb-Dino (ResNet50) $43.37\pm 0.24$ $20.23\pm 0.42$ $63.58\pm 0.36$
-CIFAR100 (20 tasks) Class-Incremental
-Methods Stream Accuracy (%) Stream Forgetting (%) Total time (s)
-CoP2L-Dino (ViT) $70.56\pm 0.99$ $21.15\pm 2.10$ $820.66\pm 176.27$
-Replay-Dino (ViT) $69.21\pm 0.26$ $29.52\pm 0.25$ $457.46\pm 2.07$
-iCaRL-Dino (ViT) $62.36\pm 1.46$ $15.88\pm 0.84$ $1359.13\pm 4.58$
-GDumb-Dino (ViT) $72.35\pm 0.30$ $11.77\pm 0.40$ $190.80\pm 0.42$
-CoP2L-Dino (ResNet50) $46.74\pm 1.63$ $22.54\pm 0.92$ $1771.12\pm 757.50$
-Replay-Dino (ResNet50) $49.44\pm 0.30$ $44.96\pm 0.36$ $293.25\pm 1.77$
-iCaRL-Dino (ResNet50) $46.47\pm 0.33$ $6.74\pm 0.31$ $858.11\pm 120.41$
-GDumb-Dino (ResNet50) $43.60\pm 0.38$ $19.93\pm 0.56$ $126.74\pm 0.46$
-TinyImageNet (40 tasks) Class-Incremental
-Methods Stream Accuracy (%) Stream Forgetting (%) Total time (s)
-CoP2L-Dino (ViT) $51.59\pm 0.24$ $36.32\pm 0.29$ $1451.93\pm 4.71$
-Replay-Dino (ViT) $47.49\pm 0.47$ $48.39\pm 0.49$ $1128.97\pm 16.95$
-iCaRL-Dino (ViT) $21.06\pm 1.11$ $34.49\pm 9.37$ $3919.23\pm 2170.08$
-GDumb-Dino (ViT) $53.93\pm 0.34$ $15.11\pm 0.19$ $471.23\pm 17.85$
-CoP2L-Dino (ResNet50) $33.30\pm 0.32$ $35.93\pm 0.27$ $3671.69\pm 29.19$
-Replay-Dino (ResNet50) $28.38\pm 0.42$ $63.36\pm 0.44$ $882.12\pm 180.18$
-iCaRL-Dino (ResNet50) $33.89\pm 0.47$ $10.42\pm 0.44$ $1905.26\pm 295.18$
-GDumb-Dino (ResNet50) $24.29\pm 0.58$ $21.29\pm 0.56$ $332.22\pm 0.98$
-CIFAR10 (5 tasks) Task-Incremental
-Methods Stream Accuracy (%) Stream Forgetting (%) Total time (s)
-CoP2L-Dino (ViT) $99.04\pm 0.05$ $-0.00\pm 0.10$ $64.51\pm 0.68$
-Replay-Dino (ViT) $99.29\pm 0.04$ $-0.01\pm 0.03$ $217.83\pm 0.76$
-LaMAML-Dino (ViT) $99.25\pm 0.06$ $0.05\pm 0.07$ $344.65\pm 1.46$
-LwF-Dino (ViT) $97.89\pm 0.41$ $1.69\pm 0.48$ $165.41\pm 0.75$
-CoP2L-Dino (ResNet50) $96.22\pm 0.18$ $0.56\pm 0.28$ $172.51\pm 4.42$
-Replay-Dino (ResNet50) $96.88\pm 0.06$ $0.42\pm 0.13$ $144.57\pm 0.51$
-LaMAML-Dino (ResNet50) $96.65\pm 0.21$ $0.73\pm 0.21$ $310.51\pm 1.64$
-LwF-Dino (ResNet50) $95.80\pm 0.15$ $0.33\pm 0.20$ $97.09\pm 0.21$
-CIFAR100 (10 tasks) Task-Incremental
-Methods Stream Accuracy (%) Stream Forgetting (%) Total time (s)
-CoP2L-Dino (ViT) $95.03\pm 0.08$ $0.90\pm 0.11$ $290.95\pm 30.98$
-Replay-Dino (ViT) $96.43\pm 0.14$ $0.34\pm 0.17$ $235.81\pm 1.10$
-LaMAML-Dino (ViT) $95.58\pm 0.16$ $0.58\pm 0.12$ $397.78\pm 3.40$
-LwF-Dino (ViT) $95.09\pm 0.10$ $0.98\pm 0.10$ $173.57\pm 0.63$
-CoP2L-Dino (ResNet50) $86.78\pm 0.16$ $2.01\pm 0.17$ $435.10\pm 7.74$
-Replay-Dino (ResNet50) $87.61\pm 0.05$ $0.86\pm 0.16$ $157.05\pm 0.14$
-LaMAML-Dino (ResNet50) $83.57\pm 0.43$ $6.18\pm 0.53$ $357.24\pm 4.58$
-LwF-Dino (ResNet50) $79.39\pm 0.13$ $0.47\pm 0.11$ $101.27\pm 0.48$
-CIFAR100 (20 tasks) Task-Incremental
-Methods Stream Accuracy (%) Stream Forgetting (%) Total time (s)
-CoP2L-Dino (ViT) $97.18\pm 0.13$ $0.56\pm 0.18$ $226.65\pm 21.43$
-Replay-Dino (ViT) $98.09\pm 0.09$ $0.25\pm 0.12$ $250.78\pm 1.45$
-LaMAML-Dino (ViT) $97.84\pm 0.10$ $0.26\pm 0.14$ $448.43\pm 2.81$
-LwF-Dino (ViT) $95.74\pm 0.19$ $2.15\pm 0.21$ $180.36\pm 0.98$
-CoP2L-Dino (ResNet50) $90.89\pm 0.22$ $2.27\pm 0.28$ $278.67\pm 6.20$
-Replay-Dino (ResNet50) $92.89\pm 0.18$ $0.47\pm 0.24$ $167.65\pm 1.26$
-LaMAML-Dino (ResNet50) $89.89\pm 0.75$ $4.12\pm 0.78$ $407.84\pm 7.38$
-LwF-Dino (ResNet50) $87.73\pm 0.18$ $0.33\pm 0.23$ $105.46\pm 0.49$
-TinyImageNet (40 tasks) Task-Incremental
-Methods Stream Accuracy (%) Stream Forgetting (%) Total time (s)
-CoP2L-Dino (ViT) $93.81\pm 0.26$ $1.50\pm 0.28$ $1039.14\pm 146.21$
-Replay-Dino (ViT) $95.66\pm 0.08$ $0.50\pm 0.09$ $632.18\pm 2.97$
-LaMAML-Dino (ViT) $94.82\pm 0.23$ $0.79\pm 0.19$ $1527.24\pm 891.43$
-LwF-Dino (ViT) $85.18\pm 0.76$ $9.72\pm 0.71$ $430.63\pm 2.14$
-CoP2L-Dino (ResNet50) $88.57\pm 0.17$ $2.46\pm 0.19$ $890.94\pm 3.44$
-Replay-Dino (ResNet50) $90.03\pm 0.26$ $1.26\pm 0.27$ $514.42\pm 102.82$
-LaMAML-Dino (ResNet50) $87.20\pm 0.34$ $4.53\pm 0.28$ $1268.50\pm 551.93$
-LwF-Dino (ResNet50) $85.03\pm 0.17$ $0.89\pm 0.21$ $284.79\pm 5.67$
-ImageNette (5 tasks) Task-Incremental
-Methods Stream Accuracy (%) Stream Forgetting (%) Total time (s)
-CoP2L-Dino (ViT) $98.22\pm 0.19$ $-0.43\pm 0.12$ $118.71\pm 24.35$
-ImageWoof (5 tasks) Task-Incremental
-Methods Stream Accuracy (%) Stream Forgetting (%) Total time (s)
-CoP2L-Dino (ViT) $93.34\pm 0.30$ $-0.65\pm 0.40$ $246.77\pm 13.43$
-Permuted-MNIST (10 tasks) Domain-Incremental
-Methods Stream Accuracy (%) Stream Forgetting (%) Total time (s)
-CoP2L $91.61\pm 0.39$ $6.69\pm 0.40$ $1381.06\pm 1232.99$
-Replay $91.28\pm 0.18$ $6.75\pm 0.21$ $1128.99\pm 49.82$
-LFL $82.76\pm 1.23$ $10.15\pm 1.36$ $92.86\pm 0.24$
-EWC $82.53\pm 0.82$ $6.50\pm 0.88$ $340.16\pm 3.81$
-SI $85.85\pm 0.61$ $11.69\pm 0.67$ $319.95\pm 1.38$
-Rotated-MNIST (10 tasks) Domain-Incremental
-Methods Stream Accuracy (%) Stream Forgetting (%) Total time (s)
-CoP2L $89.40\pm 1.00$ $8.71\pm 1.04$ $4970.86\pm 1452.34$
-Replay $88.75\pm 0.67$ $9.53\pm 0.72$ $3894.62\pm 75.11$
-LFL $53.09\pm 7.06$ $44.18\pm 7.93$ $387.61\pm 160.54$
-EWC $55.59\pm 5.57$ $36.58\pm 6.27$ $1133.19\pm 18.10$
-SI $59.90\pm 6.48$ $41.38\pm 7.18$ $1043.86\pm 7.28$
-MNIST (5 tasks) Class-Incremental
-Methods Stream Accuracy (%) Stream Forgetting (%)
-CoP2L (MLP) $95.12\pm 0.18$ $3.35\pm 0.24$
-Replay (MLP) $94.48\pm 0.28$ $6.08\pm 0.32$
-GDumb (MLP) $91.70\pm 0.31$ $3.84\pm 0.29$
-CoP2L (CNN) $96.61\pm 0.07$ $2.16\pm 0.28$
-Replay (CNN) $96.93\pm 0.29$ $3.31\pm 0.38$
-GDumb (CNN) $94.41\pm 0.13$ $2.97\pm 0.46$
-Fashion-MNIST (5 tasks) Class-Incremental
-Methods Stream Accuracy (%) Stream Forgetting (%)
-CoP2L (MLP) $84.23\pm 0.27$ $12.10\pm 0.66$
-Replay (MLP) $83.99\pm 0.52$ $15.92\pm 0.42$
-GDumb (MLP) $81.76\pm 0.36$ $10.44\pm 0.50$
-CoP2L (CNN) $87.35\pm 0.15$ $11.12\pm 0.36$
-Replay (CNN) $87.59\pm 0.11$ $13.32\pm 0.29$
-GDumb (CNN) $84.98\pm 0.21$ $8.78\pm 0.41$
-EMNIST (13 tasks) Class-Incremental
-Methods Stream Accuracy (%) Stream Forgetting (%)
-CoP2L (MLP) $78.03\pm 0.17$ $16.67\pm 0.31$
-Replay (MLP) $77.22\pm 0.44$ $22.08\pm 0.50$
-GDumb (MLP) $62.64\pm 0.20$ $15.96\pm 0.58$
-CoP2L (CNN) $87.90\pm 0.41$ $8.38\pm 0.54$
-Replay (CNN) $87.70\pm 0.50$ $12.29\pm 0.49$
-GDumb (CNN) $85.05\pm 0.27$ $7.50\pm 0.61$
-Appendix G Training, architecture details and hyperparameter settings for the experiments
-[…]
-In experiments using a Dino backbone (both ResNet50 and ViT backbones), we upsampled the images to 64x64, since Dino was pretrained on 224x224 images.
-We observed better performance from the pretrained feature extractor when the input image size is closer to the original training image size of Dino.
-We present below the two custom CNN architectures derive from the SimpleCNN architecture (from the Avalanche toolkit).
-All other architectures used for the experiments are available in the Avalanche toolkit.
-The specific CNN architecture used for MNIST experiments is defined as follows:
-MnistCnn(
-  (cnn_layers): Sequential(
-    (0): Conv2d(1, 32, kernel_size=(3, 3), stride=(1, 1))
-    (1): ReLU()
-    (2): Dropout2d(p=0.2, inplace=False)
-    (3): Conv2d(32, 64, kernel_size=(3, 3), stride=(1, 1))
-    (4): ReLU()
-    (5): Dropout2d(p=0.2, inplace=False)
-    (6): MaxPool2d(kernel_size=2, stride=2, padding=0, dilation=1, ceil_mode=False)
-  (linear_layers): Sequential(
-    (0): Linear(in_features=9216, out_features=128, bias=True)
-    (1): ReLU()
-    (2): Dropout(p=0.2, inplace=False)
-    (3): Linear(in_features=128, out_features=10, bias=True)
-The specific CNN architecture used for FMNIST and EMNIST experiments is defined as follows:
-EMNISTFashionMnistCnn(
-  (cnn_layers): Sequential(
-    (0): Conv2d(1, 32, kernel_size=(3, 3), stride=(1, 1))
-    (1): ReLU()
-    (2): BatchNorm2d(32, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
-    (3): Dropout2d(p=0.3, inplace=False)
-    (4): Conv2d(32, 64, kernel_size=(3, 3), stride=(1, 1))
-    (5): ReLU()
-    (6): BatchNorm2d(64, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
-    (7): MaxPool2d(kernel_size=2, stride=2, padding=0, dilation=1, ceil_mode=False)
-    (8): Dropout2d(p=0.3, inplace=False)
-    (9): Conv2d(64, 128, kernel_size=(3, 3), stride=(1, 1))
-    (10): ReLU()
-    (11): BatchNorm2d(128, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
-    (12): MaxPool2d(kernel_size=2, stride=2, padding=0, dilation=1, ceil_mode=False)
-    (13): Dropout2d(p=0.3, inplace=False)
-  (linear_layers): Sequential(
-    (0): Linear(in_features=3200, out_features=256, bias=True)
-    (1): ReLU()
-    (2): Dropout(p=0.3, inplace=False)
-    (3): Linear(in_features=256, out_features=10, bias=True)
-Method ${k}$ ${\omega}$ Epochs Batch size Learning rate
-CoP2L 8 15.0 10 256 0.001
-Replay NA NA 20 128 0.01
-GDumb NA NA 20 128 0.01
-Table 3: Hyperparameter setup for MNIST, FMNIST and EMNIST using a MLP architecture. NA signifies Not Applicable.
-Method ${k}$ ${\omega}$ Epochs Batch size Learning rate
-CoP2L (MNIST & FMNIST) 48 25.0 25 256 0.0009
-CoP2L (EMNIST) 105 20.0 15 130 0.0006
-Replay NA NA 20 128 0.01
-GDumb NA NA 20 128 0.01
-Table 4: Hyperparameter setup for MNIST, FMNIST and EMNIST using a CNN architecture. NA signifies Not Applicable.
-Method ${k}$ ${\omega}$ Epochs Batch size Learning rate
-CoP2L-Dino (ViT) 4 25.0 2 256 0.001
-Replay-Dino (ViT) NA NA 20 128 0.01
-iCaRL-Dino (ViT) NA NA 10 128 0.01
-GDumb-Dino (ViT) NA NA 20 128 0.01
-CoP2L-Dino (RN50) 16 25.0 2 256 0.001
-Replay-Dino (RN50) NA NA 20 128 0.01
-iCaRL-Dino (RN50) NA NA 30 128 0.01
-GDumb-Dino (RN50) NA NA 20 128 0.01
-Table 5: Hyperparameter setup for Class-Incremental setup. NA signifies Not Applicable.
-Method ${k}$ ${\omega}$ Epochs Batch size Learning rate
-CoP2L-Dino (ViT) 4 1.0 2 256 0.001
-Replay-Dino (ViT) NA NA 10 128 0.01
-LaMAML-Dino (ViT) NA NA 10 10 0.01
-LwF-Dino (ViT) NA NA 10 200 0.001
-CoP2L-Dino (RN50) 16 1.0 2 256 0.001
-Replay-Dino (RN50) NA NA 10 128 0.01
-LaMAML-Dino (RN50) NA NA 10 10 0.01
-LwF-Dino (RN50) NA NA 10 200 0.001
-Table 6: Hyperparameter setup for Task-Incremental setup. NA signifies Not Applicable.
-Method ${k}$ ${\omega}$ Epochs Batch size Learning rate
-CoP2L 16 15.0 1 256 0.001
-Replay NA NA 20 128 0.01
-LFL NA NA 3 128 0.01
-EWC NA NA 10 256 0.001
-SI NA NA 10 256 0.001
-Table 7: Hyperparameter setup for Domain-Incremental setup. NA signifies Not Applicable.
-Appendix H Dataset Details
-H.1 Dataset summary
+Appendix E Experiments
+We denote the licensing information of the principal assets used in this project.
+Notably, we use Avalanche [Carta et al., 2023] (MIT License), PyTorch [Ansel et al., 2024] (BSD 3-Clause License) and NumPy [Harris et al., 2020] (NumPy license).
+For the datasets, we use the MNIST dataset [LeCun et al., 1998] (MIT License), the CIFAR10 and CIFAR100 datasets [Krizhevsky, 2012] and subsets of the ImageNet dataset [Deng et al., 2009].
+There is no license information for CIFAR10 and CIFAR100, but it is freely available at https://www.cs.toronto.edu/˜kriz/cifar.html and the authors only ask for their technical report to be cited [Krizhevsky, 2012].
+The ImageNet dataset [Deng et al., 2009] and its subsets (TinyImageNet [mnmoustafa and Ali, 2017], ImageNette [Howard, 2019b] and ImageWoof [Howard, 2019a]) are restricted to non-commercial uses and educational purposes  [^4] The licensing information of ImageNet can be found here : https://image-net.org/download.php.
+E.1 Dataset summary
 Datasets used for class incremental learning experiments:
@@
   • ImageWoof (5 tasks), 5 tasks of 2 classes each (total of 10 classes)
+E.2 Training, architecture details and hyperparameter settings for the experiments
+Most of our experiments are conducted using the Avalanche continual learning framework [Carta et al., 2023], which provides a unified interface for datasets, benchmarks, and baseline implementations.
+We use the SimpleMLP architecture and customized SimpleCNN architectures from the Avalanche toolkit for the MNIST, FMNIST, and EMNIST experiments.
+For CIFAR10, CIFAR100 and TinyImageNet, we use a vision transformer (ViT) [Dosovitskiy et al., 2021], a ResNet50 and a ResNet18 [He et al., 2016].
+The first two models are pretrained with Dino [Caron et al., 2021], frozen and followed by a trainable MLP head.
+ResNet18 was pretrained on ImageNet [Deng et al., 2009], frozen and followed by a trainable MLP head.
+In the Task-Incremental setting, we follow the pretrained models with a linear layer and a MLP head for each task.
+All models are trained with a NVIDIA GeForce RTX 4090 GPU using the SGD optimizer.
+For all experiments, please refer to Tables˜3, 4, 5, 6 and 7 for the hyperparameter configuration.
+We denote the number of samples added to the compression set at each iteration with $k$.
+We denote the weight value used to handle the class-imbalance problem with $\omega$ (Both $k$ and $\omega$ apply only for the CoP2L (our) algorithm).
+We also provide the number of epochs, the training batch size and the learning rate for each experiment.
+For $\rm CoP2L$}{}, we have chosen the aforementioned hyperparameters by optimizing the performance on the validation set, and for the baseline methods, we have used the hyperparameter values reported in their official implementations.
+We use $\gamma=-\ln(0.5)$ as stopping criterion for mP2L, as for the cross-entropy loss, this is equivalent to achieving zero errors on the complement set $S^{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}$.
+In experiments using a Dino backbone (both ResNet50 and ViT backbones), we upsampled the images to 64x64, since the backbones were pretrained on 224x224 images.
+We use feature extractors pre-trained on ImageNet using Dino [Caron et al., 2021], which are available on https://github.com/facebookresearch/dino.
+For experiments with a ResNet18, we use a backbone pretrained on ImageNet.
+For all experiments with pretrained backbones, we freeze the backbone and follow it by a linear head.
+In Task-Incremental learning, we follow the backbone with a linear layer and a multi-head classification layer, that is a layer that uses a different linear head for each task.
+We observed better performance from the pretrained feature extractor when the input image size is closer to the original training image size of Dino.
+Table 3: Hyperparameter setup for MNIST, FMNIST and EMNIST using a MLP architecture.
+Method k{k} $\omega$} Epochs Batch size Learning rate
+CoP2L 88 $15.0$ $10$ $256$ $0.001$
+Replay - - 2020 $128$ $0.01$
+GDumb - - 2020 $128$ $0.01$
+Table 4: Hyperparameter setup for MNIST, FMNIST and EMNIST using a CNN architecture.
+Method k{k} $\omega$} Epochs Batch size Learning rate
+CoP2L (MNIST & FMNIST) $48$ $25.0$ $25$ $256$ $0.0009$
+CoP2L (EMNIST) $105$ $20.0$ $15$ $130$ $0.0006$
+Replay - - 2020 $128$ $0.01$
+GDumb - - 2020 $128$ $0.01$
+Table 5: Hyperparameter setup for Class-Incremental setup.
+Method k{k} $\omega$} Epochs Batch size Learning rate
+CoP2L (ViT) 44 $25.0$ 22 $256$ $0.001$
+Finetuning (ViT) - - 2020 $128$ $0.01$
+DER (ViT) - - 2020 $128$ $0.01$
+Replay (ViT) - - 2020 $128$ $0.01$
+iCaRL (ViT) - - $10$ $128$ $0.01$
+GDumb (ViT) - - 2020 $128$ $0.01$
+CCLIS (ViT) - - $50$ $512$ $1.0$
+CoP2L (RN50) $16$ $25.0$ 22 $256$ $0.001$
+Finetuning (RN50) - - 2020 $128$ $0.01$
+DER (RN50) - - 2020 $128$ $0.01$
+Replay (RN50) - - 2020 $128$ $0.01$
+iCaRL (RN50) - - $30$ $128$ $0.01$
+GDumb (RN50) - - 2020 $128$ $0.01$
+CCLIS (RN50) - - $50$ $512$ $1.0$
+CSReL (RN50) - - $100$ $256$ $0.001$
+Table 6: Hyperparameter setup for Task-Incremental setup.
+Method k{k} $\omega$} Epochs Batch size Learning rate
+CoP2L (ViT) 44 $1.0$ 22 $256$ $0.001$
+Finetuning (ViT) - - $10$ $128$ $0.01$
+DER (ViT) - - $10$ $128$ $0.01$
+Replay (ViT) - - $10$ $128$ $0.01$
+LaMAML (ViT) - - $10$ $10$ $0.01$
+LwF (ViT) - - $10$ $200$ $0.001$
+CCLIS (ViT) - - $50$ $512$ $1.0$
+CoP2L (RN50) $16$ $1.0$ 22 $256$ $0.001$
+Finetuning (RN50) - - $10$ $128$ $0.01$
+DER (RN50) - - $10$ $128$ $0.01$
+Replay (RN50) - - $10$ $128$ $0.01$
+LaMAML (RN50) - - $10$ $10$ $0.01$
+LwF (RN50) - - $10$ $200$ $0.001$
+CCLIS (RN50) - - $50$ $512$ $1.0$
+CSReL (RN50) - - $100$ $256$ $0.001$
+Table 7: Hyperparameter setup for Domain-Incremental setup (MLP and CNN).
+Method k{k} $\omega$} Epochs Batch size Learning rate
+CoP2L $16$ $15.0$ 11 $256$ $0.001$
+Replay - - 2020 $128$ $0.01$
+LFL - - 33 $128$ $0.01$
+EWC - - $10$ $256$ $0.001$
+SI - - $10$ $256$ $0.001$
+E.3 All results with standard deviation and training time
+In this section, we provide the average accuracy values obtained on CIFAR10, CIFAR100, TinyImageNet, ImageNette and ImageWoof datasets with the standard deviation estimates.
+Note that we use different random number generator seeds to account for variability in random number generations (e.g., model parameter initializations) – but train / test splits are kept fixed with respect to seeds across different experiments.
+We present new CI results on MNIST, FMNIST and EMNIST, with additional Domain-Incremental (DI) experiments on PermutedMNIST and RotatedMNIST.
+We evaluate against Replay, Less Forgetful Learning (LFL) [Jung et al., 2018], Elastic Weight Consolidation (EWC) [Kirkpatrick et al., 2017] and Synaptic Intelligence (SI) [Zenke et al., 2017].
+Note that, for CIFAR10, CIFAR100, TinyImageNet, ImageWoof, ImageNette, PermutedMNIST and RotatedMNIST experiments we also report the total time it takes to complete each experiment (for all tasks).
+E.3.1 Class-Incremental Learning problems
+Table 8: Class-incremental learning on CIFAR10 with 5 tasks.
+Methods Average Accuracy (%) Average Forgetting (%) Total time (s)
+CoP2L (ViT) $94.45\pm 0.14$ $2.10\pm 0.28$ $118.07\pm 2.59$
+Finetuning (ViT) $27.55\pm 1.04$ $9.74\pm 1.26 88$$80.59\pm 8.33$
+Replay (ViT) $94.00\pm 0.19$ $6.11\pm 0.25$ $486.25\pm 43.07$
+DER (ViT) $95.03\pm 0.06$ $3.79\pm 0.15$ $436.07\pm 1.21$
+iCaRL (ViT) $88.60\pm 10.04$ $3.13\pm 4.20$ $884.59\pm 1.39$
+GDumb (ViT) $94.16\pm 0.24$ $3.16\pm 0.39$ $47.72\pm 0.12$
+CCLIS (ViT) $93.15\pm 0.34$ $4.44\pm 0.67$ $2527.15\pm 44.56$
+CoP2L (ResNet50) $80.98\pm 0.17$ $5.84\pm 0.37$ $527.95\pm 8.43$
+Finetuning (ResNet50) $19.63\pm 0.02$ $97.22\pm 0.15$ $819.80\pm 2.81$
+Replay (ResNet50) $81.38\pm 0.39$ $17.72\pm 0.51$ $287.38\pm 25.06$
+DER (ResNet50) $82.35\pm 0.22$ $15.31\pm 0.33$ $285.65\pm 0.69$
+iCaRL (ResNet50) $74.49\pm 0.22$ $1.79\pm 0.43$ $579.17\pm 2.39$
+GDumb (ResNet50) $80.84\pm 0.14$ $9.35\pm 0.30$ $31.98\pm 0.12$
+CCLIS (ResNet50) $76.89\pm 0.30$ $12.32\pm 0.33$ $2654.32\pm 52.34$
+CSReL (ResNet50) $34.50\pm 0.29$ $26.44\pm 0.11$ $17284.41\pm 157.43$
+Table 9: Class-incremental learning on CIFAR100 with 10 tasks.
+Methods Average Accuracy (%) Average Forgetting (%) Total time (s)
+CoP2L (ViT) $73.49\pm 0.27$ $17.06\pm 0.29$ $522.18\pm 3.90$
+Finetuning (ViT) $12.87\pm 0.15$ $93.27\pm 0.20$ $901.37\pm 1.73$
+Replay (ViT) $68.71\pm 0.25$ $30.02\pm 0.31$ $453.04\pm 11.93$
+DER (ViT) $72.29\pm 0.53$ $25.11\pm 0.53$ $489.57\pm 7.50$
+iCaRL (ViT) $65.54\pm 2.23$ $13.64\pm 0.64$ $2491.81\pm 2889.82$
+GDumb (ViT) $72.51\pm 0.26$ $11.53\pm 0.65$ $95.43\pm 0.34$
+CCLIS (ViT) $67.95\pm 0.65$ $23.68\pm 0.19$ $3023.03\pm 87.54$
+CoP2L (ResNet50) $49.53\pm 0.24$ $22.31\pm 0.20$ $1724.15\pm 23.50$
+Finetuning (ResNet50) $9.36\pm 0.05$ $89.93\pm 0.02$ $823.61\pm 2.81$
+Replay (ResNet50) $49.41\pm 0.38$ $42.12\pm 0.54$ $295.80\pm 15.88$
+DER (ResNet50) $9.37\pm 0.16 44$$4.01\pm 0.18$ $307.86\pm 13.37$
+iCaRL (ResNet50) $46.06\pm 0.40$ $8.50\pm 0.28$ $676.76\pm 56.25$
+GDumb (ResNet50) $43.37\pm 0.24$ $20.23\pm 0.42$ $63.58\pm 0.36$
+CCLIS (ResNet50) $44.22\pm 0.28$ $23.23\pm 0.10$ $3122.45\pm 78.87$
+CSReL (ResNet50) $38.75\pm 0.25$ $42.12\pm 0.48$ $9154.71\pm 278.30$
+Table 10: Class-incremental learning on CIFAR100 with 20 tasks.
+Methods Average Accuracy (%) Average Forgetting (%) Total time (s)
+CoP2L (ViT) $70.56\pm 0.99$ $21.15\pm 2.10$ $820.66\pm 176.27$
+Finetuning (ViT) $8.49\pm 0.59$ $94.64\pm 0.54$ $902.33\pm 8.58$
+Replay (ViT) $69.21\pm 0.26$ $29.52\pm 0.25$ $457.46\pm 2.07$
+DER (ViT) $77.03\pm 0.34$ $19.27\pm 0.48$ $1096.62\pm 19.46$
+iCaRL (ViT) $62.36\pm 1.46$ $15.88\pm 0.84$ $1359.13\pm 4.58$
+GDumb (ViT) $72.35\pm 0.30$ $11.77\pm 0.40$ $190.80\pm 0.42$
+CCLIS (ViT) $66.81\pm 0.34$ $24.42\pm 0.29$ $3947.38\pm 89.23$
+CoP2L (ResNet50) $46.74\pm 1.63$ $22.54\pm 0.92$ $1771.12\pm 757.50$
+Finetuning (ResNet50) $4.88\pm 0.03$ $92.81\pm 0.31$ $1247.98\pm 506.91$
+Replay (ResNet50) $9.44\pm 0.30 44$$4.96\pm 0.36$ $293.25\pm 1.77$
+DER (ResNet50) $57.53\pm 0.05$ $33.37\pm 0.05$ $510.52\pm 99.48$
+iCaRL (ResNet50) $46.47\pm 0.33$ $6.74\pm 0.31$ $858.11\pm 120.41$
+GDumb (ResNet50) $43.60\pm 0.38$ $19.93\pm 0.56$ $126.74\pm 0.46$
+CCLIS (ResNet50) $39.10\pm 0.04$ $25.78\pm 0.03$ $4021.21\pm 59.34$
+CSReL (ResNet50) $32.57\pm 0.02$ $40.23\pm 0.27$ $13404.29\pm 320.31$
+Table 11: Class-incremental learning on TinyImageNet with 40 tasks.
+Methods Average Accuracy (%) Average Forgetting (%) Total time (s)
+CoP2L (ViT) $51.59\pm 0.24$ $36.32\pm 0.29$ $1451.93\pm 4.71$
+Finetuning (ViT) $2.83\pm 0.17$ $95.36\pm 0.31$ $2304.54\pm 37.18$
+Replay (ViT) $47.49\pm 0.47$ $48.39\pm 0.49$ $1128.97\pm 16.95$
+DER (ViT) $53.92\pm 0.36$ $41.77\pm 0.45$ $1280.76\pm 131.53$
+iCaRL (ViT) $21.06\pm 1.11$ $34.49\pm 9.37$ $3919.23\pm 2170.08$
+GDumb (ViT) $53.93\pm 0.34$ $15.11\pm 0.19$ $471.23\pm 17.85$
+CCLIS (ViT) $48.55\pm 0.46$ $32.26\pm 0.43$ $8000.69\pm 355.45$
+CoP2L (ResNet50) $33.30\pm 0.32$ $35.93\pm 0.27$ $3671.69\pm 29.19$
+Finetuning (ResNet50) $1.80\pm 0.36$ $78.66\pm 1.04$ $2167.76\pm 161.91$
+Replay (ResNet50) $28.38\pm 0.42$ $63.36\pm 0.44$ $882.12\pm 180.18$
+DER (ResNet50) $28.77\pm 0.34$ $64.17\pm 0.34$ $818.11\pm 10.28$
+iCaRL (ResNet50) $33.89\pm 0.47$ $10.42\pm 0.44$ $1905.26\pm 295.18$
+GDumb (ResNet50) $24.29\pm 0.58$ $21.29\pm 0.56$ $332.22\pm 0.98$
+CCLIS (ResNet50) $18.32\pm 0.11$ $40.40\pm 0.02$ $8349.31\pm 378.39$
+CSReL (ResNet50) $12.84\pm 0.22$ $73.69\pm 0.40$ $39644.46\pm 237.23$
+Table 12: Class-incremental learning on MNIST with 5 tasks.
+Methods Average Accuracy (%) Average Forgetting (%)
+CoP2L (MLP) $95.12\pm 0.18$ $3.35\pm 0.24$
+Replay (MLP) $94.48\pm 0.28$ $6.08\pm 0.32$
+GDumb (MLP) $91.70\pm 0.31$ $3.84\pm 0.29$
+CoP2L (CNN) $96.61\pm 0.07$ $2.16\pm 0.28$
+Replay (CNN) $96.93\pm 0.29$ $3.31\pm 0.38$
+GDumb (CNN) $94.41\pm 0.13$ $2.97\pm 0.46$
+Table 13: Class-incremental learning on Fashion-MNIST with 5 tasks.
+Methods Average Accuracy (%) Average Forgetting (%)
+CoP2L (MLP) $84.23\pm 0.27$ $12.10\pm 0.66$
+Replay (MLP) $83.99\pm 0.52$ $15.92\pm 0.42$
+GDumb (MLP) $81.76\pm 0.36$ $10.44\pm 0.50$
+CoP2L (CNN) $87.35\pm 0.15$ $11.12\pm 0.36$
+Replay (CNN) $87.59\pm 0.11$ $13.32\pm 0.29$
+GDumb (CNN) $84.98\pm 0.21$ $8.78\pm 0.41$
+Table 14: Class-incremental learning on EMNIST with 13 tasks.
+Methods Average Accuracy (%) Average Forgetting (%)
+CoP2L (MLP) $78.03\pm 0.17$ $16.67\pm 0.31$
+Replay (MLP) $77.22\pm 0.44$ $22.08\pm 0.50$
+GDumb (MLP) $62.64\pm 0.20$ $15.96\pm 0.58$
+CoP2L (CNN) $87.90\pm 0.41$ $8.38\pm 0.54$
+Replay (CNN) $87.70\pm 0.50$ $12.29\pm 0.49$
+GDumb (CNN) $85.05\pm 0.27$ $7.50\pm 0.61$
+E.3.2 Task-Incremental Learning problems
+Table 15: Task-incremental learning on CIFAR10 with 5 tasks.
+Methods Average Accuracy (%) Average Forgetting (%) Total time (s)
+CoP2L (ViT) $99.04\pm 0.05$ $0.00\pm 0.10$ $64.51\pm 0.68$
+Finetuning (ViT) $99.15\pm 0.06$ $0.23\pm 0.04$ $466.97\pm 11.76$
+Replay (ViT) $99.29\pm 0.04$ $0.01\pm 0.03$ $217.83\pm 0.76$
+DER (ViT) $99.15\pm 0.08$ $0.21\pm 0.05$ $133.94\pm 1.02$
+LaMAML (ViT) $99.25\pm 0.06$ $0.05\pm 0.07$ $344.65\pm 1.46$
+LwF (ViT) $97.89\pm 0.41$ $1.69\pm 0.48$ $165.41\pm 0.75$
+CCLIS (ViT) $98.81\pm 0.89$ $0.48\pm 0.12$ $2534.15\pm 53.56$
+CoP2L (ResNet50) $96.22\pm 0.18$ $0.56\pm 0.28$ $172.51\pm 4.42$
+Finetuning (ResNet50) $96.60\pm 0.08$ $1.00\pm 0.11$ $460.78\pm 3.38$
+Replay (ResNet50) $96.88\pm 0.06$ $0.42\pm 0.13$ $144.57\pm 0.51$
+DER (ResNet50) $96.57\pm 0.10$ $1.09\pm 0.14$ $90.56\pm 0.36$
+LaMAML (ResNet50) $96.65\pm 0.21$ $0.73\pm 0.21$ $310.51\pm 1.64$
+LwF (ResNet50) $95.80\pm 0.15$ $0.33\pm 0.20$ $97.09\pm 0.21$
+CCLIS (ResNet50) $95.45\pm 0.02$ $0.79\pm 0.13$ $2637.16\pm 61.23$
+CSReL (ResNet50) $65.46\pm 0.56$ $13.66\pm 0.19$ $12354.00\pm 160.45$
+Table 16: Task-incremental learning on CIFAR100 with 10 tasks.
+Methods Average Accuracy (%) Average Forgetting (%) Total time (s)
+CoP2L (ViT) $95.03\pm 0.08$ $0.90\pm 0.11$ $290.95\pm 30.98$
+Finetuning (ViT) $95.17\pm 0.11$ $1.84\pm 0.13$ $478.64\pm 6.32$
+Replay (ViT) $96.43\pm 0.14$ $0.34\pm 0.17$ $235.81\pm 1.10$
+DER (ViT) $96.22\pm 0.12$ $0.71\pm 0.13$ $137.12\pm 1.96$
+LaMAML (ViT) $95.58\pm 0.16$ $0.58\pm 0.12$ $397.78\pm 3.40$
+LwF (ViT) $95.09\pm 0.10$ $0.98\pm 0.10$ $173.57\pm 0.63$
+CCLIS (ViT) $90.99\pm 0.45$ $4.49\pm 0.69$ $2532.34\pm 49.57$
+CoP2L (ResNet50) $86.78\pm 0.16$ $2.01\pm 0.17$ $435.10\pm 7.74$
+Finetuning (ResNet50) $88.23\pm 0.09$ $1.78\pm 0.08$ $430.99\pm 2.11$
+Replay (ResNet50) $87.61\pm 0.05$ $0.86\pm 0.16$ $157.05\pm 0.14$
+DER (ResNet50) $88.29\pm 0.20$ $1.72\pm 0.28$ $91.77\pm 1.07$
+LaMAML (ResNet50) $83.57\pm 0.43$ $6.18\pm 0.53$ $357.24\pm 4.58$
+LwF (ResNet50) $79.39\pm 0.13$ $0.47\pm 0.11$ $101.27\pm 0.48$
+CCLIS (ResNet50) $78.55\pm 0.09$ $6.76\pm 0.21$ $2659.12\pm 43.02$
+CSReL (ResNet50) $69.14\pm 0.14$ $15.17\pm 0.41$ $5670.32\pm 256.32$
+Table 17: Task-incremental learning with CIFAR100 with 20 tasks.
+Methods Average Accuracy (%) Average Forgetting (%) Total time (s)
+CoP2L (ViT) $97.18\pm 0.13$ $0.56\pm 0.18$ $226.65\pm 21.43$
+Finetuning (ViT) $97.57\pm 0.23$ $0.86\pm 0.22$ $473.71\pm 6.70$
+Replay (ViT) $98.09\pm 0.09$ $0.25\pm 0.12$ $250.78\pm 1.45$
+DER (ViT) $98.08\pm 0.10$ $0.31\pm 0.17$ $137.98\pm 0.24$
+LaMAML (ViT) $97.84\pm 0.10$ $0.26\pm 0.14$ $448.43\pm 2.81$
+LwF (ViT) $95.74\pm 0.19$ $2.15\pm 0.21$ $180.36\pm 0.98$
+CCLIS (ViT) $94.86\pm 0.54$ $2.88\pm 0.45$ $3951.13\pm 68.94$
+CoP2L (ResNet50) $90.89\pm 0.22$ $2.27\pm 0.28$ $278.67\pm 6.20$
+Finetuning (ResNet50) $93.25\pm 0.04$ $1.27\pm 0.06$ $436.93\pm 2.10$
+Replay (ResNet50) $92.89\pm 0.18$ $0.47\pm 0.24$ $167.65\pm 1.26$
+DER (ResNet50) $93.43\pm 0.02$ $1.11\pm 0.05$ $94.25\pm 0.30$
+LaMAML (ResNet50) $89.89\pm 0.75$ $4.12\pm 0.78$ $407.84\pm 7.38$
+LwF (ResNet50) $87.73\pm 0.18$ $0.33\pm 0.23$ $105.46\pm 0.49$
+CCLIS (ResNet50) $83.51\pm 0.14$ $6.60\pm 0.11$ $4079.41\pm 61.34$
+CSReL (ResNet50) $63.52\pm 0.32$ $14.49\pm 0.22$ $8364.45\pm 289.56$
+Table 18: Task-incremental learning on TinyImageNet with 40 tasks.
+Methods Average Accuracy (%) Average Forgetting (%) Total time (s)
+CoP2L (ViT) $93.81\pm 0.26$ $1.50\pm 0.28$ $1039.14\pm 146.21$
+Finetuning (ViT) $93.62\pm 0.20$ $2.61\pm 0.21$ $983.62\pm 9.49$
+Replay (ViT) $95.66\pm 0.08$ $0.50\pm 0.09$ $632.18\pm 2.97$
+DER (ViT) $95.08\pm 0.07$ $1.13\pm 0.15$ $316.37\pm 0.81$
+LaMAML (ViT) $94.82\pm 0.23$ $0.79\pm 0.19$ $1527.24\pm 891.43$
+LwF (ViT) $85.18\pm 0.76$ $9.72\pm 0.71$ $430.63\pm 2.14$
+CCLIS (ViT) $86.82\pm 0.23$ $6.14\pm 0.56$ $8012.79\pm 321.46$
+CoP2L (ResNet50) $88.57\pm 0.17$ $2.46\pm 0.19$ $890.94\pm 3.44$
+Finetuning (ResNet50) $90.62\pm 0.35$ $1.38\pm 0.40$ $905.9\pm 10.10$
+Replay (ResNet50) $90.03\pm 0.26$ $1.26\pm 0.27$ $514.42\pm 102.82$
+DER (ResNet50) $90.66\pm 0.06$ $1.12\pm 0.15$ $237.61\pm 1.12$
+LaMAML (ResNet50) $87.20\pm 0.34$ $4.53\pm 0.28$ $1268.50\pm 551.93$
+LwF (ResNet50) $85.03\pm 0.17$ $0.89\pm 0.21$ $284.79\pm 5.67$
+CCLIS (ResNet50) $75.65\pm 0.05$ $14.12\pm 0.31$ $8379.89\pm 381.29$
+CSReL (ResNet50) $50.22\pm 0.21$ $22.34\pm 0.51$ $21455.97\pm 231.76$
+Table 19: Task-incremental learning on ImageNette with 5 tasks.
+Methods Average Accuracy (%) Average Forgetting (%) Total time (s)
+CoP2L (ViT) $98.22\pm 0.19$ $0.43\pm 0.12$ $118.71\pm 24.35$
+Finetuning (ViT) $98.24\pm 0.22$ $0.53\pm 0.17$ $221.15\pm 2.16$
+Replay (ViT) $98.61\pm 0.03$ $0.02\pm 0.10$ $356.58\pm 2.89$
+LaMAML (ViT) $98.42\pm 0.02$ $0.03\pm 0.00$ $309.91\pm 5.42$
+LwF (ViT) $96.51\pm 0.67$ $1.97\pm 0.79$ $164.80\pm 1.90$
+Table 20: Task-incremental learning on ImageWoof with 5 tasks.
+Methods Average Accuracy (%) Average Forgetting (%) Total time (s)
+CoP2L (ViT) $93.34\pm 0.30$ $0.65\pm 0.40$ $246.77\pm 13.43$
+Finetuning (ViT) $93.20\pm 0.49$ $2.00\pm 0.84$ $201.91\pm 1.89$
+Replay (ViT) $94.29\pm 0.30$ $0.16\pm 0.26$ $316.87\pm 0.93$
+LaMAML (ViT) $94.00\pm 0.21$ $0.12\pm 0.38$ $276.19\pm 3.45$
+LwF (ViT) $89.75\pm 1.88$ $5.76\pm 2.61$ $175.82\pm 39.30$
+E.3.3 Domain-Incremental Learning problems
+Table 21: Domain-incremental learning on Permuted-MNIST with 10 tasks.
+Methods Average Accuracy (%) Average Forgetting (%) Total time (s)
+CoP2L (MLP) $91.61\pm 0.39$ $6.69\pm 0.40$ $1381.06\pm 1232.99$
+Replay (MLP) $91.28\pm 0.18$ $6.75\pm 0.21$ $1128.99\pm 49.82$
+LFL (MLP) $82.76\pm 1.23$ $10.15\pm 1.36$ $92.86\pm 0.24$
+EWC (MLP) $82.53\pm 0.82$ $6.50\pm 0.88$ $340.16\pm 3.81$
+SI (MLP) $85.85\pm 0.61$ $11.69\pm 0.67$ $319.95\pm 1.38$
+Table 22: Domain-incremental learning on Rotated-MNIST on 10 tasks.
+Methods Average Accuracy (%) Average Forgetting (%) Total time (s)
+CoP2L (MLP) $89.40\pm 1.00$ $8.71\pm 1.04$ $4970.86\pm 1452.34$
+Replay (MLP) $88.75\pm 0.67$ $9.53\pm 0.72$ $3894.62\pm 75.11$
+LFL (MLP) $53.09\pm 7.06$ $44.18\pm 7.93$ $387.61\pm 160.54$
+EWC (MLP) $55.59\pm 5.57$ $36.58\pm 6.27$ $1133.19\pm 18.10$
+SI (MLP) $59.90\pm 6.48$ $41.38\pm 7.18$ $1043.86\pm 7.28$
+E.4 Bound Plots
+E.4.1 Bound plots for different buffer sizes
+This section presents the computed bound values over task risks on MNIST, Fashion-MNIST and EMNIST datasets (letters) in Class-Incremental settings.
+Figure˜4 shows results for different replay buffer sizes (1000, 2000, 3000, 4000, 5000) obtained using the MLP architecture.
+Figure˜5 shows the results obtained with the CNN architecture.
+We report the estimated deviation of the bound values over different seeds using the shaded regions which show the $\pm$1 standard deviation, and also the whiskers (plotted using the seaborn toolkit’s boxplot function) which show the spread of the values excluding the outliers.
+This plotting style for showing the deviation across seeds applies to all bound plots presented in this paper.
+Figure 4: CoP2L Bounds with respect to tasks on the MNIST, Fashion-MNIST and EMNIST datasets using an MLP architecture
+Figure 5: CoP2L Bounds with respect to tasks on the MNIST, Fashion-MNIST and EMNIST datasets using a CNN architecture
+We observe from Figures 4 and 5 that the MLP architecture and the CNN architecture give non-trivial bounds that follow the trends exhibited by the test error curve.
+E.4.2 Bound plots for CIFAR10, CIFAR100, TinyImageNet and ImageNet (subsets) datasets
+In this section, we present the plots for the bound values obtained for the additional datasets, namely CIFAR10, CIFAR100, and ImageNet subsets.
+The ImageNet subsets include the Tinyimagenet [mnmoustafa and Ali, 2017], ImageWoof [Howard, 2019a], and ImageNette [Howard, 2019b] datasets.
+We provide the results on ImageWoof and ImageNette to showcase that the estimated bounds get tighter when more data is avaible per each task.
+We note that as the pretrained models used as backbones for our experiments were pretrained on the ImageNet, we provide the bounds on the ImageNet subsets simply to give insight to the reader.
+Indeed, the dependency of the backbones to the dataset over which we certify breaks the i.i.d.
+assumption of the bound.
+Figure 6: CoP2L bounds with respect to tasks on the CIFAR10 (5 tasks) dataset in Class-Incremental (left) and Task-Incremental (right) settings
+Figure 7: CoP2L bounds with respect to tasks on the CIFAR100 (10 tasks) dataset in Class-Incremental (left) and Task-Incremental (right) settings
+Figure 8: CoP2L bounds with respect to tasks on the CIFAR100 (20 tasks) dataset in Class-Incremental (left) and Task-Incremental (right) settings
+Figure 9: CoP2L bounds with respect to tasks on the TinyImageNet (40 tasks) dataset in Class-Incremental (left) and Task-Incremental (right) settings
+Figure 10: CoP2L bounds with respect to tasks on the ImageNette (5 tasks) dataset (left) and ImageWoof dataset (right) in Task-Incremental setting
+Figure 11: CoP2L bounds with respect to tasks on the PermutedMNIST (10 tasks) dataset (left) and on the RotatedMNIST dataset in Domain-Incremental setting
+Appendix F Analysis of CoP2L
+In this section we provide multiple analyses and ablation studies of CoP2L.
+F.1 Ablation on CoP2L loss terms.
+We have conducted an ablation study to understand the effect of additional weighting that counteracts class imbalance that arises in the continual learning setup when new tasks are introduced (Standard replay counteracts this by extending the batch size to have equal class representation).
+In the modified P2L (Algorithm 2), we incorporate an additional weighting term over classes to adapt P2L for class-balanced continual learning.
+In order to assess whether this idea is indeed helpful, we have run an experiment on the MNIST dataset with and without the additional weighting idea.
+In Figure 12, we show the results of this experiment.
+For both accuracy (left panel) and forgetting (right panel), we observe that the inclusion of the additional loss weighting term significantly improves the results.
+Figure 12: Ablation study on MNIST and Fashion-MNIST on the additional weighting to ensure class balance we propose in Algorithm 2 on the CoP2L loss function.
+Accuracy (left) and forgetting (right) over replay buffer size for the MNIST dataset
+F.2 Sensitivity analysis with respect to the task ordering.
+In order to also assess the sensitivity of CoP2L with respect to ordering of the tasks, on MNIST and Fashion-MNIST experiments we have conducted an experiment where we randomly shuffle the task order.
+For both datasets, we have tested our algorithm with 5 different random task ordering.
+For each task, we report the cumulative accuracy such that, $\text{Average Cumulative Acc.}_{T}=\frac{1}{O\cdot T}\sum_{o=1}^{O}\sum_{t=1}^{T}\text{Accuracy}_{T,t}^{o}$, where $\text{Accuracy}_{T,t}^{o}$ denotes the accuracy obtained at task $t$ after finishing training until task $T$, and the $o$ index denotes the task order.
+We have used $O=5$ different task orderings.
+We present the results of this experiment in Figure 13.
+On both MNIST and Fashion-MNIST datasets CoP2L is able to outperform with more tasks.
+We also observe that the variability of the results decrease more significantly for CoP2L compared to replay.
+Figure 13: Studying the variation of cumulative accuracy with respect to task ordering for CoP2L and Replay on MNIST dataset (left) and Fashion-MNIST dataset (right).
+We use 5 different task order, and show the variability with respect to each task.
+F.3 Analysis with respect to buffer size
+In Fig.
+14, we compare the accuracy and forgetting obtained with replay buffer sizes of 1000–5000 samples for CoP2L and replay.
+This corresponds to roughly 8–40% of the per-task dataset size for MNIST and Fashion-MNIST, and 10–52% for EMNIST.
+We report the final average accuracy (after T=5T{=}5 tasks for MNIST and Fashion-MNIST, and T=13T{=}13 for EMNIST) across seeds as well as per-task test accuracy.
+For MNIST, CoP2L clearly outperforms replay on average.
+For Fashion-MNIST and EMNIST, CoP2L shows a clear advantage for small buffers and remains comparable or slightly better for larger buffers.
+Across all datasets, CoP2L substantially reduces forgetting.
+Figure 14: Accuracy (left) and forgetting (right) vs. buffer size for MNIST (top), Fashion-MNIST (middle), and EMNIST (bottom). Orange triangles indicate mean performance.
+F.4 Analysis of memory cost
+We have conducted a memory comparison on Split-CIFAR10 (5 tasks) in Class Incremental Learning (CIL) and Task Incremental Learning (TIL) settings, as shown in Table˜23.
+We report Maximum allocated memory (actual usage), maximum reserved memory (allocated memory PyTorch’s caching allocator), and also the average cumulative memory over all tasks.
+We would like to note that, in the current implementation, we execute the sample compression in one go, by processing all data items that correspond to the current task in parallel.
+It is totally possible to implement this with minibatching without loss of performance.
+The table reports, in parentheses, the approximate memory that would be used by CoP2L using the same minibatch size as the reported Replay.
+Table 23: Memory usage comparison for CoP2L and Replay in CIL and TIL settings.
+Method Max Mem allocated (MB) Max Mem reserved (MB) Cumulate Mem usage (TB)
+CIL
+CoP2L 15535.57 $\pm$ 87.53 (342.85 $\pm$ 6.34) 22310.0 $\pm$ 104.67 (492.36 $\pm$ 7.98) 1261.38 $\pm$ 31.72
+Replay 458.76 $\pm$ 9.84 648.0 $\pm$ 12.35 3383.43 $\pm$ 74.51
+TIL
+CoP2L 15535.66 $\pm$ 98.34 (342.86 $\pm$ 6.78) 22310.0 $\pm$ 102.65 (492.36 $\pm$ 7.81) 197.18 $\pm$ 6.42
+Replay 458.85 $\pm$ 9.91 648.0 $\pm$ 12.41 1783.11 $\pm$ 52.63
+F.5 Ablation on the use of early stopping
+In this section, we study the use of early stopping in our CoP2L algorithm.
+We present the result in Table˜24.
+The early stopping improves the accuracy, the forgetting and the bound on each task except for the bound on Task 1 with the ViT.
+This is to be expected as the early stopping helps to alleviate the overfitting of Pick-To-Learn by selecting the checkpoint that achieves the best bound once the algorithm has effectively converged.
+In contrast, without early stopping, training proceeds until the training error reaches zero, which can lead to overfitting, bigger compression set and consequently worse generalization metrics.This can also explain the tighter bounds, as it was demonstrated by Bazinet et al.
+[2025] that, on binary subsets of the MNIST dataset, the sample-compression bound was actually minimized about halfway before Pick-To-Learn converges.
+Table 24: Illustration of the behavior of the bound on CIFAR10 after 5 tasks, with and without early stopping using the bound.
+Backbone Accuracy $(\uparrow)$ Forgetting $(\downarrow)$ Bound on Task 1 $(\downarrow)$ Bound on Task 2 $(\downarrow)$ Bound on Task 3()($\downarrow$) Bound on Task 4()($\downarrow$) Bound on Task 5()($\downarrow$) With early stopping RN18 68.12 $\pm$ 0.42 17.62 $\pm$ 0.31 82.34 $\pm$ 0.55 94.62 $\pm$ 0.47 88.17 $\pm$ 0.36 66.15 $\pm$ 0.28 53.52 $\pm$ 0.44 RN50 80.98 $\pm$ 0.25 5.84 $\pm$ 0.18 62.97 $\pm$ 0.63 83.46 $\pm$ 0.39 78.17 $\pm$ 0.41 58.03 $\pm$ 0.22 51.02 $\pm$ 0.35 ViT 94.45 $\pm$ 0.12 2.10 $\pm$ 0.09 31.02 $\pm$ 0.48 48.17 $\pm$ 0.33 40.66 $\pm$ 0.27 25.45 $\pm$ 0.19 21.30 $\pm$ 0.26 Without early stopping RN18 65.30 $\pm$ 0.51 22.71 $\pm$ 0.37 90.02 $\pm$ 0.46 96.67 $\pm$ 0.40 91.20 $\pm$ 0.35 71.06 $\pm$ 0.29 59.15 $\pm$ 0.43 RN50 77.36 $\pm$ 0.34 12.11 $\pm$ 0.26 78.05 $\pm$ 0.52 91.65 $\pm$ 0.38 83.52 $\pm$ 0.31 60.04 $\pm$ 0.24 53.28 $\pm$ 0.36 ViT 94.34 $\pm$ 0.11 2.63 $\pm$ 0.07 29.46 $\pm$ 0.45 49.46 $\pm$ 0.30 42.16 $\pm$ 0.28 25.99 $\pm$ 0.21 23.72 $\pm$ 0.25
+F.6 Class-wise accuracy
+We report the accuracies obtained over earlier tasks with CoP2L and with replay.
+We observe that CoP2L is able to obtain a more balanced performance over earlier tasks compared to standard replay, as can be seen from the last rows of Table 25, and Table 26.
+Table 25: The accuracies obtained on earlier tasks with CoP2L on CIFAR10.
+Class 1 Class 2 Class 3 Class 4 Class 5 Class 6 Class 7 Class 8 Class 9 Class 10 Task 1 98.10 $\pm$ 0.22 97.30 $\pm$ 0.31 Task 2 98.60 $\pm$ 0.18 96.80 $\pm$ 0.27 83.70 $\pm$ 0.45 85.50 $\pm$ 0.36 Task 3 95.70 $\pm$ 0.29 96.70 $\pm$ 0.21 79.40 $\pm$ 0.52 77.40 $\pm$ 0.47 82.90 $\pm$ 0.33 71.10 $\pm$ 0.58 Task 4 93.50 $\pm$ 0.34 97.00 $\pm$ 0.19 72.60 $\pm$ 0.49 68.70 $\pm$ 0.55 84.10 $\pm$ 0.28 71.20 $\pm$ 0.46 84.40 $\pm$ 0.31 80.10 $\pm$ 0.37 Task 5 85.30 $\pm$ 0.41 91.50 $\pm$ 0.26 71.40 $\pm$ 0.53 66.50 $\pm$ 0.60 80.20 $\pm$ 0.35 69.70 $\pm$ 0.48 90.30 $\pm$ 0.24 83.60 $\pm$ 0.39 89.50 $\pm$ 0.30 83.10 $\pm$ 0.42
+Table 26: The accuracies obtained on earlier tasks with standard replay algorithm on CIFAR10.
+Class 1 Class 2 Class 3 Class 4 Class 5 Class 6 Class 7 Class 8 Class 9 Class 10 Task 1 98.10 $\pm$ 0.24 98.40 $\pm$ 0.19 Task 2 93.10 $\pm$ 0.38 97.20 $\pm$ 0.27 94.70 $\pm$ 0.33 92.40 $\pm$ 0.41 Task 3 93.80 $\pm$ 0.36 96.80 $\pm$ 0.22 74.90 $\pm$ 0.55 53.80 $\pm$ 0.61 94.70 $\pm$ 0.29 90.50 $\pm$ 0.34 Task 4 91.70 $\pm$ 0.40 96.50 $\pm$ 0.25 70.90 $\pm$ 0.57 57.20 $\pm$ 0.63 69.80 $\pm$ 0.48 76.20 $\pm$ 0.45 96.30 $\pm$ 0.21 95.60 $\pm$ 0.26 Task 5 77.30 $\pm$ 0.52 77.90 $\pm$ 0.49 73.00 $\pm$ 0.54 59.90 $\pm$ 0.60 75.10 $\pm$ 0.47 80.90 $\pm$ 0.39 92.00 $\pm$ 0.28 87.70 $\pm$ 0.36 98.20 $\pm$ 0.18 96.10 $\pm$ 0.23
+F.7 Plasticity forgetting tradeoff
+Denoting $A(t,\theta_{t})$ the accuracy obtained on a task $t$ of a model $f_{\theta_{t}}$ after finishing training on task $t$, the plasticity is given by
+$\overline{P}(\theta_{T})\,{=}\,\frac{1}{T}\sum_{t=1}^{T}A(t,\theta_{t})\,$,
+We observe that CoP2L is able to retain a low level of forgetting while having a relatively good level of plasticity.
+As shown in Table 27, CoP2L consistently attains substantially lower forgetting than Finetuning, Replay, and Dark Experience Replay, and also outperforms GDumb on this metric.
+Although iCaRL achieves slightly lower forgetting, this comes at the cost of reduced plasticity.
+Overall, this trade-off highlights the favorable balance achieved by CoP2L.
+We observe the same behavior across both RN50 and ViT backbones.
+Table 27: Plasticity vs Forgetting Tradeoff on CIFAR10 dataset
+Backbone Method Plasticity ($\uparrow$) Forgetting ($\downarrow$)
+RN50 CoP2L 85.57 $\pm$ 0.34 05.55 $\pm$ 0.28
+RN50 Finetuning 97.21 $\pm$ 0.22 97.22 $\pm$ 0.41
+RN50 Replay 95.50 $\pm$ 0.37 17.09 $\pm$ 0.33
+RN50 iCaRL 75.79 $\pm$ 0.45 01.54 $\pm$ 0.19
+RN50 Gdumb 87.86 $\pm$ 0.31 08.95 $\pm$ 0.36
+RN50 DER 94.78 $\pm$ 0.27 15.60 $\pm$ 0.42
+ViT CoP2L 96.33 $\pm$ 0.18 03.36 $\pm$ 0.21
+ViT Finetuning 99.36 $\pm$ 0.09 89.74 $\pm$ 0.38
+ViT Replay 98.82 $\pm$ 0.14 06.01 $\pm$ 0.26
+ViT iCaRL 94.48 $\pm$ 0.23 01.55 $\pm$ 0.17
+ViT Gdumb 96.86 $\pm$ 0.20 03.39 $\pm$ 0.24
+ViT DER 98.09 $\pm$ 0.16 03.86 $\pm$ 0.29